### v2

# Characteristic initial value problems for the Einstein-Maxwell-scalar field equations in spherical symmetry

Abstract
The characteristic initial boundary problem is discussed in spherical symmetry for the Einstein-Maxwell-scalar field equations. It is formulated for an affine-null metric and the resulting field equations are cast into a hierarchical system of partial differential equations. The initial boundary value problem for a family of null hypersurfaces is specified for a timelike-null foliation at the central geodesic of spherical symmetry as well as for a double-null foliation where the corresponding boundary is a null hypersurface. For the latter, two distinct boundary value formulations arise – one where the null boundary has zero Misner-Sharp mass and another one where the corresponding Misner-Sharp mass is nonzero. As an application, the nonextremal and the extremal Reissner-Nordström solution in null coordinates for a charged black hole and the Fisher-Janis-Newman-Winicour solution are derived.
I Introduction
The choice of spacetime foliation with correspondingly adapted coordinates to solve the Einstein field equations determines whether the resulting partial differential equations are either a complete set of hyperbolic partial differential equations or a mixed hyperbolic/elliptic system. The hypersurfaces of the hyperbolic part of this system along which the information travels are the characteristic surfaces. Characteristic surfaces in General Relativity may be for example null cones along which gravitational waves propagate or null hypersurfaces tied to the spacetime’s causal structure. Solution methods for Einstein equations using the geometric information of the characteristic are commonly referred to as characteristic methods. The employment of characteristic methods to study the field equations of General Relativity very often brought a deeper understanding to various aspects of the theory. Three important examples are the discovery of the Bondi mass loss formula by Bondi, Sachs and collaborators Bondi (1960); Bondi et al. (1962); Sachs (1962a), the first long term stable numerical evolutions of a black hole Gómez et al. (1998) via computer simulations and recently the proof of incorrectness of the third law of black hole thermodynamics Kehle and Unger (2022). The Bondi mass loss formula was found in an asymptotic patch of a spacetime where the radial coordinate parametrizes the null rays of outgoing null cones. The long term stable black hole simulations were possible by numerically representing the spacetime manifold with a countable number of finite size patches discretized by finite difference operators adapted to either in- or outgoing null hypersurfaces. The technical base of the disproof of the third law was the novel technique called “characteristic gluing” Aretakis et al. (2021), where different characteristic (null) hypersurfaces are stitched together using gauge invariant quantities. These three examples have in common that the patch(es) is (are) charted with at least one null coordinate, i.e. a coordinate constant along a null hypersurface, and that boundary information at one side of the patch needs to be provided to archive the physical result. For the Bondi mass loss case, this boundary data is the specification of an asymptotic inertial observer, an initial value for the mass and the so-called news function which is the retarded time derivative of the gravitational shear. For the black hole simulation, the boundary data were the specification of black hole data at an inner world tube excising the black hole singularity, where the first patch of the numerically discretized null hypersurface is located. While in the disproof of the third law, the boundary data between the patches are specification of quantities in Minkowski space and the horizon of a Schwarzschild/Reissner-Nordström black hole. The difference for the latter with respect to the former two is that the employed initial boundary value problem is based on the two intersecting null hypersurfaces instead of only one null hypersurface. A further difference is that in the last example the Einstein equations are solved with a coupling to a charged scalar field, while the first two concerned the full Einstein equations in vacuum. Although in spherical symmetry, the coupling to the scalar field captures dynamical features and adds the necessary degrees of freedom so that the discussion of a characteristic initial boundary value problem (CIBVP) can be made without loosing too much structural information.
Here, we investigate characteristic methods for Einstein equations minimally coupled to a charged, massless and necessary complex scalar field in spherical symmetry. In contrast to the numerous works Angelopoulos et al. (2024); Kehle and Unger (2024a, b, 2022); Kehle and Van de Moortel (2024) related to characteristic gluing Aretakis et al. (2021) as well as the studies of formation of sub-extremal black hole formation of Reisner-Nordström black holes Murata et al. (2013); Gelles and Pretorius (2025) with characteristic methods using a double null foliation, we will investigate the initial boundary value problem in affine null metric coordinates Winicour (2013). The affine null metric formulation is build on a foliation with respect to a family of null hypersurfaces and has the advantage that a simple evolution algorithm can be found, which is beneficial for numerical work. For this evolution scheme, the field equations can be cast into one set of hierarchical hypersurface equations and one first order transport equation with respect to the chosen null coordinate. The set of derived equations is similar to a previously used one for a real scalar field Crespo et al. (2019) having the additional advantage that it can be used to investigate black hole spacetimes beyond the formation of an event horizon and track the apparent horizon and the scalar field inside the black hole Mädler et al. (2024). This interior black hole evolution is possible because the hypersurface-evolution equations remain regular where the expansion rate of the outgoing null rays vanishes, i.e. at an apparent horizon. Here, we extend and modify the scheme of Crespo et al. (2019) such that this regularisation is maintained if the scalar field is complex and a Maxwell field is added to the energy momentum tensor. We will also show that for the derived set of equations, CIBVPs can be formulated either on the central world line in spherical symmetry or by two different versions of double null CIBVPs at a null boundary. As an example for these cases, we discuss the asymptotic CIBVP showing that the entire asymptotic behaviour of a charged scalar field is governed by the specification of an initial value for the scalar field monopole, mass, charge and a free function - the scalar field news function. We further derive the subextremal and extremal Reisner-Nordström black hole as well as the Fisher-Janis-Newman-Winicour (FJNW) solution in the affine null-metric coordinate chart. Those solution turn out to be expressible by simple rational functions.
In Sec. II, the basic field equations and additional geometric quantities for the Einstein-Maxwell scalar field equations in spherical symmetry and affine null metric coordinates are presented. The regularised set of hypersurface-evolution equations following from the field equations being the CIBVPs is given in Sec. III. The CIBVP in an asymptotic region is presented in III.1, the one at the central geodesic is formulated in III.2, while the two different CIBVPs with a null boundary surface are given in Sec. III.3. In Sec. IV, we discuss the properties for asymptotic timelike Killing vectors, which are needed for the derivation of the Reisner-Nordström black hole in Sec. V.1 and the FJNW solution in Sec. V.2. Sec. VI contains a summary and conclusion of the work.
We use geometric units with $c=1$ and the Einsteinian gravitational constant is $\kappa=8\pi G$. The metric signature is $+2$, capital latin indices $A,B,C,...$ are two dimensional with values $x^{A}=(x^{2},x^{3})$ and lowercase latin indices are four dimensional taking values $x^{a}=(w,\lambda,x^{A})$. The covariant derivative of the metric is $\nabla_{a}$ and partial derivatives are denoted with the comma notation $\frac{\partial f}{\partial x^{a}}=f_{,a}$. Calligraphic capital letters, $\mathcal{J},\mathcal{L},\mathcal{C},...$ are always assumed complex fields, while Roman capital letters, like $Q,Z,V,L,\ldots$ are purely real. Numeral subscripts in parenthesis, like $V_{(1)}$, correspond to expansion coefficients of a power series evaluated at $\lambda=0$ with the number in the parenthesis referring to the power $\lambda^{k}$. Numeral subscripts in a bracket, like $\Phi_{[1]}$, correspond to expansion coefficients of an asymptotic power series and the number in the bracket refers to the order of the inverse power $1/\lambda^{k}$.
II Field equations for the Einstein-Maxwell-scalar field equations
In a four dimensional spacetime, consider the Einstein-Hilbert action $S$ given by
$S=\int d^{4}x\sqrt{-g}\left[\frac{g^{ab}R_{ab}}{2\kappa}-\frac{F_{ab}F^{ab}}{16\pi}-\frac{1}{2}(\overline{\mathcal{D}^{a}\Phi})(\mathcal{D}_{a}\Phi)\right]$ (1)
where $\kappa=8\pi G$ is the Einsteinian gravitational constant, $\sqrt{-g}$ the determinant of the metric $g_{ab}$ with well defined inverse $g^{ab}$ and corresponding Levi-Civita covariant derivative $\nabla_{c}$, $R_{ab}$ is the Ricci tensor wrt $g_{ab}$, $F_{ab}=\nabla_{a}A_{b}-\nabla_{b}A_{a}$ is a Maxwell field determined by the 4-covector $A_{b}$, $\mathcal{D}_{c}=\nabla_{c}+iqA_{c}$ a gauge invariant derivative for the massless complex scalar field $\Phi$ that is coupled to the Maxwell field with the gauge charge $q$. The overbar denotes complex conjugation. From the variation of the action with respect to $g^{ab}$, $A_{a}$, $\bar{\Phi}$ and $\Phi$, we can derive the trace reversed Einstein equations, $E_{ab}$, the Maxwell equations, $E_{a}$, and the scalar wave equations $E$ and $\bar{E}$, which are respectively
$E_{ab}:=$ 00 0 $=R_{ab}-\frac{\kappa}{4\pi}\left(g^{cd}F_{ac}F_{bd}-\frac{1}{4}g_{ab}F^{cd}F_{cd}\right)$ (2)
$-\kappa\left[\frac{1}{2}(\overline{\mathcal{D}_{a}\Phi})(\mathcal{D}_{b}\Phi)+\frac{1}{2}(\overline{\mathcal{D}_{b}\Phi})(\mathcal{D}_{a}\Phi)\right]$
$E_{a}:=$ 00 0 $=\nabla_{a}F^{a}_{\;b}+4\pi j_{b}$ (3)
$E:=$ 00 0 $={\mathcal{D}^{a}}\mathcal{D}_{a}\Phi$ (4)
$\bar{E}:=$ 00 0 $=\overline{\mathcal{D}^{a}\mathcal{D}_{a}\Phi}$ (5)
with the current
$j_{b}=\frac{iq}{2}\left(\overline{\Phi}\mathcal{D}_{a}\Phi-\Phi\overline{\mathcal{D}_{a}\Phi}\right)$ (6)
[…]
with $g_{AB}$ being the 2-metric of the spacelike hypersurface of the co-dimension 2 boundary $\Sigma=\partial\mathtt{V}$ where the integral curves of null vectors $\ell^{a}\partial_{a}$ and $n^{a}\partial_{a}$ orthogonal to $\Sigma$ intersect. Let us consider the metric
$ds^{2}=-V(w,\lambda)dw^{2}+2\epsilon dwd\lambda+r^{2}(w,\lambda)q_{AB}dx^{A}dx^{B}$ (8)
where $q_{AB}$ is any metric of a unit sphere [^2] In standard spherical coordinates $x^{A}=(\theta,\varphi)$, $q_{AB}=\mathrm{diag}(1,\sin^{2}\theta)$ with the three associated Killing vectors $\xi_{1}=\partial_{\phi}$, $\xi_{2}=\cos\phi\partial_{\theta}-\cot\theta\sin\phi\partial_{\phi}$ and $\xi_{3}=\sin\phi\partial_{\theta}+\cot\theta\cos\phi\partial_{\phi}$. In principle, this two metric could also be a flat metric, like ${q_{AB}=\mathrm{diag}(1,1)}$, or a unit hypersphere ${q_{AB}=\mathrm{diag}(1,\sinh^{2}\theta)}$. Of the field equations, only (17) would change by the substitution of the term +1 by either $+0$ or $-1$, respectively. The former may be interesting for plane wave front propagation, like pp waves, while the latter for hyperbolical black holes or fluids, see e.g. Herrera et al. (2020, 2021).. An understanding of the the binary parameter $\epsilon$ can be found in flat space where $V=1$ and $r=\lambda$, then for $\epsilon=-1$ the coordinate surfaces $w=const$ represents outgoing null hypersurfaces $u$, while for $\epsilon=1$ they are ingoing null hypersurfaces $v=const$.
The inverse metric has the components
$g^{w\lambda}=\epsilon\;\;,\;\;g^{\lambda\lambda}=V\;\;,\;\;g^{AB}=\frac{1}{r^{2}}q^{AB}\;,$ (9)
where we used $\epsilon^{2}=1$. Furthermore, the form of the metric (8) remains invariant under the coordinate changes
[…]
For the metric we define two null vectors
$\ell^{a}\partial_{a}=\partial_{\lambda}\;\;,\;\;n^{a}\partial_{a}=-\epsilon\partial_{w}-\frac{V}{2}\partial_{\lambda}\;,$ (11)
whose associated one forms are $\ell_{a}dx^{a}=\epsilon dw$ and $n_{a}dx^{a}=\epsilon(V/2)dw-d\lambda$ and which are orthogonal to one another $n^{a}\ell_{a}=-1$. The null vector field $\ell^{a}\partial_{a}$ is tangent to affine parameterized null geodesics meaning $\ell^{a}\nabla_{a}\ell^{a}=0$, while the vector field $n^{a}\partial_{a}$ is in general tangent to non-affine parametrized null geodesics where
$n^{a}\nabla_{a}n^{b}=-\frac{1}{2}V_{,\lambda}n^{a}\;\;.$ (12)
[…]
$Q=-\epsilon r^{2}\alpha_{,\lambda}.$ (13)
The non-zero components of $E_{ab}$ are
$(E_{ww})$ ::: $0=r_{,ww}-\frac{V(r^{2}V_{,\lambda})_{,\lambda}}{4r}-\frac{\epsilon}{2}\left(r_{,w}V_{,\lambda}-r_{,\lambda}V_{,w}\right)\;,+\frac{1}{2}\kappa r\left[|\Phi_{,w}|^{2}+\frac{VQ^{2}}{8\pi r^{4}}+q^{2}\alpha^{2}|\Phi|^{2}+iq\alpha\Bigl{(}\Phi\bar{\Phi}_{,w}-\bar{\Phi}\Phi_{,w}\Bigr{)}\right]\;,$
$(E_{w\lambda})\;$ ::: $0=r_{,w\lambda}+\frac{\epsilon}{4r}(r^{2}V_{,\lambda})_{,\lambda}-\frac{1}{2}\kappa r\left[\frac{\epsilon Q^{2}}{8\pi r^{4}}-\frac{1}{2}[\overline{\Phi}_{,w}\Phi_{,\lambda}+\overline{\Phi}_{,\lambda}\Phi_{,w}]-\frac{iq\alpha}{2}(\Phi\overline{\Phi}_{,\lambda}-\overline{\Phi}\Phi_{,\lambda})\right]\;,$
$(E_{\lambda\lambda})\;$ ::: $0=r_{,\lambda\lambda}+\frac{1}{2}\kappa r|\Phi_{,\lambda}|^{2}\;,$ (16)
$(q^{AB}E_{AB})\;$ ::: $0=-\left(Vrr_{,\lambda}+2\epsilon rr_{,w}\right)_{,\lambda}+1-\frac{\kappa}{8\pi}\frac{Q^{2}}{r^{2}}$ (17)
and those of $E_{a}$
$(E_{w})\;$ ::: $0=Q_{,w}+\epsilon VQ_{,\lambda}+4\pi q^{2}r^{2}\alpha|\Phi|^{2}+2i\pi qr^{2}\Bigl{(}\Phi\bar{\Phi}_{,w}-\bar{\Phi}\Phi_{,w}\Bigr{)}\;,$ (18)
$(E_{\lambda})\;$ ::: $0=Q_{,\lambda}+2\pi r^{2}qi(\overline{\Phi}\Phi_{,\lambda}-\Phi\overline{\Phi}_{,\lambda})\;,$ (19)
while $E$ gives
$0=(r^{2}\Phi_{,w})_{,\lambda}+(r^{2}\Phi_{,\lambda})_{,w}+\epsilon(r^{2}V\Phi_{,\lambda})_{,\lambda}+2iqr\alpha(r\Phi)_{,\lambda}-i\epsilon q\Phi Q\;.$ (20)
In spherical symmetry the Misner-Sharp mass Misner and Sharp (1964)
$M=\frac{r}{2}(1-g^{ab}r_{,a}r_{b})=\frac{r}{2}\Bigl{(}1-2\epsilon r_{,w}r_{,\lambda}-Vr_{,\lambda}^{2}\Bigr{)}$ (21)
is an invariant concept of quasilocal mass. Assuming asymptotically flat spacetimes, its asymptotic limit
$m_{B}:=\lim_{\lambda\rightarrow\infty}M$ (22)
is the Bondi mass $m_{B}$. With the complex null vector $m^{a}\partial_{a}=r^{-1}q^{A}\partial_{A}$ in which the complex dyad $q^{A}\partial_{A}$ is any dyad for the unit sphere metric normalized like ${q^{A}q_{A}=q^{A}\bar{q}_{A}-1=0}$, we calculate the Weyl scalar ${\Psi_{2}=C_{abcd}\ell^{a}m^{b}\bar{m}^{c}n^{d}}$ Griffiths and Podolsky (2009) like
$\Psi_{2}=$ $-\frac{1}{6r^{2}}-\frac{\epsilon}{3}\left(\frac{r_{,\lambda}}{r}\right)_{,w}-\frac{V}{6}\left(\frac{r_{,\lambda}}{r}\right)_{,\lambda}+\frac{r^{2}}{12}\left(\frac{V_{,\lambda}}{r^{2}}\right)_{,\lambda}\;,$ (23)
whose asymptotic limit provides an alternative way to calculate the Bondi mass in a frame approaching an inertial frame at large values for the affine parameter
$m_{B}=-\lim_{\lambda\rightarrow\infty}\left[\lambda^{3}\Psi_{2}-\frac{1}{12}\kappa\epsilon\lambda^{2}\left(\Phi\bar{\Phi}_{,w}+\Phi_{,w}\bar{\Phi}\right)\right]\;.$ (24)
III The characterisitic initial-boundary value problem
The formulation of the an initial boundary value problem for (II) - (20) is a consequence of the twice contracted Bianchi identities, $\nabla^{a}(E_{ab}-\frac{1}{2}g_{ab}(g^{cd}E_{cd}))$ and its electromagnetic counterpart $\nabla^{a}E_{a}$ Tamburino and Winicour (1966) and it follows the idea of Bondi et al. (1962); Sachs (1962a, b). The spirit is to pick from $(E_{ab},E_{a},E)$ a set of main equations to hold everywhere, then the twice contracted Bianchi identities and this set of main equation imply that the complement equations of the main equations either hold trivially as a consequence of the main equations or they must be satisfied at a given value of $\lambda=\lambda_{\mathcal{B}}=const$. The former is called trivial equation and the latter are the so-called supplementary equations. For the set (II) - (20), the following assignations are made:
  • main equations
    $E_{\lambda\lambda}\;,\;\;q^{AB}E_{AB}\;,\;\;E_{\lambda}\;,\;\;E\;,$ (25)
    which given by (16), (17), (19) and (20), respectively.
  • supplementary equations :
    $E_{ww}\;,\;\;E_{w}$ (26)
    given by (II) and (18), respectively
  • trivial equation :
    $E_{w\lambda}$ (27)
    represented by (LABEL:eq:E01).
The main equations (25) do not form a typical hierarchy for the field equations like in other characteristic formulations, see e.g. Bondi et al. (1962); Newman and Penrose (1962); Mädler and Winicour (2016). This is due to the appearence of the $r_{,w}$ derivative in (17).
Definition of the following new fields
$\rho$ $=$ $r_{,w}\;\;,\;\;$ (28a)
$Z$ $=$ $Vrr_{,\lambda}+2\epsilon r\rho-\lambda\;\;,\;\;$ (28b)
$\mathcal{L}$ $=$ $\frac{2r\left(r_{,\lambda}\Phi_{,w}-\rho\Phi_{,\lambda}\right)+\epsilon(Z+\lambda)\Phi_{,\lambda}}{r_{,\lambda}}$ (28c)
cast (25) into
00 0 $=$ $r_{,\lambda\lambda}+\frac{1}{2}\kappa r|\Phi_{,\lambda}|^{2}\;,$ (29a)
$Q_{,\lambda}$ $=$ $-2\pi r^{2}q\mathcal{J}\;,$ (29b)
$\alpha_{,\lambda}$ $=$ $-\epsilon\frac{Q}{r^{2}}\;,$ (29c)
$Z_{,\lambda}$ $=$ $-\frac{\kappa}{8\pi}\frac{Q^{2}}{r^{2}}\;,$ (29d)
$\mathcal{L}_{,\lambda}$ $=$ $-\frac{\epsilon(\lambda+Z)\Phi_{,\lambda}}{r}-2iq\alpha(r\Phi)_{,\lambda}+i\frac{\epsilon q\Phi Q}{r}\;,$ (29e)
$V_{,\lambda\lambda}$ $=$ $-\frac{1}{\lambda}\left(\frac{\lambda^{2}}{r^{2}}\right)_{,\lambda}+\frac{2Zr_{,\lambda}}{r^{3}}+\frac{\kappa}{2\pi}\frac{Q^{2}}{r^{4}}$ (29f)
$-\frac{\kappa\epsilon}{2r}\left(\overline{\Phi_{,\lambda}}\mathcal{L}+\overline{\mathcal{L}}\Phi_{,\lambda}\right)+\kappa q\epsilon\alpha\mathcal{J}\;,$
$\Phi_{,w}$ $=$ $\frac{\mathcal{L}}{2r}-\frac{\epsilon}{2}V\Phi_{,\lambda}\;,$ (29g)
where we introduced the auxillary variable
$\mathcal{J}=i(\overline{\Phi}\Phi_{,\lambda}-\overline{\Phi}_{,\lambda}\Phi).$ (29h)
This system is completed with the two supplementary equations (II) and (18) as well as (17) evaluated at the boundary $\lambda=\lambda_{\mathcal{B}}=const$, which can be a world line Madler and Muller (2013), a finite size world tube Winicour (2011) or a null hypersurface Mädler (2019).
A solution of (29) requires initial data
$\Phi(0,\lambda)=\mathcal{F}_{\Phi}(\lambda)$ (30)
on a null hypersurface $w=0$ for values $\lambda\geq\lambda_{\mathcal{B}}$, where $\mathcal{F}_{\Phi}(\lambda)$ is an arbitrary function. It further requires the following boundary conditions at $\lambda=\lambda_{\mathcal{B}}$ for all values $w>0$
$\begin{split}&r(w,\lambda_{\mathcal{B}})\;,\;\;r_{,\lambda}(w,\lambda_{\mathcal{B}})\;,\;\;Q(w,\lambda_{\mathcal{B}})\;,\;\;\alpha(w,\lambda_{\mathcal{B}})\;,\;\;\\ &\mathcal{L}(w,\lambda_{\mathcal{B}})\;,\;\;V(w,\lambda_{\mathcal{B}})\;,\;\;V_{,\lambda}(w,\lambda_{\mathcal{B}})\;.\end{split}$ (31)
In particular, the boundary conditions for the variables $r$, $Q$, $\alpha$ and $V$ follow from physical considerations, while those for auxillary fields $Z$ and $\mathcal{L}$ they follow from the definitions (28) and are given by
$Z|_{\lambda=\lambda_{\mathcal{B}}}$ $=$ $(Vrr_{,\lambda}+2\epsilon rr_{,w}-\lambda)|_{\lambda=\lambda_{\mathcal{B}}}\;\;,\;\;$ (32a)
$\mathcal{L}|_{\lambda=\lambda_{\mathcal{B}}}$ $=$ $\left(2r\Phi_{,w}+\epsilon rV\Phi_{,\lambda}\right)|_{\lambda=\lambda_{\mathcal{B}}}\;\;.$ (32b)
It seems “unpleasant” that the boundary conditions (32) for $Z$ and $\mathcal{L}$, contain $w$ derivatives of the data $\Phi$ as well as the conformal factor $r$ at the boundary $\lambda=\lambda_{\mathcal{B}}$. We will see in the next sections that this can be (partially) resolved by specifying the nature of the boundary hypersurface at $\lambda_{\mathcal{B}}$.
Since we can always translate the origin of an affine parameter by a transformation like $\lambda\rightarrow a(w)\lambda+b(w)$, we set hereafter $\lambda_{\mathcal{B}}=0$. In the following three sections we discuss the asymptotic initial value problem, the initial value problem at a world line and the double null initital boundary value problem for two intersecting null hypersurfaces.
III.1 Asymptotic initial value problem for large values of the affine parameter
Let the scalar field have the following asymptotic expansion
$\Phi=\frac{\varphi_{[1]}+i\psi_{[1]}}{\lambda}+\frac{\varphi_{[2]}+i\psi_{[2]}}{\lambda^{2}}+O(\lambda^{-3})\;,$ (33)
where the functions with subscipt indices in brackets labelling the order of $\lambda^{-i}$ are real functions depending on $w$. The function $\mathcal{L}$ is also decomposed into a real and imaginary part
$\mathcal{L}=L+iP$ (34)
with $L$ and $P$ also being real functions. Insertion of (33) into (29h) yields
$\mathcal{J}=\frac{2(\varphi_{[1]}\psi_{[2]}-\varphi_{[2]}\psi_{[1]})}{\lambda^{4}}+O(\lambda^{-5})$ (35)
and subsequent integration of (29) gives the asymptotic solution
$r(w,\lambda)=$ $H\lambda+R_{\infty}-\left(\frac{\kappa H}{4}\right)\frac{(\varphi_{[1]}^{2}+\psi_{[1]}^{2})}{\lambda}+O(\lambda^{-2})\;,$ (36a)
$Q(w,\lambda)=$ $Q_{\infty}+\frac{4\pi qH^{2}(\varphi_{[1]}\psi_{[2]}-\varphi_{[2]}\psi_{[1]})}{\lambda}+O(\lambda^{-2})\;,$ (36b)
$\alpha(w,\lambda)=$ $\alpha_{\infty}+\left(\frac{\epsilon Q_{\infty}}{H^{2}}\right)\frac{1}{\lambda}+O(\lambda^{-2})\;,$ (36c)
$Z(w,\lambda)=$ $Z_{\infty}+\frac{\kappa Q^{2}_{\infty}}{8\pi H^{2}}\frac{1}{\lambda}+O(\lambda^{-2})\;,$ (36d)
$L(w,\lambda)=$ $L_{\infty}-\frac{\epsilon\varphi_{[1]}}{H\lambda}+q\frac{\epsilon Q_{\infty}\psi_{[1]}+2\alpha_{\infty}H(R_{\infty}\psi_{[1]}+H\psi_{[2]})}{H\lambda}+O(\lambda^{-2})\;,$ (36e)
$P(w,\lambda)=$ $P_{\infty}-\frac{\epsilon\psi_{[1]}}{H\lambda}-q\frac{\epsilon Q_{\infty}\varphi_{[1]}+2\alpha_{\infty}H(R_{\infty}\varphi_{[1]}+H\varphi_{[2]})}{H\lambda}+O(\lambda^{-2})\;,$ (36f)
$V(w,\lambda)=$ $U_{\infty}\lambda+V_{\infty}+\frac{2(HZ_{\infty}-R_{\infty})+\epsilon\kappa H^{2}(L_{\infty}\varphi_{[1]}+P_{\infty}\psi_{[1]})}{2H^{3}\lambda}+O(\lambda^{-2})\;,$ (36g)
where $H$, $R_{\infty}$, $Q_{\infty}$, $\alpha_{\infty}$, $Z_{\infty}$, $L_{\infty}$, $P_{\infty}$, $U_{\infty}$ and $V_{\infty}$ are functions of integrations depending on $w$. Using (17), we relate the functions integration $U_{\infty}$ and $V_{\infty}$
$U_{\infty}$ $=$ $-\frac{2\epsilon H^{\prime}}{H}\;,$ (37)
$V_{\infty}$ $=$ $\frac{1-2\epsilon HR_{\infty}^{\prime}}{H^{2}}\;\;.$ (38)
The evolution equations (29g) relate the constants $L_{\infty}$ and $P_{\infty}$ to the $w$ derivatives of $\varphi_{[1]}$ and $\psi_{[1]}$, respectively,
$L_{\infty}$ $=$ $2(H\varphi_{[1]})^{\prime}\;,$ (39)
$P_{\infty}$ $=$ $2(H\psi_{[1]})^{\prime}\;.$ (40)
The Bondi mass (22) can be found from the limit of the Misner-Sharp mass (21)
$m_{B}=\frac{1}{2}(R_{\infty}-HZ_{\infty})\;.$ (41)
From this result, alternative calculation of the Bondi mass using the hypersurface variables is
$m_{B}=-\frac{1}{2}\lim_{\lambda\rightarrow\infty}\left[\lambda^{2}\left(\frac{r}{\lambda}\right)_{,\lambda}+HZ\right]\;\;.$ (42)
The metric for large values of the affine parameter becomes
$\begin{split}ds^{2}=&-\left[-\frac{2H^{\prime}\lambda}{H}+\frac{1-2\epsilon HR^{\prime}_{\infty}}{H}\right]dw^{2}+2\epsilon dwd\lambda\\ &+(H\lambda)^{2}q_{AB}dx^{A}dx^{B}+O(\lambda^{-1})\;.\end{split}$ (43)
The asymptotic transformation
$x(w,\lambda)=H(w)\lambda+R_{\infty}(w)+O(\lambda^{-1})$ (44)
with $w$ and the angles kept unaltered casts (43) into
$\begin{split}ds^{2}=&-\left(\frac{dw}{H}\right)^{2}+2\epsilon dx\left(\frac{dw}{H}\right)\\ &+x^{2}\left[1+O(x^{-2})\right]^{2}q_{AB}dx^{A}dx^{B}\;\;.\end{split}$ (45)
Then if we define $dw_{B}=dw/H$ or alternatively
$\frac{dw_{B}}{dw}=\frac{1}{H}\;,$ (46)
the metric (8) approaches
$ds^{2}=-(dw_{B})^{2}+2\epsilon dw_{B}dx+x^{2}q_{AB}dx^{A}dx^{B}+O(x^{-1})\;\;,$ (47)
which is the metric of an inertial observer for large values of the affine parameter. The coordinate frame in which $H=1$ is called Bondi frame and $w_{B}$ referred to as Bondi time corresponding to the inertial time of this observer. Consequently, in a regime where $H\neq 0$, it measures the redshift between asymptotic inertial observers and observes at finite affine parameter distance. In particular, if $H\neq 0$, we can always find a coordinate transformation of the type (46) such that $H=1$. If, however, $H=0$ for some time $w=w_{E}$, then $w_{E}$ corresponds to the time of formation of an event horizon, since the surface forming rays of the surface $dw=const$ do not have an endpoint in the asymptotic region. The redshift can also conveniently be calculated via
$H=\lim_{\lambda\rightarrow\infty}r_{,\lambda}$ (48)
In a Bondi frame, the metric fields behave as
$\begin{split}r&=\lambda+R_{\infty}(w_{B})+O(\lambda^{-1})\\ V&=1-2\epsilon R_{\infty}^{\prime}(w_{B})+O(\lambda^{-2})\\ \end{split}$ (49)
and the scalar field $\Phi$ has the expansion,
$\Phi=\frac{\mathcal{C}(w_{B})}{r}+\frac{\mathcal{C}_{NP}}{r^{2}}+O(r^{-3})$ (50)
where $\mathcal{C}(w_{B})=\mathcal{C}_{\varphi}(w_{B})+i\mathcal{C}_{\psi}(w_{B})$ is the scalar monopole and $\mathcal{C}_{NP}$ is the conserved Newman-Penrose constant. The two can alternatively be determined via the limits
$\mathcal{C}$ $=$ $\lim_{r\rightarrow\infty}r\Phi=\lim_{\lambda\rightarrow\infty}H\lambda\Phi=H(\varphi_{[1]}+i\psi_{[1]})$ (51)
$\mathcal{C}_{NP}$ $=$ $-\!\lim_{r\rightarrow\infty}r^{2}(r\Phi)_{,r}=-\!\lim_{\lambda\rightarrow\infty}\frac{r^{2}(r\Phi)_{,\lambda}}{r_{,\lambda}}$ (52)
$=$ $R_{\infty}\mathcal{C}+H^{2}(\varphi_{[2]}+i\psi_{[2]})$
The derivative
$\mathscr{N}(w_{B}):=\frac{d\mathcal{C}}{dw_{B}}\;\;$ (53)
is the news function of the scalar field, whose naming will become justified below. The relation between the Bondi time $w_{B}$, and the asymptotic time $w$ is given by asymptotic transformation (46) therefore
$\mathcal{N}(w_{B}):=\frac{d\mathcal{C}}{dw}\frac{dw}{dw_{B}}=H\frac{d\mathcal{C}}{dw}$ (54)
The asymptotic limit of the Weyl scalar $\Psi_{2}$ gives
$\begin{split}\lim_{\lambda\rightarrow\infty}\lambda^{3}\Psi_{2}=&\frac{2HZ_{\infty}-R_{\infty}}{2H^{3}}\\ &+\frac{\kappa\epsilon}{6H^{2}}\left[H\varphi_{[1]}\frac{d}{dw}H\varphi_{[1]}+H\psi_{[1]}\frac{d}{dw}H\psi_{[1]}\right]\end{split}$ (55)
in a Bondi frame ($H=1$, $r=\lambda$), we can deduce that
$\begin{split}\lim_{r\rightarrow\infty}\Psi_{2}=&-m_{B}(w_{B})\\ &+\frac{1}{12}\kappa\epsilon\left[\mathcal{C}(w_{B})\bar{\mathcal{N}}(w_{B})+\bar{\mathcal{C}}(w_{B})\mathcal{N}(w_{B})\right]\;,\end{split}$ (56)
which gives an alternative means to calculate the Bondi mass in a Bondi frame
$m_{B}=-\lim_{r\rightarrow\infty}\left\{r^{3}\Psi_{2}+\frac{1}{12}\kappa\epsilon r^{2}\left[\Phi\partial_{w}\bar{\Phi}+\bar{\Phi}\partial_{w}\Phi\right]\right\}.$ (57)
The expression (57) corresponds to an expression presented in Bicak et al. (2010). Evaluation of the supplementary equations (18) and (II) in a Bondi frame ($H=1$) gives us the asymptotic balance laws
$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}\left\{\left(\frac{d\varphi_{[1]}}{dw_{B}}\right)^{2}+\left(\frac{d\psi_{[1]}}{dw_{B}}\right)^{2}+2q\alpha_{\infty}\left[\phi_{[1]}\left(\frac{d\psi_{[1]}}{dw_{B}}\right)-\psi_{[1]}\left(\frac{d\phi_{[1]}}{dw_{B}}\right)\right]+q^{2}\alpha_{\infty}^{2}\left(\phi_{[1]}^{2}+\psi_{[1]}^{2}\right)\right\}$ (58a)
$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-4\pi q^{2}\alpha_{\infty}^{2}\left(\phi_{[1]}^{2}+\psi_{[1]}^{2}\right)+4\pi q\left[\phi_{[1]}\left(\frac{d\psi_{[1]}}{dw_{B}}\right)-\psi_{[1]}\left(\frac{d\phi_{[1]}}{dw_{B}}\right)\right]$ (58b)
which we can also write a in terms of the complex news function and complex scalar monopole
$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}\left|\mathcal{N}+iq\alpha_{\infty}\mathcal{C}\right|^{2}$ (59a)
$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-4\pi q^{2}\alpha_{\infty}^{2}|\mathcal{C}|^{2}-2i\pi q\left(\mathcal{N}\bar{\mathcal{C}}-\bar{\mathcal{N}}\mathcal{C}\right)$ (59b)
As mentioned Sec. II, there is a gauge transformation to set [^3] Since $\alpha_{\infty}=\alpha_{\infty}(w_{B})$ this gauge transformation needs to be done on each null hypersurface $w_{B}=const$ for every value of $w_{B}$. $\alpha_{\infty}=0$. Using $\alpha_{\infty}=0$ in (59) casts the the Bondi mass conservation law and asymptotic the charge conservation law into the simple expressions
$\frac{dm_{B}}{dw_{B}}=\frac{\epsilon\kappa}{2}|\mathcal{N}|^{2}\;\;,\;\;\frac{dQ_{\infty}}{dw_{B}}=-2i\pi q\left(\mathcal{N}\bar{\mathcal{C}}-\bar{\mathcal{N}}\mathcal{C}\right)\;\;$ (60)
From the definition of the news function $\mathcal{N}$ in (53) as the derivative of the scalar monopole $\mathcal{C}$ and (60), we deduce that the asymptotic initial value problem for the charged version ($q\neq 0$ ) requires specification of three scalar values and one arbitrary function depending on $w_{B}$ while the uncharged version $(q=0)$ only requires specification of two scalars and one function depending on $w_{B}$. In both of these cases, the free function is the news function $\mathcal{N}$. Its name arises from the same motivation as in the gravitational case, see (Bondi et al., 1962, Sec. 3), the function $\mathcal{N}$ determines the complete (temporal) behaviour of all the other physically measurable fields, $\mathcal{C}(w_{B})$, $m_{B}(w_{B})$ and $Q_{\infty}(w_{B})$.
Both of the two versions of these asymptotic initial value problems can be discussed together. Given a scalar news function $\mathcal{N}$ for $w_{B}\geq 0$. The three scalars are specified at the initial time $w_{B}=0$. They are $\mathcal{C}(0)$, $m_{B}(0)$ and $Q_{\infty}(0)$, which are the initial scalar monopole, the initial Bondi mass and the initial value of the global charge, respectively. The behavior of these fields for $w_{B}>0$ is then determined by the following hierarchy of ordinary differential equations of the asymptotic initial value problem
$\frac{d\mathcal{C}}{dw_{B}}$ $=$ $\mathcal{N}(w_{B})\;\;,$ (61a)
$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}|\mathcal{N}(w_{B})|^{2}\;\;,$ (61b)
$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-2i\pi q\left[\mathcal{N}(w_{B})\bar{\mathcal{C}}(w_{B})-\bar{\mathcal{N}}(w_{B})\mathcal{C}(w_{B})\right]\;\;.$ (61c)
In the uncharged version, where the gauge charge $q$ vanishes, the last equation of (61) implies $Q_{\infty}(w_{B})=0$ as the initial value of the global charge $Q_{\infty}(0)$ is obviously zero. Thus only two initial values and one free function is required for the asymptotic initial value problem in this case.
III.2 Local initial-boundary formulation at a world line
In spherical symmetry there exists a designated worldline, which is the geodesic tracing the centers of spherical symmetry Synge (1960). We denote this geodesic with $c(w)$ and choose $w$ as an affine parameter for its parameterisation. The tangent vector $c^{a}\partial_{a}$ is a unit timelike vector and $w$ measures the proper time along $c(w)$. This proper time is assumed to be the proper time of a freely falling Fermi observer along $c(w)$ Manasse and Misner (1963); Poisson (2009). Following Madler and Muller (2013), in a convex normal neighborhood of $c(w)$ we extend the definition of $w$ be constant along null cones emanating from $c(w)$. The surface forming rays of the $w=const$ null cones are labeled by the angular coordinates $x^{A}$ within each null cone. The null condition on $w=const$ implies $g^{ww}=0$, since $g^{ab}\nabla_{a}w\nabla_{b}w=0$. The co-vector $\ell_{a}=\epsilon\nabla_{a}w$ defines by metric duality the tangent vectors $\ell^{a}=\epsilon g^{ab}\nabla_{b}w=\epsilon g^{aw}$ of the surface forming null rays. The null rays are labeled by the angular coordinates, $\ell^{a}\nabla_{a}x^{A}=0$ and therefore $g^{wA}=0$.
By these coordinate conditions, only the coordinate $x^{1}=\lambda$ varies along the rays of a given null cone $w=const$. Since the cones emanate from $c(w)$, we take the origin $\lambda=0$ to be on $c(w)$. We normalize the parameter $\lambda$ along the rays such that it is an affine parameter with the normalization determined at $c(w)$. In particular, $c^{a}\nabla_{a}\lambda=\ell^{a}\nabla_{a}\lambda=\epsilon g^{w\lambda}=1$ with $|\epsilon|=1$ at $c(w)$ so that $\epsilon=-1$ designates that the respective null cone $w=const$ are outgoing null cones with null rays pointing into the future of $c(w)$ and $\epsilon=1$ are ingoing null cones pointing into the past of $c(w)$.
By these conditions, the metric near $c(w)$ has the form [^4] RefMadler and Muller (2013) uses a reversed nomenclature for the binary parameter $\epsilon$. Madler and Muller (2013)
$\begin{split}ds^{2}=&-\left[1+O(\lambda^{2})\right](dw)^{2}+2\epsilon dwd\lambda\\ &+\lambda^{2}q_{AB}dx^{A}dx^{B}+O(\lambda^{3}).\end{split}$ (62)
This implies the regularity and boundary conditions for $\lambda=0$
$\begin{split}&r(w,0)=0\;\;,\;\;r_{,\lambda}(w,0)=1\;\;,\;\;\\ &V(w,0)=1\;\;,\;\;V_{,\lambda}(w,0)=0\;\;,\;\;\end{split}$ (63)
which also imply
$Z(w,0)=\mathcal{L}(w,0)=0\;\;,\;\;$ (64)
The supplementary equation (II) holds at $c(w)$ due to the regularity conditions. Eq. (63) and (18) imply
$Q_{,w}=O(\lambda^{2})\;\;\Rightarrow\;\;Q(w,0)=Q_{,\lambda}(w,0)=0$ (65)
At the vertex, we assume that the scalar field $\Phi$ has the regular Taylor expansion
$\Phi=\Phi_{(0)}+\Phi_{(1)}\lambda+\frac{1}{2}\Phi_{(2)}\lambda^{2}+O(\lambda^{3})\;\;,$ (66)
where the subscript $(k)$ in parenthesis indicate that the symbol carrying the subindex is the $\lambda^{k}$ coefficient function depending on $w$ of the Taylor series of the respective field evaluated at $\lambda=0$.
The scalar field expansion (66) and (29h) give
$\begin{split}\mathcal{J}=&\mathcal{J}_{(0)}+\mathcal{J}_{(1)}\lambda+O(\lambda^{2})\\ =&i\left(\overline{\Phi}_{(1)}\Phi_{(0)}-\overline{\Phi}_{(0)}\Phi_{(1)}\right)+i\left(\overline{\Phi}_{(2)}\Phi_{(0)}-\overline{\Phi}_{(0)}\Phi_{(2)}\right)\lambda\\ &+O(\lambda^{2})\;.\end{split}$ (67)
Insertion of these data into (29) and integration gives
$r(w,\lambda)$ $=$ $\lambda-\frac{\kappa\lambda^{3}}{12}|\Phi_{(1)}|^{2}-\frac{\kappa\lambda^{4}}{24}\Bigl{(}\Phi_{(2)}\overline{\Phi}_{(1)}+\Phi_{(1)}\overline{\Phi}_{(2)}\Bigr{)}+O(\lambda^{5})\;,$ (68a)
$Q(w,\lambda)$ $=$ $-\pi q\left(\frac{2}{3}\mathcal{J}_{(0)}+\frac{1}{2}\mathcal{J}_{(1)}\lambda\right)\lambda^{3}+O(\lambda^{5})\;,$ (68b)
$\alpha(w,\lambda)$ $=$ $\alpha_{0}+\frac{\pi\epsilon q}{3}\left(\mathcal{J}_{(0)}+\frac{1}{2}\mathcal{J}_{(1)}\lambda\right)\lambda^{2}+O(\lambda^{4})\;,$ (68c)
$Z(w,\lambda)$ $=$ $-\frac{\pi\kappa q^{2}\mathcal{J}_{(0)}}{90}\Big{(}\mathcal{J}_{(0)}+\frac{5}{4}\mathcal{J}_{(1)}\lambda\Big{)}\lambda^{5}+O(\lambda^{7})\;,$ (68d)
$\mathcal{L}(w,\lambda)$ $=$ $-\left(\epsilon\Phi_{(1)}+2iq\alpha_{0}\Phi_{(0)}\right)\lambda-\left(\frac{\epsilon\Phi_{(2)}}{2}+2iq\alpha_{0}\Phi_{(1)}\right)\lambda^{2}+O(\lambda^{3})\;,$ (68e)
$V(w,\lambda)$ $=$ $1+\frac{\kappa}{3}|\Phi_{(1)}|^{2}\lambda^{2}+O(\lambda^{3})\;,$ (68f)
$\Phi_{,w}$ $=$ $-\epsilon\Phi_{(1)}-iq\alpha_{0}\Phi_{(0)}-\left[\frac{3}{4}\epsilon\Phi_{(2)}+iq\alpha_{0}\Phi_{(1)}\right]\lambda+O(\lambda^{2})\;.$ (68g)
The gauge freedom of the electromagnetic field allows us in principle to set $\alpha_{0}$ everywhere along $c(w)$. Moreover, from (69) it is not difficult to deduce that
$\Phi_{(0),w}=$ $-\epsilon\Phi_{(1)}-iq\alpha_{0}\Phi_{(0)}\;,$ (69)
$\quad\Phi_{(1),w}=$ $-\frac{3}{4}\epsilon\Phi_{(2)}-iq\alpha_{0}\Phi_{(1)}\;.$ (70)
We can see that the time evolution of the coefficient $\Phi_{(i)}$ along $c(w)$ is driven by a function depending on the expansion coefficients containing all expansion coefficients up to $\Phi_{(i+1)}$. This is behavior is a result of the choice of $w$ being constant along a null hypersurface. The hierarchical structure of (69) is the scalar field equivalent to the similar structure in the gravitational case Madler and Muller (2013); Mädler (2013) for gravitational waves. Since $\Phi_{(1)}=\lim_{\lambda\rightarrow 0}\Phi_{,\lambda}$ and $\Phi_{(2)}=\lim_{\lambda\rightarrow 0}\Phi_{,\lambda\lambda}$, we learn from (68) that the evolution of the initial data $\Phi$ along $c(w)$ and in its neighbourhood is completely determined by the data on a cone $w=const$, provided the data and its derivatives are finite in the limit $\lambda\rightarrow 0$. Indeed it is not difficult to derive similar equations like in (69) for $\Phi_{(k)}$ with $k\geq 2$ while using an appropriate higher order expansion than (66) and integrating (29).
From such an exercise, we can deduce the following local characteristic initial-boundary value problem at the vertices of null cones emanating from a geodesic world line: Suppose, the complex scalar field $\Phi(w,\lambda)$ on a null cone $w=0$ has a regular Taylor series expansion,
$\Phi(0,\lambda)=\sum^{\infty}_{k=0}\frac{\lambda^{k}}{k!}\frac{\partial^{k}\Phi}{\partial\lambda^{k}}\Bigg{|}_{\lambda=0}\;,$ (71)
evaluated at the world line, the boundary conditions of the fields are (63) and (64), $\alpha_{0}$ is an arbitrary function of $w$ for $w>0$, and the gauge charge $q$ is provided, then the metric, the electromagnetic potential $\alpha$ and the charge function $Q$ can be completely calculated on the cone $w=0$ using the hypersurface equations (29a)-(29f). Their local solution at $\lambda=0$ is of the form (68a), (68b), (68c) or (68f). The evolution equations (29g) determine the time derivative of the Taylor coefficients of the initial data, as seen in e.g. (69).
We remark, that this regularity assumption at $w=0$ does not imply that the data necessarily remain regular for $w>0$. If $r\rightarrow 0$ for $\lambda\rightarrow\lambda_{C}>0$ a (massive) caustic forms at finite value for the affine parameter. This caustic formation for $w>0$ is in general driven by the value of norm squared, $|\Phi_{,\lambda}|^{2}$, of the derivative of the scalar field. An example for caustic initical data for a real scalar field was recently given in Mädler et al. (2024), where the data were employed in an evolution through an event horizon of a supercritical scalar field. The thus formulated initial value problem in spherical symmetry and with a scalar field is an analogue to the gravitational one of Friedrich Friedrich (1986).
III.3 Initial-boundary value problem at a null hypersurface
Consider a family of null hypersurfaces $w=const$ designated by $\mathscr{N}^{\epsilon}(w)$, hereafter. The generators $\ell^{a}\partial_{a}$ of these null hypersurface are affinely parameterized with the affine parameter $\lambda$. We choose the origin, $\lambda=0$, of the affine parameter to be at a null hypersurface $\mathcal{B}^{\epsilon}$. The surface $\mathcal{B}^{\epsilon}$ is a null boundary for every null hypersurface $w=const$ of $\mathscr{N}^{\epsilon}$ and is the boundary surface for the double null CIBVPs to be presented, see Fig.(1).
Figure 1: A family of null hypersurfaces $\mathscr{N}^{\epsilon}(w)$ with corresponding boundary null hypersurface $\mathcal{B}^{\epsilon}(w)$ is shown for $\epsilon=-1$ (left) and $\epsilon=-1$ (right). The former is a characteristic initial value problem with the initial data surface $\mathscr{N}^{-}(0)$ and the latter gives rise to a characteristic past value problem with the past data on $\mathscr{N}^{+}(0)$. The common intersection between $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}(w)$ is $\Sigma(0)$.
Near $\mathcal{B}^{\epsilon}$ on $\mathscr{N}^{\epsilon}$, the metric must behave like
$ds^{2}=O(\lambda)(dw)^{2}+2\epsilon dwd\lambda+r^{2}(w,0)q_{AB}dx^{A}dx^{B}+O(\lambda^{3}).$ (72)
This implies $V(w,0)=0$ for all values of $w$ and where $\lambda=0$. Consequently, $V=O(\lambda)$ near $\mathcal{B}^{\epsilon}$ on $\mathscr{N}^{\epsilon}$. For the conformal factor $r$ there are two possibilities regarding its evaluation on $\mathcal{B}^{\epsilon}$: i) $r(w,0)=0$ and ii) $r(w,0)\neq 0$.
Physically these cases mean that the null hypersurface $\mathcal{B}^{\epsilon}$ has either a trivial (case (i)) or non-trivial Misner-Sharp mass (case (ii)), which becomes clear considering (21) in the limit $\lambda\rightarrow\infty$
$\lim_{\lambda\rightarrow\infty}M=\frac{r(w,0)}{2}\left[1-2\epsilon r_{,w}(w,0)r_{,\lambda}(w,0)\right]\;\;.$ (73)
The two options lead to different implications for the formulation of an initial boundary value problem. These options are considered separately, hereafter.
III.3.1 Initial boundary formulation at a null hypersurface with $r(w,0)=0$
If $V(w,0)=r(w,0)=0$, the line element near the origin of the affine parameter $\lambda$ can be written like
$\begin{split}ds^{2}=&-\left[V_{(1)}(w)\lambda+O(\lambda^{2})\right](dw)^{2}+2\epsilon dwd\lambda\\ &+[r_{(1)}(w)\lambda]^{2}\left[1+O(\lambda)\right]q_{AB}dx^{A}dx^{B}\;,\end{split}$ (74)
where it is required that $r_{1}\neq 0$. Locally there exists a coordinate transformation, $\hat{\lambda}=r_{1}\lambda$ and $\hat{w}=\int(dw/r_{1})$, so that the line element takes the form
$\begin{split}ds^{2}=&-\left\{\left[r_{(1)}V_{(1)}+2\epsilon r_{(1),\hat{w}}\right]\hat{\lambda}+O(\hat{\lambda}^{2})\right\}(d\hat{w})^{2}\\ &+2\epsilon d\hat{w}d\hat{\lambda}+\hat{\lambda}^{2}\left[1+O(\hat{\lambda})\right]q_{AB}dx^{A}dx^{B}.\end{split}$ (75)
Consequently, we can set wlog $r_{(1)}=1$ which is assumed herefter in this subsection. With the assumption of a regular expansion of the scalar field like in (66), we find that $Z(w,0)=\mathcal{L}(w,0)=0$. The residual gauge freedom of the electromagnetic field $\alpha$, allows us to set $\alpha(w,0)=0$. In addition, the supplementary equation (18) implies $Q(w,0)=const$ for $\lambda=0$ and we may choose $Q(w,0)=0$. The metric supplementary equation (II), gives in the limit $\lambda\rightarrow 0$
$0=V_{(1),w}+\epsilon\left[-V_{(1)}^{2}+\kappa|\Phi_{(0),w}|^{2}\right]$ (76)
This equation can be solved along the null boundary $\mathcal{B}^{\epsilon}$ for $w>0$, provided an initial value $V_{(1)}(0)$ is given at $w=0$ and provided the function $\Phi_{(0),w}$ is known for all values $w\geq 0$. The function $V_{(1)}$ measures the inaffinity of the null geodesics $n^{a}\partial_{a}$ ruling the null hypersurface $\mathcal{B}^{\epsilon}$, which follows from (12). For a non-vanishing and $w$-dependent scalar field on $\mathcal{B}^{\epsilon}$, (76) shows that the evolution of the inaffinity along $\mathcal{B}^{\epsilon}$ is driven by the free function $\Phi_{(0),w}$. Thus, the function $\Phi_{(0),w}$ can be seen as some kind of analogue of the news function at $\lambda=0$ instead of in the asymptotic limit $\lambda\rightarrow\infty$.
The characteristic initial value problem can now be constructed in the following way. Let $w=0$ be a null hypersurface $\mathscr{N}^{\epsilon}(0)$ whose generators are tangent to affinly parameterized null geodesics with tangent $\ell^{a}\partial_{a}$ and affine parameter $\lambda$. At the origin of the affine parameter, $\lambda=0$, the null hypersurface $\mathscr{N}^{\epsilon}(0)$ intersects a null hypersurface $\mathcal{B}^{\epsilon}$ which has the generators $n^{a}\partial_{a}$ that obey the normalization $n^{a}\ell_{a}=-1$. The intersection of the two null hypersurfaces $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}$ is a two dimensional spacelike space with spherical topology and unit Gaussian curvature, called $\Sigma$. The intrinsic metric of $\Sigma$ is the one of the unit sphere, $q_{AB}$. For this double null characteristic initial value problem, the following specification of data determines a solution of the Einstein-scalar field equations in spherical symmetry:
  • specify the type of null hypersurface $w=0$, by choosing either $\epsilon=1$ (ingoing null hypersurface) or $\epsilon=-1$ (outgoing null hypersurface)
  • specify the gauge charge $q$
  • on $\mathscr{N}^{\epsilon}(0)$ specify a free complex function $\mathcal{F}_{\Phi}(\lambda)$ that can be expressed by a regular Taylor series at $\lambda=0$
  • on $\mathcal{B}^{\epsilon}$ specify a free complex function $\mathcal{N}_{\mathcal{B}}(w)$
  • on $\Sigma$ specify the following fields: $\Phi(0,0)$ and $V_{(1)}(0)$
  • on $\mathcal{B}^{\epsilon}$, i.e. for $\lambda=0$, the following fields are trivial: $r$, $V$, $Q$, $\alpha$, $Z$ and $\mathcal{L}$
  • on $\mathcal{B}^{\epsilon}$, the field $r_{,\lambda}$ is nontrivial, i. e. $\lim_{\lambda\rightarrow 0}r_{,\lambda}=1$ for all values $w\geq 0$
With these specifications the field equations are solved in the following way:
  1. 1.
    with $\mathcal{N}_{\mathcal{B}}(w)$ and $\Phi(0,0)$ determine the scalar field everywhere on $\mathcal{B}^{\epsilon}$ via the equation
    $\Phi_{,w}(w,0)=\mathcal{N}_{\mathcal{B}}(w).$ (77)
  2. 2.
    with $\mathcal{N}_{\mathcal{B}}(w)$ and $V_{(1)}(0)$ determine the inaffinity everywhere on $\mathcal{B}^{\epsilon}$ via the equation (which follows from (76))
    $V_{(1),w}=-\epsilon\left[V_{(1)}^{2}+\kappa|\mathcal{N}_{\mathcal{B}}|^{2}\right].$ (78)
  3. 3.
    with $\mathcal{F}_{\Phi}(\lambda)$ and $\Phi(0,0)$ determine the scalar field everywhere on $\mathscr{N}^{\epsilon}$ via the equation
    $\Phi_{,\lambda}(0,\lambda)=\mathcal{F}_{\Phi}(\lambda).$ (79)
  4. 4.
    as all boundary data for $r$, $Q$, $\alpha$, $Z$, $\mathcal{L}$ and $V$ are known on $\mathcal{B}^{\epsilon}$, the hypersurface equations (29a) - (29f) can be hierarchically integrated for all the fields relevant to determine the $w$-derivative of $\Phi$ in equation (29g)
  5. 5.
    integration of the $w$-derivative of $\Phi$ by equation (29g) (using for example a finite difference scheme) determines $\Phi(\Delta w,\lambda)$ at some time $w=\Delta w>0$ on a null hypersurface $\mathscr{N}^{\epsilon}(\Delta w)$
  6. 6.
    the new values $\Phi(\Delta w,\lambda)$ serve as initial data on $\mathscr{N}^{\epsilon}(\Delta w)$ and are used to determine the metric and electromagnetic fields following step 4.
This algorithm can be suitably implemented in a numerical algorithm.
III.3.2 Initial boundary formulation at a null hypersurface with $r(w,0)\neq 0$
If the areal distance function is non-trivial on the bounday $\mathcal{B}^{\epsilon}$, i.e. $r(w,0)\neq 0$, a different boundary value formulation will come about. Near the boundary the metric has the behaviour
$\begin{split}ds^{2}=&-\left[V_{(1)}(w)\lambda+O(\lambda^{2})\right]dw^{2}+2\epsilon dwd\lambda\\ &+\left[r_{(0)}(w)+r_{(1)}(w)\lambda+O(\lambda^{2})\right]^{2}q_{AB}dx^{A}dx^{B}\;\;,\end{split}$ (80)
with $r_{(0)}(w)=r(w,0)\neq 0$. Suppose that there is a coordinate transform like
$w=\hat{w}+G(\hat{w})\hat{\lambda}+O(\hat{\lambda}^{2})\;\;,\;\;\lambda=(\hat{\lambda}-r_{(0)})/r_{(1)}+O(\hat{\lambda}^{2})\;\;.$ (81)
that casts (80) into a form similar to the expansion of (74). If that were the case, the initial boundary value problem could be discussed like in the previous section. It is indeed possible to find a coordinate transformation for (80) such that the resulting metric has $g_{AB}=\hat{\lambda}^{2}q_{AB}+O(\hat{\lambda}^{3})$, but such a transformation gives $g_{\hat{w}\hat{w}}|_{\hat{\lambda}=0}=O(1)\neq 0$ in the new coordinates. This violates the null condition at the boundary $\hat{\lambda}=0$ and cannot be removed without requiring $r_{(0),w}=0$. Since the coefficient $r_{(0)}$ corresponds to a mass, which becomes clear considering (73), we discard the restriction $r_{(0),w}=0$ as it implies that null boundary $\mathcal{B}^{\epsilon}$ has a mass which remains constant wrt a $w$ evolution. Thus, having $r_{(0)}\neq 0$ on $\mathcal{B}^{\epsilon}$ indicates that $\mathcal{B}^{\epsilon}$ is a null hypersurface with a given mass. The choice of (80) with $r_{(0)}>0$ corresponds to choosing a spherically symmetric Gaussian null coordinate system Moncrief and Isenberg (1983) at $\mathcal{B}^{\epsilon}$ since $g_{AB}$ is positive definite on $\mathcal{B}^{\epsilon}$. In the next section it will be shown that $\mathcal{B}^{\epsilon}$ with $r_{(0)}\neq 0$ can be related to a horizon of a black hole. Although there is no local coordinate transformation to cast (80) into a form like (74), it can still be simplified because $V_{(1)}$ can be set to zero everywhere on $\mathcal{B}^{\epsilon}$. The null vector generating $\mathcal{B}^{\epsilon}$ is $n^{a}\partial_{a}$ and the generating curves are non-affine null geodesics, which becomes clear when projecting (12) to $\mathcal{B}^{\epsilon}$ whose only component is
$n^{a}\nabla_{a}n^{w}=-\frac{1}{2}V_{(1)}n^{w}.$ (82)
The coordinate transformation $w=w(\hat{w})$ brings (83) into the form
$n^{a}\nabla_{a}n^{\hat{w}}=-\left(\frac{d}{d\hat{w}}\ln\left|\frac{dw}{d\hat{w}}\right|+\frac{1}{2}V_{(1)}\right)n^{\hat{w}}.$ (83)
If the coordinate transformation is chosen such that the term in the parenthesis of (83) vanishes, the non-affine null geodesics ruling $\mathcal{B}^{\epsilon}$ become affine null geodesics and $g_{\hat{w}\hat{w}}=O(\lambda^{2})$ near $\mathcal{B}^{\epsilon}$. This removal of the inaffinity is in fact a classical feature of non-affine geodesics Schouten (1954); Synge (1960); Poisson (2009) and not restricted to null geodesics. We make this choice in this section. In addition, we make the gauge choice $\alpha=0$ everwhere on $\mathcal{B}^{\epsilon}$. Evaluation of the supplementary equations (II) and (18) on $\mathcal{B}^{\epsilon}$ yields
00 0 $=$ $\left(r_{,ww}+\frac{1}{2}\kappa r|\Phi_{,w}|^{2}\right)\bigg{|}_{\mathcal{B}^{\epsilon}}\;\;,$ (84a)
00 0 $=$ $\left[Q_{,w}+2\pi qr^{2}i(\overline{\Phi}\Phi_{,w}-\Phi\overline{\Phi}_{,w})\right]\bigg{|}_{\mathcal{B}^{\epsilon}}\;\;.$ (84b)
The two equations (84a) and (84b) determine the value of $r$ and $Q$ everywhere on $\mathcal{B}^{\epsilon}$ provided the scalar field $\Phi$ and its $w$-derivative are known on $\mathcal{B}^{\epsilon}$. A solution of (29a) also requires the specification of $r_{,\lambda}$ on $\mathcal{B}^{\epsilon}$. The evaluation of the main equation $q^{AB}E_{AB}$, from (17), while using $V_{,\lambda}|_{\lambda=0}=0$ yields
00 0 $=$ $\left[2\epsilon(rr_{,\lambda})_{,w}-1+\frac{\kappa}{8\pi}\frac{Q^{2}}{r^{2}}\right]\bigg{|}_{\mathcal{B}^{\epsilon}}\;\;.$ (84c)
Equation (84c) can be used to determine $r_{,\lambda}$ everwhere on $\mathcal{B}^{\epsilon}$ provided $r$ and $Q$ are known on $\mathcal{B}^{\epsilon}$ and $r\neq 0$ on $\mathcal{B}^{\epsilon}$. The start values for solving the hypersurface equations (29d) and (29e) for $Z$ and $\mathcal{L}$ follow from evaluation of their definitions on $\mathcal{B}^{\epsilon}$
$Z\Big{|}_{\mathcal{B}^{\epsilon}}=2\epsilon(rr_{,w})\Big{|}_{\mathcal{B}^{\epsilon}}\;\;,\;\;\mathcal{L}\Big{|}_{\mathcal{B}^{\epsilon}}=2(r\Phi_{,w})\Big{|}_{\mathcal{B}^{\epsilon}}\;\;,$ (85)
which show that the initial values for $Z$ and $\mathcal{L}$ are known once $r$ and $\Phi_{,w}$ are known on $\mathcal{B}^{\epsilon}$. With the information collected above, a double null characteristic initial value problem can be formulated with a nontrivial conformal factor $r$ on the null boundary $\mathcal{B}^{\epsilon}$.
In a spherically symmetric spacetime, let $\mathscr{N}^{\epsilon}(w)$ be a family of null hypersurface $w=const$ with generators $\ell^{a}\partial_{a}$ that are affinely parameterized with an affine parameter $\lambda$. The affine parameter $\lambda$ is normalized with a parameter $\epsilon$ via $\ell^{a}\nabla_{a}\lambda=\epsilon$ with $\epsilon^{2}=1$. The binary parameter $\epsilon$ indicates whether a null hypersurface $w=const$ is an outgoing null hypersurface ($\epsilon=-1$) or an ingoing null hypersurface ($\epsilon=1$). The null hypersurfaces $\mathscr{N}^{\epsilon}(w)$ intersect a null hypersurface $\mathcal{B}^{\epsilon}(w)$, where $\lambda=0$ on $\mathscr{N}^{\epsilon}(w)$. The generators of $\mathcal{B}^{\epsilon}(w)$ are null vectors $n^{a}\partial_{a}$. These generators obey the cross normalisation $\ell^{a}n_{a}=-1$ and the integral curves (null geodesics) of $n^{a}$ are affinely parameterized with the parameter $w$ on $\mathcal{B}^{\epsilon}(w)$. The intersections of $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}(w)$ are spatial spherical cross sections $\Sigma(w)$ of co-dimension 2 with an intrinsic two metric $g_{AB}$. The volume element $\sqrt{\det(g_{AB})}=r^{2}(w,0)\sqrt{q}$ with $\sqrt{q}$ being the volume element of a unit sphere metric, $q_{AB}$, is nowhere zero on $\mathcal{B}^{\epsilon}(w)$ and for $\lambda=0$ on $\mathscr{N}^{\epsilon}(w)$. The function $r(w,0)$ is extended for values $\lambda>0$ on every null hypersurface $w=const$ to be be an areal distance so that the volume integral for a value $\lambda_{c}>0$ is given by $4\pi r^{2}(w,\lambda_{c})$. A solution of the Einstein-Maxwell-scalar field equations on the domain $\mathcal{B}^{\epsilon}(w)\cup\mathscr{N}^{\epsilon}(w)$ can be found for the double-null CIBVP with the following specifications:
  • provide a value for $\epsilon$ to specify the nature of the initial data surface $\mathscr{N}^{\epsilon}(0)$ with $w=0$
  • specify the value of the gauge charge $q$
  • on $\mathscr{N}^{\epsilon}(0)$ specify a free (sufficiently differentiable) complex function $F_{\Phi}(\lambda)$
  • on $\mathcal{B}^{\epsilon}(0)$ specify a free (sufficiently differentiable) complex function $\mathcal{N}_{\mathcal{B}}(w)$
  • the following fields are trivial everywhere on $\mathcal{B}^{\epsilon}(w)$
    $\alpha(w,0)\;\;,\;\;V(w,0)\;\;,\;\;V_{,\lambda}(w,0).$ (86)
  • on the common intersection $\Sigma(0)=\mathcal{B}^{\epsilon}(0)\cap\mathscr{N}^{\epsilon}(0)$ specify the values for the following fields
    $\begin{split}&\Phi(0,0)\;,\;\;r(0,0)\;,\;\;r_{,w}(0,0)\;,\;\;\\ &r_{,\lambda}(0,0)\;,\;\;Q(0,0).\end{split}$ (87)
    where in particular $r(0,0)>0$.
This (double null) CIBVP can be solved in the following manner
  1. 1.
    with $\mathcal{N}_{\mathcal{B}}(w)$ and $\Phi(0,0)$ determine the scalar field everywhere on $\mathcal{B}^{\epsilon}(w)$ via the equation
    $\Phi_{,w}(w,0)=\mathcal{N}_{\mathcal{B}}(w)$ (88)
  2. 2.
    with the boundary data $\mathcal{N}_{\mathcal{B}}(w)$, the scalar field $\Phi(w,0)$ on $\mathcal{B}^{\epsilon}(w)$ and the initial values $r(0,0)$, $r_{,w}(0,0)$, $Q(0,0)$ and $r_{\lambda}(0,0)$ solve the hierarchical set of boundary equations
    00 0 $=$ $r_{,ww}(w,0)+\frac{1}{2}\kappa r(w,0)|\mathcal{N}_{\mathcal{B}}|^{2}\;\;,$ (89)
    00 0 $=$ $Q_{,w}(w,0)+2i\pi qr^{2}(w,0)\left[\overline{\Phi}(w,0)\mathcal{N}_{\mathcal{B}}-\Phi(w,0)\overline{\mathcal{N}}_{\Phi}\right]\bigg{|}_{\mathcal{B}^{\epsilon}}\;\;.$ (90)
    00 0 $=$ $\left[2\epsilon(rr_{,\lambda})_{,w}-1+\frac{\kappa}{8\pi}\frac{Q^{2}}{r^{2}}\right]\bigg{|}_{\mathcal{B}^{\epsilon}}.$ (91)
    to obtain the fields $r$, $Q$ and $r_{,\lambda}$ on $\mathcal{B}^{\epsilon}(w)$. With the values of $r$, $r_{,w}$ and $\mathcal{N}_{\mathcal{B}}(w)$ on $\mathcal{B}^{\epsilon}$ determine the fields $Z$ and $\mathcal{L}$ everywhere on $\mathcal{B}^{\epsilon}(w)$. This determines all necessary boundary values to integrate the hypersurface equations (29a)-(29f).
  3. 3.
    on the initial null hypersurface $w=0$ utilize the initial value $\Phi(0,0)$ on $\Sigma(0)$ and the initial data $\mathcal{F}_{\Phi}$ to integrate
    $\Phi_{,\lambda}(0,\lambda)=\mathcal{F}_{\Phi}(\lambda)$ (92)
    for the initial data $\Phi(0,\lambda)$.
  4. 4.
    on $\mathscr{N}^{\epsilon}(0)$, i.e. the null hypersurface $w=0$, integrate hierarchically the hypersurface equations (29a)-(29f) to obtain the fields $r(0,\lambda)$, $Q(0,\lambda)$, $\alpha(0,\lambda)$, $Z(0,\lambda)$, $\mathcal{L}(0,\lambda)$ and $V(0,\lambda)$ while using $\mathcal{F}_{\Phi}(\lambda)$, $\Phi(0,\lambda)$ and the respective boundary value fields evaluated at $w=0$.
  5. 5.
    with $r(0,\lambda)$, $\mathcal{L}(0,\lambda)$, $V(0,\lambda)$ and $\Phi_{,\lambda}(0,\lambda)$ algebraically calculate $\Phi_{,w}(0,\lambda)$ on $w=0$ using equation (29g), integrate this $w$ derivative (using for example a finite difference scheme) to find $\Phi(\Delta w,\lambda)$ at some time $w=\Delta w>0$ on a null hypersurface $\mathscr{N}^{\epsilon}(\Delta w)$
  6. 6.
    the new values $\Phi(\Delta w,\lambda)$ serve as initial data on $\mathscr{N}^{\epsilon}(\Delta w)$ and are used to determine the metric and electromagnetic fields following step 4 above but using the coordinate time $w=\Delta w$ instead of $w=0$.
Comparing this formulation of the double null CIBVP with the one of the previous section Sec. III.3.1, reveals that the two are very similar. In fact, the only essential difference is step 2 - the determination of the nontrivial boundary fields in $\mathcal{B}^{\epsilon}(w)$. The reason for this difference is that for one of the formulations $r=0$ on $\mathcal{B}^{\epsilon}(w)$ while for the other one it is essentially non-trivial. The two different formulations give rise for different areas of application. The first scheme, where $r=0$ on $\mathcal{B}^{\epsilon}(w)$, is well suited for studying the evolution of initial data with a regular center. The second scheme is suited for the study of (charged) black holes in which $\mathcal{B}^{\epsilon}(w)$ is a horizon with an associated horizon mass of the black hole. Indeed, in Sec. V.1 the second scheme is used to derive the Reisner-Nordström solution of a charged black hole.
IV Symmetry considerations
For the subsequent sections, we consider possible timelike Killing vectors $\xi$. Killing equations are given by
$0=\nabla^{a}\xi^{b}+\nabla^{b}\xi^{a}=-g^{ab}_{\;\;,c}\xi^{c}+g^{cb}\xi^{a}_{\;\;,c}+g^{ac}\xi^{b}_{\;\;,c}\;.$ (93)
An ansatz for the Killing vector $\xi$ left invariant under spatial rotations is $\xi=\xi^{w}\partial_{w}+\xi^{\lambda}\partial_{\lambda}$ with the components $\xi^{w}$ and $\xi^{\lambda}$ only dependent on $w$ and $\lambda$. Using (9) we find
$0=$ $2\epsilon\xi^{w}_{\;,\lambda}\;,$ (94a)
$0=$ $\epsilon\xi^{w}_{\;,w}+\epsilon\xi^{\lambda}_{\;,\lambda}+V\xi^{w}_{,\lambda}\;,$ (94b)
$0=$ $-\frac{2}{r^{3}}\left(r_{,w}\xi^{w}+r_{,\lambda}\xi^{\lambda}\right)\;,$ (94c)
$0=$ $-V_{,w}\xi^{w}-V_{,\lambda}\xi^{\lambda}+2\epsilon\xi^{\lambda}_{\;,w}+2V\xi^{\lambda}_{\;,\lambda}\;.$ (94d)
From the first two, we find
$\xi=A(w)\partial_{w}+\left[-A_{,w}(w)\lambda+B(w)\right]\partial_{\lambda}\;,$ (95)
where $A$ and $B$ are free functions. It is convenient to introduce the scalar function
[…]
which is invariant under the action of the vector field $\xi$ that is the Lie derivative $\mathscr{L}_{\xi}z=\xi^{a}\partial_{a}z=0$. In principle we may choose $B=0$, because of the metric’s invariance of the coordinate transformation of the type (10). However, we will keep it general to allow a broader family of null affine coordinates. Inserting (95) into (94c) yields
$0=\frac{2}{r^{3}}\left[r_{,w}A-(\lambda A_{,w}-B)r_{,\lambda}\right]\;,$ (97)
whose general solution can be found as
$r=\hat{r}(z)\;,$ (98)
where $\hat{r}(\cdot)$ is any non-vanishing differentiable function. Finally, inserting (95) into (94d) gives us
$\begin{split}0=&V_{,w}A-V_{,\lambda}(A_{,w}\lambda-B)-2\epsilon\left(A_{,ww}\lambda-B_{,w}\right)+2VA_{,w}\end{split}$ (99)
with the general solution
$V=\frac{\hat{V}[z(w,\lambda)]}{A^{2}(w)}-\frac{2\epsilon(A_{,w}\lambda-B)}{A}\;,$ (100)
in which $\hat{V}(\cdot)$ is an arbitrary differentible function.
Some choices for $A$ and $B$ relevant in this work are listed in Table  1.
[…]
Table 1: Selective choices for the $A$ and $B$ to build a Killing vector.
The coordinate dependence of the metric potentials $r$ and $V$ for the different expressions for Killing vectors in Table 1 is general. For further study of particular applications it is practical to study the norm of the Killing vector either for large values of the affine parameter or its origin. The former will give us restrictions to the metric by requiring the Killing vector be asymptotically timelike, while the latter gives us necessary conditions, if the origin of the affine parameter is on a Killing horizon with a given surface gravity.
The norm of the Killing vector (95) is
$\xi^{a}\xi_{a}=-VA^{2}-\epsilon(A^{2})_{,w}\lambda+2\epsilon AB$ (101)
and in its asymptotic limit ($\lambda\rightarrow\infty$) we find using (36g), (37) and (LABEL:eq:Vinf)
$\xi^{a}\xi_{a}=$ $-2\epsilon A^{2}\left[\left(\ln\frac{A}{H}\right)_{,w}\right]\lambda$
+$+\left\{2\epsilon B-\frac{[1-2\epsilon H(R_{\infty})_{,w}]A}{H^{2}}\right\}A+O(\lambda^{-1}).$ (102)
Hence, as seen in (IV), the asymptotic behavior of the Killing vector norm is determined by the asymptotic behavior of the areal distance through the functions $H$ and $R_{\infty}$. Therefore, in order to have a finite norm of this asymptotic limit, the linear term in $\lambda$ should vanish. One posibility is to set $A=0$, which from (101) implies a Killing vector field with vanishing norm of the form $\xi=B\partial_{\lambda}$. However, taking into account (98) we see that it implies that the vector field is zero everywhere or $\xi$ is a null Killing vector with $r=r(w)$ everywhere, which is inadmissible.
The other choice for the finiteness of the norm of the Killing vector for $\lambda\rightarrow\infty$ is
$\left(\ln\frac{A}{H}\right)_{,w}=0\;,$ (103)
whose integration immediately gives
$A=cH\;,$ (104)
where $c$ is a nontrivial constant.
If the Killing vector is required to be an asymptotically timelike unit Killing vector, one finds the condition
$-1=\left[2\epsilon B-\frac{c-2c\epsilon H(R_{\infty})_{,w}}{H}\right]cH$ (105)
or
$B=\frac{c\epsilon}{2H}-c(R_{\infty})_{,w}-\frac{\epsilon}{2cH}.$ (106)
The relations (104) and (106) show that the components of a timelike Killing vector can be completely fixed by the asymptotic solution of the field equation. In particular, it is determined by the asymptotic behavior of the areal distance $r$ and an asymptotic time like Killing vector should have the form
$\begin{split}\xi=&cH(w)\partial_{w}+\\ &\left[-cH_{,w}\lambda+\frac{c\epsilon}{2H}-c(R_{\infty})_{,w}-\frac{\epsilon}{2cH}\right]\partial_{\lambda}+O(\lambda^{-1}).\end{split}$ (107)
In particular, regarding the choices of $A$ and $B$ in table 1 and requiring the Killing vectors to be asymptotically timelike unit Killing vectors implies using (98), (103) and (106).
  1. i)
    … case $A=1$ and $B=0$, we find $c=1/H$. Moreover, since $r=r(\lambda)$, $H=const$ and $R_{\infty}=const$. Consequently, $\xi^{a}\xi_{a}=-1$ in the asymptotic limit can only be archived if $H=1$, i.e. the frame at large radii is a Bondi frame and the areal distance has the asymptotic expansion
    $r=\lambda+R_{\infty}+O(\lambda^{-1}).$ (108)
  2. ii)
    …case $A=kw$ and $B=0$ gives $kw=cH(w)$. As Table 1 shows $r=r(w\lambda)$, the asymptotic expansion of $r$ in (36a) becomes $r(w\lambda)=\frac{kw\lambda}{c}+R_{\infty}(w)+O(\lambda^{-1})$ which requires $R_{\infty}=const$. But then the normalisation $\xi^{a}\xi_{a}=-1$ implies $c^{2}=1$, so that we can choose $c=1$. Consequently $r$ has the asymptotic expansion
    $r=kw\lambda+R_{\infty}+O(\lambda^{-1}).$ (109)
  3. iii)
    …case $A=\frac{1}{2}w^{2}(w+k)^{-2}$ and $B=1$ gives ${w^{2}=2c(w+k)^{2}H}$. Then fixing $H$ in the asymptotic behavior of $r$ gives for (36a)
    $r(z)=\frac{w^{2}c}{2(w+k)^{2}}\lambda+R_{\infty}(w)+O(\lambda^{-1})\;\;.$ (110)
    Table 1 indicates that
    $r(z)=r\left[\frac{1}{2}\left(\frac{w}{w+k}\right)^{2}\lambda-w\right],$
    which implies by comparison with (110) that
    $R_{\infty}(w)=-cw+r_{\infty}\;\;,r_{\infty}=const.$ (111)
    To be consistient with the invariant, any solution for $r$ of the field equations with this Killing vector, which is in addition an asymptotical timelike unit vector, must have the large $\lambda$ expansion of the form
    $r(z)=r_{\infty}+c\left[w-\frac{w^{2}}{2(w+k)^{2}}\lambda\right]+O(\lambda^{-1}).$ (112)
    The determination of $c$ for the present case is postponed to V.1, because of the subtlety of the behavior of the Killing vector for $w=0$. There, it is a non-vanishing null tangent vector to the hypersurface, consequently, $\xi|_{w=0}=\partial_{\lambda}$. Requiring regularity of this surface (which implies regularity of $V$ given by (100) can be used to fix the value of the constant $c$ (see next section for an explicit example in an extremal Reissner-Nordströn black hole).
Of interest is also the behaviour of the norm of the Killing vector for $\lambda=0$, i.e on $\mathcal{B}^{\epsilon}$. At the origin ($\lambda=0$) of the affine parameter, the norm (101) takes the value
$\xi^{a}\xi_{a}|_{\mathcal{B}^{\epsilon}}=-A^{2}V|_{\mathcal{B}^{\epsilon}}+2\epsilon AB.$ (113)
If $V|_{\lambda=0}=0$ the surface $\lambda=0$ is a null hypersurface $\mathcal{B}^{\epsilon}$ and if $B=0$ and $A\neq 0$ the null hypersurface $\mathcal{B}^{\epsilon}$ is a Killing horizon [^5] If $A=0$, we also have $\xi^{a}\xi_{a}=0$, but it is not a Killing horizon because only a vector field of the form $\xi\propto\partial_{w}$ can be a null generator of a Killing horizon at a null surface determined by $\lambda=0$. However, the possibility $A=B=0$ can be satisfied at an eventual bifurcating Killing horizon, where the Killing vector field vanishes.. Indeed, $B\neq 0$ in case iii) is crucial for the discussion of the (vanishing) surface gravity of an extremal black hole as is the case of extremal Reissner-Nordström black holes discussed in the next section. It turns out that for case iii) there also exists a Killing horizon not necessarily located at $\lambda=0$ but given implicitly by an expression of the form $F(w,\lambda)=0$. In the other cases, the corresponding surface gravity, $\kappa_{H}$, can be found from the invariant relation Wald (1984)
$(\kappa_{H})^{2}=-\frac{1}{2}(\nabla^{a}\xi^{b})(\nabla_{a}\xi_{b})\Big{|}_{\mathcal{B}^{\epsilon}}=\left(A_{,w}-\frac{A}{2}V_{,\lambda}|_{\mathcal{B}^{\epsilon}}\right)^{2}.$ (114)
These properties of the Killing vector will be used in the next section, where the Reissner-Nordström black hole is discussed.
V Exact solutions for the ciBVP
V.1 Charged black hole solutions in null coordinates
For finding charged black hole solutions we work with the double null foliation of Sec. III.3 where the null boundary has nontrivial Misner-Sharp mass.
We assume that the scalar field on the initial slice $\mathscr{N}^{\epsilon}(0)$ and the boundary $\mathcal{B}^{\epsilon}(w)$ is trivial meaning $\mathcal{F}_{\Phi}=\mathcal{N}_{\mathcal{B}}=0$. We use $\kappa=8\pi$ and the most general values for the remaining data at the common intersection $\Sigma(0)$
$\begin{split}&r(0,0)=r_{0}>0\;\;,\;\;r_{,w}(0,0)=r_{N}\;\;,\;\;\\ &Q(0,0)=Q_{0}\;\;,\;\;r_{,\lambda}(0,0)=\Theta\end{split}$ (115)
The boundary equations on $\mathcal{B}^{\epsilon}$ are
$\frac{r_{,ww}}{r}\bigg{|}_{\mathcal{B}^{\epsilon}}$ $=$ $0,$ (116a)
$Q_{,w}\big{|}_{\mathcal{B}^{\epsilon}}$ $=$ $0,$ (116b)
$(rr_{,\lambda})_{,w}\big{|}_{\mathcal{B}^{\epsilon}}$ $=$ $\frac{\epsilon}{2}\left(1-\frac{Q^{2}}{r^{2}}\right)\bigg{|}_{\mathcal{B}^{\epsilon}}.$ (116c)
Integration of (116a) and (116b) the first two gives us
$r(w,0)=$ $r_{\mathcal{B}}:=r_{0}+r_{N}w\;\;,$ (117a)
$Q(w,0)=$ $Q_{0}\;\;,\;\;$ (117b)
inserting these values into (116c) and integration yields
$\begin{split}(r_{,\lambda})\big{|}_{\mathcal{B}^{\epsilon}}=&r_{1}(w):=\frac{r_{0}\Theta}{r_{\mathcal{B}}}+\frac{\epsilon}{2}\left(\frac{r_{0}r_{\mathcal{B}}-Q_{0}^{2}}{r_{0}r_{\mathcal{B}}^{2}}\right)\left(\frac{r_{\mathcal{B}}-r_{0}}{r_{N}}\right)\\ =&\frac{\Theta r_{0}}{r_{N}w+r_{0}}+\frac{(r_{0}r_{N}w-Q_{0}^{2}+r_{0}^{2})\epsilon w}{2r_{0}(r_{N}w+r_{0})^{2}}\;\;.\end{split}$ (118)
The boundary values (117) and (118) provide the values for $Z$ and $\mathcal{L}$ at $\mathcal{B}^{\epsilon}$
$Z(w,0)=2\epsilon r_{\mathcal{B}}r_{N}\;\;,\;\;\mathcal{L}(w,0)=0\;\;.$ (119)
Now the hypersurface equation for the given initial boundary value data are
$r_{,\lambda\lambda}$ $=$ $0,$ (120a)
$Q_{,\lambda}$ $=$ $0,$ (120b)
$\alpha_{,\lambda}$ $=$ $-\epsilon\frac{Q}{r^{2}},$ (120c)
$Z_{,\lambda}$ $=$ $-\frac{Q^{2}}{r^{2}},$ (120d)
$\mathcal{L}_{,\lambda}$ $=$ $0,$ (120e)
$V_{,\lambda\lambda}$ $=$ $-\frac{1}{\lambda}\left(\frac{\lambda^{2}}{r^{2}}{}\right)_{,\lambda}+\frac{2Zr_{,\lambda}}{r^{3}}+\frac{4Q^{2}}{r^{4}}.$ (120f)
The integration of (120) with the boundary data (117), (118) and (119) gives us
$r$ $=$ $r_{\mathcal{B}}+r_{1}\lambda,$ (121a)
$Q$ $=$ $Q_{0},$ (121b)
$\alpha$ $=$ $\epsilon\frac{Q_{0}}{r_{1}}\left(\frac{1}{r}-\frac{1}{r_{\mathcal{B}}}\right),$ (121c)
$Z$ $=$ $2\epsilon r_{\mathcal{B}}r_{N}+\frac{Q_{0}^{2}}{r_{1}}\left(\frac{1}{r}-\frac{1}{r_{\mathcal{B}}}\right),$ (121d)
$V$ $=$ $\left(\frac{\lambda}{r}\right)^{2}\frac{Q^{2}_{0}(r+r_{\mathcal{B}})-(1-2\epsilon r_{N}r_{1})rr_{\mathcal{B}}^{2}}{r_{\mathcal{B}}^{3}}$ (121e)
$=$ $\left(\frac{\lambda}{r}\right)^{2}\left[\frac{Q_{0}^{2}}{r_{\mathcal{B}}^{2}}+\frac{\left(Q^{2}_{0}-r_{\mathcal{B}}^{2}+2\epsilon r_{N}r_{1}r_{\mathcal{B}}^{2}\right)r}{r_{\mathcal{B}}^{3}}\right].$
Consequently the most general form of metric a function of the initial values reads
$\begin{split}ds^{2}=&-\left(\frac{\lambda}{r}\right)^{2}\left[\frac{Q_{0}^{2}}{r_{\mathcal{B}}^{2}}+\left(Q^{2}_{0}-r_{\mathcal{B}}^{2}+2\epsilon r_{N}r_{1}r_{\mathcal{B}}^{2}\right)\left(\frac{r}{r_{\mathcal{B}}^{3}}\right)\right]dw^{2}\\ &+2\epsilon dwd\lambda+r^{2}q_{AB}dx^{A}dx^{B}\;,\\ r=&r_{0}+r_{N}w+\left[\frac{\Theta r_{0}}{r_{N}w+r_{0}}+\frac{(r_{0}r_{N}w-Q_{0}^{2}+r_{0}^{2})\epsilon w}{2r_{0}(r_{N}w+r_{0})^{2}}\right]\lambda,\end{split}$ (122)
in which $Q_{0}$ corresponds to a global charge. This solution was also found in Gallo et al. (2021) using a different approach.
On $\mathcal{B}^{\epsilon}$, it reduces to
$ds^{2}=(r_{0}+r_{N}w)^{2}q_{AB}dx^{A}dx^{B}$ (123)
and for $w=0=dw$ that is on $\mathscr{N}^{\epsilon}(0)$
$ds^{2}=(r_{0}+\Theta\lambda)^{2}q_{AB}dx^{A}dx^{B}$ (124)
Since $w$ and $\lambda$ are the affine parameters on the null hypersurfaces $\mathcal{B}^{\epsilon}(w)$ and $\mathscr{N}^{\epsilon}(0)$, respectively, the initial data parameters $r_{N}$ and $\Theta$ correspond to the expansion rates of the generators of these surfaces.
Calculation of the Misner-Sharp mass (21) gives us
$M=$ $\frac{1}{2}\left[r_{\mathcal{B}}-2\epsilon r_{\mathcal{B}}r_{N}r_{1}+Q^{2}_{0}\left(\frac{1}{r_{\mathcal{B}}}-\frac{1}{r}\right)\right]$ (125)
with the limiting values
$\lim_{w\rightarrow 0}M=$ $\frac{1}{2}\left[r_{0}-2\epsilon r_{0}r_{N}\Theta+Q^{2}_{0}\left(\frac{1}{r_{0}}-\frac{1}{r_{0}+\Theta\lambda}\right)\right],$ (126)
$\lim_{\lambda\rightarrow 0}M=$ $\frac{1}{2}(r_{0}+r_{N}w)\left(1-2\epsilon r_{N}r_{1}\right),$ (127)
so that the composite limit
$\lim_{w\rightarrow 0\atop\lambda\rightarrow 0}M=$ $\frac{r_{0}}{2}(1-2\epsilon r_{N}\Theta)$ (128)
shows that $r_{0}=2M$ if either $r_{N}$ or $\Theta$ vanishes.
As the limit
$\lim_{\lambda\rightarrow\infty}M=$ $\frac{r_{\mathcal{B}}}{2}\left[1-2\epsilon r_{N}r_{1}+\frac{Q^{2}_{0}}{r_{\mathcal{B}}^{2}}\right]$ (129)
gives rise to the Bondi mass, we find the most general expression for the Bondi mass of the solution to be
$m_{B}=\frac{Q_{0}^{2}+r_{0}^{2}}{2r_{0}}-\epsilon r_{0}r_{N}\Theta\;.$ (130)
If either $r_{N}=0$ or $\Theta=0$, the initial value $r_{0}$ can be fixed by the Bondi mass, $m_{B}$, to find
$r_{0,\pm}=m_{B}\pm\sqrt{m_{B}^{2}-Q^{2}_{0}}$ (131)
where the numerical value $r_{0,+}$ is the same value as the value of the radius of the outer event horion of a Reissner-Nordström black hole and $r_{0,-}$ the value of the inner event horizon.
V.1.1 Sub-extremal black holes
Suppose $r_{N}=0$, then the areal distance has the form
$r=r_{0}+\left[\Theta-\epsilon\frac{Q^{2}_{0}-r^{2}_{0}}{2r_{0}^{3}}w\right]\lambda\;\;.$ (132)
Since $w$ parameterises affine geodesics on the boundary $\mathcal{B}^{\epsilon}$ where $\lambda=0$, we have the freedom of choice in the origin of the affine parameter. The transformation
$w\rightarrow\tilde{w}=w-\epsilon\Theta\frac{2r_{0}^{3}}{Q_{0}^{2}-r_{0}^{2}}$ (133)
gives the symmetric expression
$r=\tilde{r}(\lambda\tilde{w}):=r_{0}-\epsilon\frac{Q^{2}_{0}-r^{2}_{0}}{2r_{0}^{3}}(\tilde{w}\lambda)$ (134a)
and
$V$ $=-2\lambda^{2}\frac{\epsilon(\tilde{w}\lambda)(Q_{0}^{2}-r_{0}^{2})^{2}-2r_{0}^{4}(2Q^{2}_{0}-r_{0}^{2})}{\left[\epsilon(\tilde{w}\lambda)(Q_{0}^{2}-r_{0}^{2})-2r_{0}^{4}\right]^{2}}$
$\equiv\lambda^{2}\tilde{V}(\tilde{w}\lambda)$ (134b)
thereby showing the two functions depend on the product $w\lambda$. For simplicity, we will write $\tilde{w}$ as $w$, hereafter.
Therefore, the metric
$\begin{split}ds^{2}=-\lambda^{2}\tilde{V}(\lambda w)dw^{2}+2\epsilon dwd\lambda+\tilde{r}^{2}(\lambda w)q_{AB}dx^{A}dx^{B}\end{split}$ (135)
is invariant under the transformation $(w,\lambda)\rightarrow(aw,a^{-1}\lambda)$ for any number $a$, which implies a Killing vector of the form $\xi=\epsilon k(w\partial_{w}-\lambda\partial_{\lambda})$, where $k$ is a positive normalization constant determined from the calculation of the norm $\xi^{a}\xi_{a}$ and the factor $\epsilon$ is included in order to choose a future timelike Killing vector field. In accordance with the results of Sec. 95, this also gives us $A(w)=kw$ and $B=0$ for the two $w-$dependent functions of the general Killing vector in equation (95). Calculation of the asymptotic limits of the norms gives us
$\lim_{\lambda\rightarrow 0}\xi^{a}\xi_{a}=$ $0\;\;,$ (136)
$\lim_{\lambda\rightarrow\infty}\xi^{a}\xi_{a}=$ $-\left[\frac{2kr_{0}^{3}}{Q_{0}^{2}-r_{0}^{2}}\right]^{2}\;\;.$ (137)
Therefore, the Killing vector is an asymptotic timelike unit Killing vector if the Killing vector has the form
$\xi=\frac{\epsilon}{2r_{0}}\left(1-\frac{Q^{2}_{0}}{r_{0}^{2}}\right)(w\partial_{w}-\lambda\partial_{\lambda})\;\;,$ (138)
As mentioned above, if the norm of the Killing vector, $\xi^{a}\xi_{a}$ vanishes for $\lambda=0$, the null hypersurface $\mathcal{B}^{\epsilon}$ is a Killing horizon. If $\epsilon=1$, this boundary is the future Killing horizon $H^{+}$, and the past event horizon $H^{-}$ for the choice $\epsilon=-1$. Its surface gravity $\kappa_{H}$ can be found from the invariant relation Wald (1984)
$\kappa_{H}^{2}=-\frac{1}{2}(\nabla^{a}\xi^{b})(\nabla_{a}\xi_{b})\Big{|}_{\mathcal{B}^{\epsilon}}=k^{2}=\left[\frac{1}{2r_{0}}\left(1-\frac{Q^{2}_{0}}{r_{0}^{2}}\right)\right]^{2}\;\;.$ (139)
Using the value of $r_{0,+}$ of (131), the surface gravity reduces to the value of the Reissner-Nordström black hole
$\kappa_{H}=\frac{\sqrt{m^{2}_{B}-Q_{0}^{2}}}{2mr_{0,+}-Q_{0}^{2}}=\frac{\sqrt{m^{2}_{B}-Q_{0}^{2}}}{2m_{B}\left(m_{B}+\sqrt{m^{2}_{B}-Q_{0}^{2}}\right)-Q_{0}^{2}}\;,$ (140)
which also gives the Schwarzschild value ${\kappa_{H}=1/(4m_{B})}$ in the case of vanishing global charge, $Q_{0}=0$.
Hence, the Killing vector field can be written as
$\xi=\epsilon\kappa_{H}(w\partial_{w}-\lambda\partial_{\lambda}).$ (141)
Note also that the metric (135) is regular at the null surface $w=0$ given by $\mathcal{N}^{\epsilon}(0)$. This surface is the future (past) Killing horizon $H^{+}$ ($H^{-}$) if $\epsilon=-1$ ($\epsilon=1$) with Killing vector field $\xi=\kappa_{H}\lambda\partial_{\lambda}$ ($\xi=\kappa_{H}w\partial_{w}$). On the other hand, $\xi$ vanishes at the 2-sphere given by $w=\lambda=0$, which constitutes the bifurcating Killing horizon.
V.1.2 Extremal black holes
The surface gravity $\kappa_{H}$ in (140) vanishes if $Q_{0}=m_{B}=r_{0}$, but this goes along with the vanishing of the Killing vector in (138). In fact, in that limit the resulting metric associated to the $\{w,\lambda\}$ coordinates introduced in the last subsection is not the extremal Reissner-Nordström metric, but the so called Bertotti-Robinson metric Gallo et al. (2021). Thus the Killing vector (138) cannot be used to find an extremal Reissner-Nordström black hole solution where the global charge to mass ratio is unity. Indeed, as we will show here, the resulting extremal metric is not invariant under the transformation $(w,\lambda)\rightarrow(aw,a^{-1}\lambda)$, and therefore the resulting Killing vector field can not be written as in (138).
To single out an extremal black hole solution in the general solution (121) for $r$ and $V$ which also has a non-vanishing Killing vector at the horizon of the black hole, consider (130) while setting $m_{e}:=Q_{0}=m_{B}.$ This gives
$m_{e}=\frac{m_{e}^{2}+r_{0}^{2}}{2r_{0}}-\epsilon r_{0}r_{N}\Theta$ (142)
allowing us to express any two of the initial data $(r_{0},\,r_{N},\Theta)$, by the third, i.e.
$r_{0}=$ $\frac{m_{e}}{2\epsilon r_{N}\Theta-1}\left(-1\pm\sqrt{2\epsilon r_{N}\Theta}\right)\;,$ (143)
$r_{N}=$ $\frac{(m_{e}-r_{0})^{2}}{2\epsilon r_{0}^{2}\Theta}\;,$ (144)
$\Theta=$ $\frac{(m_{e}-r_{0})^{2}}{2\epsilon r_{0}^{2}r_{N}}\;.$ (145)
Since in the Schwarzschild case $r_{0}$ equates twice the Bondi mass, it is desirable to relate $r_{0}$ one-to-one with $m_{e}$. In (143) this is achieved only by either setting $\Theta=0$ or $r_{N}=0$ giving in both cases $r_{0}=m_{e}$. In (144), the choice $r_{0}=m_{e}$ gives $r_{N}=0$ and $\Theta\neq 0$ left arbitrary, while for (145) the same choice yields $\Theta=0$ and $r_{N}\neq 0$ left arbitrary.
Thus an extremal black hole solution can be found if $r_{0}=m_{e}=Q_{0}$ and either $r_{N}=0$ and $\Theta\neq 0$ or $\Theta=0$ and $r_{n}\neq 0$. Now consider the solution for the areal distance and metric potential $V$ for these for cases: for the case $(r_{N}=0\;,\;\Theta\neq 0)$ we find
$r=$ $m_{e}+\Theta\lambda$ (146a)
$V=$ $\left(\frac{\lambda}{m_{e}+\Theta\lambda}\right)^{2}$ (146b)
and for the case $(\Theta=0\;,\;r_{N}\neq 0)$
$r=$ $m_{e}+r_{N}w+\left[\frac{\epsilon r_{N}w^{2}}{2(r_{N}w+m_{e})^{2}}\right]\lambda\;,$ (147)
$V=$ ${-\frac{4m_{e}\lambda^{2}[\epsilon\lambda r_{N}^{2}w^{3}-(m_{e}-2r_{N}w)(m_{e}+r_{N}w)^{3}]}{(r_{N}w+m_{e})[w^{4}r_{N}^{2}\lambda^{2}+4\epsilon r_{N}w^{2}(m_{e}+r_{N}nw)^{3}\lambda+4(m_{e}+r_{N}w)^{6}]}}\;.$ (148)
The line element of first solution is
$\begin{split}ds^{2}=&-\left(\frac{\lambda}{m_{e}+\Theta\lambda}\right)^{2}dw^{2}+2\epsilon dwd\lambda\\ &+(m_{e}+\Theta\lambda)^{2}q_{AB}dx^{A}dx^{B}\;,\end{split}$ (149)
but in fact it looks more familiar, if $r(\lambda)$ in (146a) is inverted to $\lambda(r)$, then we find
$V(r)=\frac{1}{\Theta^{2}}\left(1-\frac{m_{e}}{r}\right)^{2}\;\;,\;\;$ (150)
so that the coordinate transformation $\hat{w}=w/\Theta$, $\lambda=(r-m_{e})/\Theta$ leaves us with the well known line element
$ds^{2}=-\left(1-\frac{m_{e}}{r}\right)^{2}d\hat{w}^{2}+2\epsilon d\hat{w}dr+r^{2}q_{AB}dx^{A}dx^{B}.$ (151)
of an extremal Reissner-Nordström black hole with either ingoing($\epsilon=1$) or outgoing polar null coordinates. In such case, $\hat{w}$ is usually represented as $v$, and the past event horizon is reached as $v\to-\infty.$ For $\epsilon=1$, this metric covers the interior region of the past event horizon and the exterior region (in that case $\hat{w}$ is identified with the retarded null coordinate $u$). This coordinate does not cover the future horizon, which is asymptotically reached as $u\to\infty$. Indeed, calculation of the outgoing expansion rate of the surface forming null vector $\ell^{a}\partial_{a}=\partial_{\lambda}$ using (149) gives $2\Theta/(m_{e}+\Theta\lambda)$ which does not vanish for any value of $w$. Thus there is no null surface $w=const$ labelling the future event horizon. In a similar way, for ingoing null coordinates ($\epsilon=1$), the resulting coordinates cover the exterior and interior of the future event horizon.
Thus the solution with initial data $r_{N}=0$ and $\Theta\neq 0$ cannot describe an extremal black hole in a regular manner, i.e. including both horizons. However, we will see that the description of an extremal charged black hole with both horizons is possible for initial data with $\Theta=0$ and $r_{N}\neq 0$.
Consider now the case $\Theta=0$ which gives (147) and (147) for the non trivial components of the metric.
From (147) and (36a), we read off the relevant functions, $H$ and $R_{\infty}$ of the asymptotic expansion of $r$ as
$H(w)$ $=$ $\frac{\epsilon r_{N}w^{2}}{2(r_{N}w+m_{e})^{2}}\;,$ (152)
$R_{\infty}(w)$ $=$ $m_{e}+r_{N}w\;\;.$ (153)
Note also, that the surface $w=0$ is a regular null surface determined by $r=m_{e}$ and where $V=\frac{\lambda^{2}}{m^{2}_{e}}$. On the other hand, from (152) and (107), we find the following asymptotic behaviour of the Killing vector near $w=0$
$\xi=\left[\frac{c^{2}-1}{cr_{N}w^{2}}{(2r_{N}w+m_{e})m_{e}-\frac{r_{N}}{c}}\right]\partial_{\lambda}+O(w)\;,$ (154)
which implies $c=\pm 1$ for the Killing vector be regular and finite at $w=0$. Setting c=1, we obtain from (105) that $B=-r_{N}.$ Therefore, from (95) and (103) we arrive to the final expression for the Killing vector $\xi$
$\xi=\frac{\epsilon r_{N}w^{2}}{2(r_{N}w+m_{e})^{2}}\partial_{w}-\left[c\frac{\epsilon r_{N}m_{e}(w\lambda)}{(r_{N}w+m_{e})^{3}}+r_{N}\right]\partial_{\lambda}.$ (155)
The asymptotic norm of the Killing vector gives a timelike unit vector by construction. Observe the difference to the nonextremal case (138), the null surface $\lambda=0$ is not a Killing horizon for (155). In fact, the boundary surface has an expansion rate proportional to $r_{N}$. An additional surface (to $w=0$) where the Killing vector becomes null is at
$\lambda(w)=-\frac{2\epsilon}{w(w+m_{e})^{2}}.$ (156)
which follows from the calculation of (101). This is the Killing horizon $r=m_{e}$.
Indeed, the parameter $r_{N}$ in (147) and (148) can be eliminated using the scaling relations $w\rightarrow w/r_{N}$ and $\lambda\rightarrow r_{N}\lambda$, so that the final solution for the extremal Reissner-Nordström black hole only depends on the physical parameter $m_{e}$, the Bondi mass, so that
$r=$ $m_{e}+w+\epsilon\left[\frac{w^{2}}{2(w+m_{e})^{2}}\right]\lambda$ (157a)
$V=$ $-\frac{4m_{e}\lambda^{2}}{(m_{e}+w)}\left[\frac{\epsilon\lambda w^{3}-(m_{e}-2w)(m_{e}+w)^{3}}{\lambda^{2}w^{4}+4\epsilon\lambda w^{2}(m_{e}+w)^{3}+4(m_{e}+w)^{6}}\right]$ (157b)
$\xi$ $=\frac{\epsilon w^{2}}{2(w+m_{e})^{2}}\partial_{w}-\left[\frac{\epsilon m_{e}(w\lambda)}{(w+m_{e})^{3}}+1\right]\partial_{\lambda}.$ (157c)
This solution was also obtained by us in Gallo et al. (2021) using an alternative hierarchical system of equations, note however that the $w$ used here must be identified with the $-w$ used in that reference. In difference to the traditional outgoing/ingoing Eddington-Finkelstein coordinates, the $\{w,\lambda,x^{A}\}$ is a regular coordinate system at both horizons.
V.2 FJNW solution in null coordinates
The simplest Killing vector (107) is where $A=1$ and $B=0$. The Killing vector becomes $\xi=\partial_{w}$ and its additional requirement to be asymptotically timelike restricts the asymptotic frame to be a Bondi frame, i.e. $H=1$ so that $r\sim\lambda$ for large values of $\lambda$. Hereafter, we consider a real and uncharged scalar field solution possessing this symmetry. We will find the FJNW solution Fisher (1948); Janis et al. (1968); Wyman (1981); Virbhadra (1997); Ota et al. (2022) in affine null coordinates, which has not been derived explicitly as a solution of an hierarchical system of the Einstein-scalar field equations in affine null coordinates.
Considering $\xi=\partial_{w}$ for an uncharged ($q=0$), real scalar field, the system of equations (29) becomes
$r_{,\lambda\lambda}+\frac{\kappa}{2}r\Phi_{,\lambda}^{2}=0\;,$ (158)
$Z_{,\lambda}=0\;,$ (159)
$L_{,\lambda}=-\frac{\epsilon(\lambda+Z)\Phi_{,\lambda}}{r}\;,$ (160)
$V_{,\lambda\lambda}=-\frac{1}{\lambda}\left(\frac{\lambda^{2}}{r^{2}}\right)_{,\lambda}+\frac{2Zr_{,\lambda}}{r^{3}}-\frac{\kappa\epsilon}{r}\Phi_{,\lambda}L\;,$ (161)
where we used that $\mathcal{L}\rightarrow L$ becomes real. The wave equation gives (20)
$(r^{2}V\Phi_{,\lambda})_{,\lambda}=0$ (162)
In the absence of a scalar field, $\Phi=0$, the solution is trivial, yielding the Schwarzschild solution $V=1+b/r$ with $r=\lambda$, where $b$ is a constant Mädler and Gallo (2023). Insertion of these Schwarzschild values into the wave equation (162) gives us after integration
$\Phi_{,\lambda}=\frac{a}{\lambda^{2}+b\lambda}\;,$ (163)
whose subsequent integration yields
$\Phi=-\frac{a}{b}\ln\Bigl{(}1+\frac{b}{\lambda}\Bigr{)}=-\frac{a}{\lambda}+O(\lambda^{{-2}}),$ (164)
where we applied the asymptotic condition, $\lim_{r\rightarrow\infty}\Phi=0$, for the scalar field. As it is also shown in (164), the parameter $a$ is the scalar monopole. Although $\Phi$ in (164) is a test scalar field, we can use this ansatz for the scalar field initial data to solve the complete Einstein-scalar field equations in a consistent way . In particular, note that if $r^{2}V$ has the same form for the complete system like in the Schwarzschild case, the scalar field will have the same form as in Schwarzschild. The solution of (158) gives us
$r=c_{1}\lambda\Bigl{(}1+\frac{b}{\lambda}\Bigr{)}^{\frac{1}{2}(1-\delta)}+c_{2}\lambda\Bigl{(}1+\frac{b}{\lambda}\Bigr{)}^{\frac{1}{2}(1+\delta)}\;,$ (165)
with
$\delta=\sqrt{1-\frac{2\kappa a^{2}}{b^{2}}}.$ (166)
Reality of $\delta$ requires $2\kappa a^{2}<b^{2}$. Note that if $a=0$, the scalar field vanishes, and therefore the solution must reduce to the Schwarzschild metric. In that case (165) reduces to
$r=(c_{1}+c_{2})\lambda+c_{2}b.$
For $a=0$, both $r$ and $\lambda$ serve as affine parameters. Thus, we can freely choose $r=\lambda$, which leads to $c_{1}=1$ and $c_{2}=0$. For this reason, in the following we have
$r=\lambda\Bigl{(}1+\frac{b}{\lambda}\Bigr{)}^{\frac{1}{2}(1-\delta)}\;.$ (167)
The asymptotic limit of (167) gives
$r=\lambda+\frac{b}{2}\left(1-\delta\right)+O(\lambda^{-1})$ (168)
which from (36a) implies
$H$ $=$ $1,$ (169)
$R_{\infty}$ $=$ $\frac{b}{2}\left(1-\delta\right).$ (170)
Next, using the notation as in (36d), integration of (159) yields $Z(\lambda)=Z_{\infty}=const,$ which is related to the Bondi mass by (41), i.e.
$Z_{\infty}=R_{\infty}-2m_{B}=\frac{1}{2}\left[b(1-\delta)-4m_{B}\right].$ (171)
Then integration of (160) yields
$\begin{split}L=&L_{\infty}+\frac{2a\epsilon}{(\delta-1)b}\left(1+\frac{b}{\lambda}\right)^{-\frac{1-\delta}{2}}\\ &+\frac{2\epsilon aZ_{\infty}}{b(\delta+1)}\left(\frac{1}{\lambda}-\frac{2}{b(\delta-1)}\right)\left(1+\frac{b}{\lambda}\right)^{-\frac{1-\delta}{2}}.\end{split}$ (172)
The integration constant $L_{\infty}$ vanishes because of the Killing vector $\partial_{w}$ and considering (39). Next, integration of (161) gives us
$\begin{split}V=&V_{1}\lambda+V_{0}+\frac{2\lambda}{b(\delta-1)}\left(1+\frac{b}{\lambda}\right)^{\delta}\\ &+\frac{2Z_{\infty}\left[b(\delta-1)-2\lambda\right]}{b^{2}(\delta^{2}-1)}\left(1+\frac{b}{\lambda}\right)^{\delta}\end{split}$ (173)
Performing an asymptotic expansion of $V$ as in (36g) gives
$\begin{split}V=&\left[V_{1}+\frac{(2\delta+2)b-4Z_{\infty}}{b^{2}(\delta^{2}-1)}\right]\lambda+\left[V_{0}+\frac{2b\delta-2Z_{\infty}}{(\delta-1)b}\right]\\ &+O(\lambda^{-1})\end{split}$ (174)
Taking into account (37), (38) and the Killing symmetry implies $V_{\infty}=1$ and $U_{\infty}=0$, which gives us
$V_{0}=$ $\frac{2Z_{\infty}-(1+\delta)b}{(\delta-1)b}$ (175)
$V_{1}=$ $\frac{2(1+\delta)b-4Z_{\infty}}{b^{2}(1-\delta^{2})}$ (176)
These relations allow us to write $V$ as
$\begin{split}V=&\frac{2\lambda\Bigl{[}b(\delta+1)-2Z_{\infty}\Bigr{]}}{b^{2}(\delta^{2}-1)}\left[\left(\frac{\lambda+b}{\lambda}\right)^{\delta}-1\right]\\ &+\frac{2Z_{\infty}}{b(\delta+1)}\left(\frac{\lambda+b}{\lambda}\right)^{\delta}.\end{split}$ (177)
Finally, the asymptotic expansion of (177) which gives
$V=1+\frac{b\delta}{\lambda}+O(\lambda^{-2}),$ (178)
we can connect the Bondi mass with $\delta$ and $b$, namely $\delta b=-2m_{B}$. Non-negativity of the Bondi mass requires $b\leq 0$. Hence, from (171) we can express $Z_{\infty}$ in terms of $b$ and $\delta$ as
$Z_{\infty}=\frac{b(\delta+1)}{2}.$
By replacing this values into (177) we arrive to our final expression for $V$
$V=\left(1+\frac{b}{\lambda}\right)^{\delta}.$ (179)
The metric is the one of the classical FJNW solution Fisher (1948); Janis et al. (1968); Wyman (1981); Virbhadra (1997); Ota et al. (2022)
$\begin{split}ds^{2}=&-\Bigl{(}1+\frac{b}{\lambda}\Bigr{)}^{\delta}dw^{2}+2\epsilon dwd\lambda\\ &+\lambda^{2}\Bigl{(}1+\frac{b}{\lambda}\Bigr{)}^{1-\delta}q_{AB}dx^{A}dx^{B}.\end{split}$ (180)
The corresponding nontrivial components of the contravariant metric are
$g^{w\lambda}=\epsilon\;\;,\;\;g^{\lambda\lambda}=\Bigl{(}1+\frac{b}{\lambda}\Bigr{)}^{\delta}\;\;,\;\;g^{AB}=\lambda^{-2}\Bigl{(}1+\frac{b}{\lambda}\Bigr{)}^{\delta-1}q^{AB}.$ (181)
VI Summary
We discussed the CIBVP in spherical symmetry of the Einstein equations minimally coupled to a complex and charged massless scalar field. Charting spacetime with an affine null metric at a family of null hypersurfaces, the field equations can be cast into a hierarchical system of hypersurface-evolution equations of which the evolution equation is a first order transport equation for the complex initial data of this system. The initial data consist of an arbitrary profile of the scalar field on an initial null hypersurface. The boundary values required to solve the hypersurface equations are determined from the type of boundary at which the family of null hypersurfaces is specified. Different CIBVPs were discussed; i) an asymptotic CIBVP for large values of the affine parameter, ii) a CIBVP at the central geodesic of spherical symmetry, and iii) two realizations of double null CIBVPs.
The solution of the asymptotic CIBVP is represented by two functions, one for a redshift factor $H$ and another being the complex scalar news function $\mathcal{N}_{\mathcal{B}}$ (which is the $w$ derivative of the scalar field’s monopole) and in addition there are three scalar values at a given coordinate value $w=w_{0}$, representing the complex monopole of the scalar field, the initial Bondi mass and the initial value of the global charge of the system. The redshift factor indicates whether an event horizon forms, i.e. $H\rightarrow 0$, or whether the solution approaches asymptotically Minkowkian values, $H\rightarrow 1$. Given a continous redshift function $H>0$ for $w\geq w_{0}$, an asymptotic inertial coordinate system can be established in which $H=1$ for $w\geq 0$. Such coordinate frame is called a Bondi frame. Indeed, if $\epsilon=-1$ and the asymptotic observer is in a Bondi frame, the $w$ coordinate corresponds to the retarded time $u$, while if $\epsilon=1$ the meaning of the $w$ coordinate is the advanced time $v$. Evaluation of the field equations in a Bondi frame significantly simplifies the solution of the field equation. In such Bondi frame, it is easy to see that the news function completely determines the future($\epsilon=-1$)/past($\epsilon=+1$) behaviour of the scalar monopole, the Bondi mass and the global charge. This behaviour is the spherically symmetric scalar field analogue to the asymptotic CIBVP for electro-vacuum space times, see e.g. van der Burg (1969); Bicak and Pravdova (1998).
If the boundary degenerates to the central world line of spherical symmetry so that the family of null hypersurfaces are a family of null cones whose vertices are traced by the world line, the regularity conditions at the vertices completely fix the boundary conditions to constant values along the world line. As a consequence, the local evolution of the scalar field along the world line is fully determined by the (first) derivative of the scalar field at the world line. As such, by specification of finite initial data at the vertex of an initial null cone $w=w_{0}$ and requiring the local regularity conditions for the hypersurface variables for $w>w_{0}$, one can completely determine the evolution of the scalar field and the resulting structure of the metric for all values $w>w_{0}$. Numerical simulations of the real analog of the presented hypersurface-evolution system at a world line was presented first in Crespo et al. (2019) and whose continuation Mädler et al. (2024) also showed the regular center can be maintained in spacetimes forming black holes.
The double null CIBVP was considered for the two cases where the Misner-Sharp mass of the null boundary surface is either trivial or non-trivial. In both cases, the only required initial data on given null hypersurface $w=w_{0}$ is the first $\lambda-$derivative of the scalar field on the initial data surface. At the common intersection of the initial data surface $w=w_{0}$ and the null boundary, a value for the scalar field must be specified. Moreover, an analogue to the news function, referred to as boundary news function, needs to be specified to be able to find the boundary values required to solve the hypersurface-evolution system. Like for the asymptotic CIBVP, this boundary news function needs to be specified a priori and is completely independent from the initial data profile on $w_{0}$.
For the case, where null boundary surface has vanishing Misner-Sharp mass, in fact all but two of these boundary values are zero. One nontrivial and constant boundary value is the first $\lambda-$derivative of the areal distance that can be set to unity for $w\geq w_{0}$. The other one is the first derivative of $g^{\lambda\lambda}$ evaluated at the null boundary, but this field only needs to be specified as initial value at the cross section between initial data surface and null boundary. This boundary value is in fact the inaffinity of the generating null vectors of the boundary surface where $w=w_{0}$. Its values for $w>w_{0}$ on the boundary are completely determined by the boundary news function. To summarize, the double-null initial-boundary value problem with a null boundary of zero Misner sharp mass requires (apart from trivial and unity values) the specification of the scalar field value and inaffinity at the intersection of the initial data surface and the null boundary, the specification of a free news function on the boundary and the specification of the $\lambda-$derivative of the scalar field on the initial data surface.
If the Misner-Sharp mass of the boundary is nonzero, the values giving rise to the mass and expansion rates of the generators of the boundary and initial data surface are to be specified at the cross section of the two surfaces. In addition, a value for the charge and the scalar field ought to be specified at this intersection. These initial values are propagated along the null boundary after specifying the boundary news function. The set of four differential equations governing this propagation is a hierarchical set of (ordinary) differential equations along the generators of the null boundary. The evolution of the boundary values is completely decoupled from the evolution of the initial data via the hypersurface evolution equations and it can therefore be done stand alone and prior to any evolution of the initial data. This is in contrast to the timelike null problem at the world line where the evolution of the scalar field along the world line requires data (the $\lambda$ derivative) depending on the neighborhood of the world line.
The different initial boundary value formulations can be used for numerically investigating different physical scenaria. The worldline-null problem is suited for studies of charged black hole formation and critical phenomena based on regular and asymptotically flat initial data of a charged scalar field. Currently some of these studies are ongoing as an extension of the previous work employing a real scalar field Mädler et al. (2024). This work used a coordinate compactification of the affine parameter together with some rescaling of the metric field to map null infinity to a finite value. Regarding the formation of charged black holes, it is natural to ask how close to unity the charge-to-mass ratio of such black holes can be found by evolving charged scalar fields. Such simulations would test numerically the recent proof of the third law of thermodynamics Kehle and Unger (2022), which is still outstanding. A confirmation of the disproof of the third law requires that extremal charged black holes (for which the charge-to-mass ratio is equal to one) form in finite (computational) time. But, previous investigations only reach values of $\sim 0.6$ for the charge-to-mass ratio Torres and Alcubierre (2014). The double null CIBVP with a massive null boundary is well suited to investigate scattering Baake and Rinne (2016) or perturbations of a spherically symmetric charged black hole Gelles and Pretorius (2025) since the horizon of the black hole can be taken as a boundary surface. Given the recent success of using null geodesics stability analysis via Lyapynov exponents Ianniccari et al. (2024) to explain black hole formation, critical phenomena Choptuik (1993) or even black hole shadows Gallo and Mädler (2025), it may be interesting to explore how the presented CIBVP can be adopted to investigate such phenomena numerically. A formulation of a double null CIBVP at a horizon may be an option for such study.
To better understand a spherically symmetric charged black hole in affine null coordinates, we derived the subextremal and extremal Reissner-Nordström solution using the developed affine-null metric double null CIBVP. For that aim, we also looked at the general solution of Killing equations. It turns out that there is an entire family of Killing vectors depending on two free functions depending on the $w$ coordinate. In deed, a restriction of the Killing vectors to be asymptotically timelike, while using the general solution of the asymptotically flat initial boundary value problem, showed that the frame at large radii does not need to necessarily be a Bondi frame, where the redshift factor is unity. This is in fact only the case if the Killing vector has the simple form $\partial_{w}$. For subextremal and extemal black holes, the results of Sec. V.1 show that corresponding Killing vectors have much more elaborated functional dependence.
Assuming the simplest Killing vector, we were also able to explicitly derive the FJNW solution based on a guess of the scalar field in a Schwarzschild background. It would be interesting to see if a charged version of the FJNW solution can be singled out within the presented formalism.
Acknowledgements.
The work of R.G. is supported by ANID FONDECYT Regular No. 1220965. E.G. acknowledges financial support from CONICET and SeCyT-UNC. T.M. thanks J. Winicour, G. Khanna and S. Liebling for discussions of aspects on the subject. We also thank O. Baake for reading and commenting on the manuscript.

### v3

@@
 The initial boundary value problem for a family of null hypersurfaces is specified for a timelike-null foliation at the central geodesic of spherical symmetry as well as for a double-null foliation where the corresponding boundary is a null hypersurface.
-For the latter, two distinct boundary value formulations arise – one where the null boundary has zero Misner-Sharp mass and another one where the corresponding Misner-Sharp mass is nonzero.
+For the latter, two distinct boundary value formulations arise; one where the null boundary has zero Misner-Sharp mass and another one where the corresponding Misner-Sharp mass is nonzero.
 As an application, the nonextremal and the extremal Reissner-Nordström solution in null coordinates for a charged black hole and the Fisher-Janis-Newman-Winicour solution are derived.
@@
 The Bondi mass loss formula was found in an asymptotic patch of a spacetime where the radial coordinate parametrizes the null rays of outgoing null cones.
-The long term stable black hole simulations were possible by numerically representing the spacetime manifold with a countable number of finite size patches discretized by finite difference operators adapted to either in- or outgoing null hypersurfaces.
+The long term stable black hole simulations were possible by numerically representing the spacetime manifold with a countable number of finite size patches discretised by finite difference operators adapted to either in- or outgoing null hypersurfaces.
 The technical base of the disproof of the third law was the novel technique called “characteristic gluing” Aretakis et al.
 (2021), where different characteristic (null) hypersurfaces are stitched together using gauge invariant quantities.
-These three examples have in common that the patch(es) is (are) charted with at least one null coordinate, i.e.
-a coordinate constant along a null hypersurface, and that boundary information at one side of the patch needs to be provided to archive the physical result.
-For the Bondi mass loss case, this boundary data is the specification of an asymptotic inertial observer, an initial value for the mass and the so-called news function which is the retarded time derivative of the gravitational shear.
-For the black hole simulation, the boundary data were the specification of black hole data at an inner world tube excising the black hole singularity, where the first patch of the numerically discretized null hypersurface is located.
-While in the disproof of the third law, the boundary data between the patches are specification of quantities in Minkowski space and the horizon of a Schwarzschild/Reissner-Nordström black hole.
+These three examples have in common that the patch(es) is (are) equipped with at least one null coordinate, i.e.
+a coordinate constant along a null hypersurface, and that boundary information at one side of the patch needs to be provided to achieve the physical result.
+For the Bondi mass loss case, this boundary data are the specification of an asymptotic inertial observer, an initial value for the mass and the so-called news function which is the retarded time derivative of the gravitational shear.
+For the black hole simulation, the boundary data were the specifications of black hole data at an inner world tube excising the black hole singularity, where the first patch of the numerically discretised null hypersurface is located.
+While in the disproof of the third law, the boundary data between the patches are specifications of quantities in Minkowski space and the horizon of a Schwarzschild/Reissner-Nordström black hole.
 The difference for the latter with respect to the former two is that the employed initial boundary value problem is based on the two intersecting null hypersurfaces instead of only one null hypersurface.
 A further difference is that in the last example the Einstein equations are solved with a coupling to a charged scalar field, while the first two concerned the full Einstein equations in vacuum.
-Although in spherical symmetry, the coupling to the scalar field captures dynamical features and adds the necessary degrees of freedom so that the discussion of a characteristic initial boundary value problem (CIBVP) can be made without loosing too much structural information.
+In spherical symmetry, the coupling to the scalar field captures dynamical features and adds the necessary degrees of freedom so that the discussion of a characteristic initial boundary value problem (CIBVP) can be made without losing too much structural information.
 Here, we investigate characteristic methods for Einstein equations minimally coupled to a charged, massless and necessary complex scalar field in spherical symmetry.
@@
 (2024); Kehle and Unger (2024a, b, 2022); Kehle and Van de Moortel (2024) related to characteristic gluing Aretakis et al.
-(2021) as well as the studies of formation of sub-extremal black hole formation of Reisner-Nordström black holes Murata et al.
-(2013); Gelles and Pretorius (2025) with characteristic methods using a double null foliation, we will investigate the initial boundary value problem in affine null metric coordinates Winicour (2013).
-The affine null metric formulation is build on a foliation with respect to a family of null hypersurfaces and has the advantage that a simple evolution algorithm can be found, which is beneficial for numerical work.
-For this evolution scheme, the field equations can be cast into one set of hierarchical hypersurface equations and one first order transport equation with respect to the chosen null coordinate.
+(2021) as well as the studies of formation of subextremal Reissner-Nordström black holes Murata et al.
+(2013); Gelles and Pretorius (2025) with characteristic methods using a double-null foliation, we will investigate the CIBVP in affine-null metric coordinates Winicour (2013).
+The affine-null metric formulation is built on a foliation with respect to a family of null hypersurfaces and has the advantage that a simple evolution algorithm can be found, which is beneficial for numerical work.
+As we will demonstrate, for this evolution scheme, the field equations can be cast into one set of hierarchical hypersurface equations and one first order transport equation with respect to the chosen null coordinate.
 The set of derived equations is similar to a previously used one for a real scalar field Crespo et al.
@@
 at an apparent horizon.
-Here, we extend and modify the scheme of Crespo et al.
-(2019) such that this regularisation is maintained if the scalar field is complex and a Maxwell field is added to the energy momentum tensor.
-We will also show that for the derived set of equations, CIBVPs can be formulated either on the central world line in spherical symmetry or by two different versions of double null CIBVPs at a null boundary.
-As an example for these cases, we discuss the asymptotic CIBVP showing that the entire asymptotic behaviour of a charged scalar field is governed by the specification of an initial value for the scalar field monopole, mass, charge and a free function - the scalar field news function.
-We further derive the subextremal and extremal Reisner-Nordström black hole as well as the Fisher-Janis-Newman-Winicour (FJNW) solution in the affine null-metric coordinate chart.
-Those solution turn out to be expressible by simple rational functions.
+In fact, with an extension of Crespo et al.
+(2019), it is possible to investigate the Choptuik phenomenon Choptuik (1993) for charged scalar fields by globally foliating spacetime to include null infinity on the numerical grid through coordinate compactification111 It was shown in Mädler and Gallo (2025) that the conformal Einstein-scalar field equations in affine-null coordinates Friedrich (1986); Penrose (1963) are equivalent to corresponding physical Einstein-scalar field equations, provided the latter employ a compactified affine parameter and suitable normalizations of the metric and scalar field.
+Thus, null infinity can be represented on a numerical grid using a compactified coordinate..
+Traditionally, such investigations have been carried out only within the 3+1 formulation of the Einstein-Maxwell-scalar field equations Gundlach and Martin-Garcia (1996), which does not usually allow this.
+These studies can also address the formation process of charged, spherically symmetric black holes.
+Indeed, the CIVBP presented below can be used to numerically test the mathematical results of Kehle (2019), particularly regarding the disproof of the third law of black hole thermodynamics, that is, as a consequence of Kehle and Unger (2022), extremal black holes should form in finite time.
+In a numerical simulation, this would imply identifying an extremal black hole in the output of a simulation where the black hole forms within finite simulation time.
+In this context, it is significantly simpler to study the formation of extremal black holes in the Einstein-Maxwell-scalar field system in spherical symmetry rather than to study the formation of extremal Kerr black holes.
+This is because, in the former, the simulations can be performed with a single spherically symmetric dynamical field (i.e., the scalar field) , whereas the latter requires at least axial symmetry for the Einstein equations, involving two dynamical fields (i.e.
+the gravitational wave degrees of freedom).
+Affine-null coordinates can also be useful for constructing outgoing null coordinate systems that not only describe the exterior of black holes, but are also sufficiently regular to describe the region inside the event horizon.
+Coordinate systems of this kind for the Schwarzschild case include, for example, the Israel coordinates Israel (1966), which were later generalized to include a broader family of metrics inGallo et al.
+(2021).
+Other null coordinate systems have been constructed numerically for the Kerr black hole in Argañaraz and Moreschi (2021), and used to describe the evolution of a scalar field penetrating the black hole interior in Argañaraz and Moreschi (2022).
+A fully analytical construction of an affine-null coordinate system to describe slowly rotating Kerr black holes is also possible, as demonstrated in Mädler and Gallo (2023).
+The general formalism presented in this work for spherically symmetric spacetimes accommodates such regular affine-null coordinate systems as a particular case, thereby allowing for the study of matter fields radiating in the exterior region while simultaneously tracking their evolution as they penetrate into the interior of black holes.
+In this work, we extend and modify the scheme of Crespo et al.
+(2019) so that the regularization of the fields involved in the hierarchical affine-null system of equations is preserved when the scalar field is complex and a Maxwell field is included in the energy-momentum tensor.
+We will also show that for the derived set of equations, CIBVPs can be formulated either on the central world line in spherical symmetry or by two different versions of double-null CIBVPs at a null boundary.
+As an example for these cases, we discuss the asymptotic CIBVP showing that the entire asymptotic behavior of a charged scalar field is governed by the specification of an initial value for the scalar field monopole, mass, charge and a free function - the scalar field news function.
+We further derive the subextremal and extremal Reissner-Nordström black hole as well as the Fisher-Janis-Newman-Winicour (FJNW) solution in the affine-null-metric coordinate chart.
+Those solutions turn out to be expressible by simple rational functions.
 In Sec.
-II, the basic field equations and additional geometric quantities for the Einstein-Maxwell scalar field equations in spherical symmetry and affine null metric coordinates are presented.
-The regularised set of hypersurface-evolution equations following from the field equations being the CIBVPs is given in Sec.
+II, the basic field equations and additional geometric quantities for the Einstein-Maxwell-scalar field equations in spherical symmetry and affine-null metric coordinates are presented.
+The regularized set of hypersurface-evolution equations following from the field equations being the CIBVPs is given in Sec.
 III.
-The CIBVP in an asymptotic region is presented in III.1, the one at the central geodesic is formulated in III.2, while the two different CIBVPs with a null boundary surface are given in Sec.
+The CIBVP in an asymptotic region is presented in Sec.
+III.1, the one at the central geodesic is formulated in III.2, while the two different CIBVPs with a null boundary surface are given in Sec.
 III.3.
 In Sec.
-IV, we discuss the properties for asymptotic timelike Killing vectors, which are needed for the derivation of the Reisner-Nordström black hole in Sec.
+IV, we discuss the properties for asymptotic timelike Killing vectors, which are needed for the derivation of the Reissner-Nordström black hole in Sec.
 V.1 and the FJNW solution in Sec.
 V.2.
-Sec.
-VI contains a summary and conclusion of the work.
-We use geometric units with $c=1$ and the Einsteinian gravitational constant is $\kappa=8\pi G$.
-The metric signature is $+2$, capital latin indices $A,B,C,...$ are two dimensional with values $x^{A}=(x^{2},x^{3})$ and lowercase latin indices are four dimensional taking values $x^{a}=(w,\lambda,x^{A})$.
+Section VI contains a summary and conclusion of the work.
+We use geometric units with $c=1$ and the Einstein gravitational constant is $\kappa=8\pi G$.
+The metric signature is $+2$, capital Latin indices $A,B,C,\ldots$ are two dimensional with values $x^{A}=(x^{2},x^{3})$ and lowercase Latin indices are four dimensional taking values $x^{a}=(w,\lambda,x^{A})$.
 The covariant derivative of the metric is $\nabla_{a}$ and partial derivatives are denoted with the comma notation $\frac{\partial f}{\partial x^{a}}=f_{,a}$.
-Calligraphic capital letters, $\mathcal{J},\mathcal{L},\mathcal{C},...$ are always assumed complex fields, while Roman capital letters, like $Q,Z,V,L,\ldots$ are purely real.
-Numeral subscripts in parenthesis, like $V_{(1)}$, correspond to expansion coefficients of a power series evaluated at $\lambda=0$ with the number in the parenthesis referring to the power $\lambda^{k}$.
-Numeral subscripts in a bracket, like $\Phi_{[1]}$, correspond to expansion coefficients of an asymptotic power series and the number in the bracket refers to the order of the inverse power $1/\lambda^{k}$.
+Calligraphic capital letters, $\mathcal{J},\mathcal{L},\mathcal{C},...$ are always assumed complex fields, while Roman capital letters, such as $Q,Z,V,L,\ldots$ are purely real.
+Numeral subscripts in parentheses, like $V_{(1)}$, correspond to expansion coefficients of a power series evaluated at $\lambda=0$ with the number in the parentheses referring to the power $\lambda^{k}$.
+Numeral subscripts in brackets, such as $\Phi_{[1]}$, correspond to expansion coefficients of an asymptotic power series and the number in the bracket refers to the order of the inverse power $1/\lambda^{k}$.
 II Field equations for the Einstein-Maxwell-scalar field equations
-In a four dimensional spacetime, consider the Einstein-Hilbert action $S$ given by
-$S=\int d^{4}x\sqrt{-g}\left[\frac{g^{ab}R_{ab}}{2\kappa}-\frac{F_{ab}F^{ab}}{16\pi}-\frac{1}{2}(\overline{\mathcal{D}^{a}\Phi})(\mathcal{D}_{a}\Phi)\right]$ (1)
-where $\kappa=8\pi G$ is the Einsteinian gravitational constant, $\sqrt{-g}$ the determinant of the metric $g_{ab}$ with well defined inverse $g^{ab}$ and corresponding Levi-Civita covariant derivative $\nabla_{c}$, $R_{ab}$ is the Ricci tensor wrt $g_{ab}$, $F_{ab}=\nabla_{a}A_{b}-\nabla_{b}A_{a}$ is a Maxwell field determined by the 4-covector $A_{b}$, $\mathcal{D}_{c}=\nabla_{c}+iqA_{c}$ a gauge invariant derivative for the massless complex scalar field $\Phi$ that is coupled to the Maxwell field with the gauge charge $q$.
+In a four-dimensional spacetime, consider the Einstein-Hilbert action $S$ given by
+$S=\int d^{4}x\sqrt{-g}\left[\frac{g^{ab}R_{ab}}{2\kappa}-\frac{F_{ab}F^{ab}}{16\pi}-\frac{1}{2}(\overline{\mathcal{D}^{a}\Phi})(\mathcal{D}_{a}\Phi)\right],$ (1)
+where $\kappa=8\pi G$ is the Einsteinian gravitational constant, $\sqrt{-g}$ is the determinant of the metric $g_{ab}$ with a well defined inverse $g^{ab}$ and a corresponding Levi-Civita covariant derivative $\nabla_{c}$, $R_{ab}$ is the Ricci tensor with respect to $g_{ab}$, $F_{ab}=\nabla_{a}A_{b}-\nabla_{b}A_{a}$ is a Maxwell field determined by the four-covector $A_{b}$, and $\mathcal{D}_{c}=\nabla_{c}+iqA_{c}$ is a gauge invariant derivative for the massless complex scalar field $\Phi$ that is coupled to the Maxwell field with the gauge charge $q$.
 The overbar denotes complex conjugation.
-From the variation of the action with respect to $g^{ab}$, $A_{a}$, $\bar{\Phi}$ and $\Phi$, we can derive the trace reversed Einstein equations, $E_{ab}$, the Maxwell equations, $E_{a}$, and the scalar wave equations $E$ and $\bar{E}$, which are respectively
+From the variation of the action with respect to $g^{ab}$, $A_{a}$, $\bar{\Phi}$ and $\Phi$, we can derive the trace reversed Einstein equations, $E_{ab}$, the Maxwell equations, $E_{a}$, and the scalar wave equations $E$ and $\bar{E}$, which are, respectively
 $E_{ab}:=$ 00 0 $=R_{ab}-\frac{\kappa}{4\pi}\left(g^{cd}F_{ac}F_{bd}-\frac{1}{4}g_{ab}F^{cd}F_{cd}\right)$ (2)
-$-\kappa\left[\frac{1}{2}(\overline{\mathcal{D}_{a}\Phi})(\mathcal{D}_{b}\Phi)+\frac{1}{2}(\overline{\mathcal{D}_{b}\Phi})(\mathcal{D}_{a}\Phi)\right]$
-$E_{a}:=$ 00 0 $=\nabla_{a}F^{a}_{\;b}+4\pi j_{b}$ (3)
-$E:=$ 00 0 $={\mathcal{D}^{a}}\mathcal{D}_{a}\Phi$ (4)
-$\bar{E}:=$ 00 0 $=\overline{\mathcal{D}^{a}\mathcal{D}_{a}\Phi}$ (5)
+$-\kappa\left[\frac{1}{2}(\overline{\mathcal{D}_{a}\Phi})(\mathcal{D}_{b}\Phi)+\frac{1}{2}(\overline{\mathcal{D}_{b}\Phi})(\mathcal{D}_{a}\Phi)\right],$
+$E_{a}:=$ 00 0 $=\nabla_{a}F^{a}_{\;b}+4\pi j_{b},$ (3)
+$E:=$ 00 0 $={\mathcal{D}^{a}}\mathcal{D}_{a}\Phi,$ (4)
+$\bar{E}:=$ 00 0 $=\overline{\mathcal{D}^{a}\mathcal{D}_{a}\Phi},$ (5)
 with the current
-$j_{b}=\frac{iq}{2}\left(\overline{\Phi}\mathcal{D}_{a}\Phi-\Phi\overline{\mathcal{D}_{a}\Phi}\right)$ (6)
+$j_{b}=\frac{iq}{2}\left(\overline{\Phi}\mathcal{D}_{a}\Phi-\Phi\overline{\mathcal{D}_{a}\Phi}\right),$ (6)
 […]
 […]
-with $g_{AB}$ being the 2-metric of the spacelike hypersurface of the co-dimension 2 boundary $\Sigma=\partial\mathtt{V}$ where the integral curves of null vectors $\ell^{a}\partial_{a}$ and $n^{a}\partial_{a}$ orthogonal to $\Sigma$ intersect.
-Let us consider the metric
-$ds^{2}=-V(w,\lambda)dw^{2}+2\epsilon dwd\lambda+r^{2}(w,\lambda)q_{AB}dx^{A}dx^{B}$ (8)
+with $g_{AB}$ being the two-metric of the spacelike hypersurface of the co-dimension two boundary $\Sigma=\partial\mathtt{V}$, which is the intersection of the integral curves of null vectors $\ell^{a}\partial_{a}$ and $n^{a}\partial_{a}$ orthogonal to $\Sigma$.
+Consider the metric
+$ds^{2}=-V(w,\lambda)dw^{2}+2\epsilon dwd\lambda+r^{2}(w,\lambda)q_{AB}dx^{A}dx^{B},$ (8)
 where $q_{AB}$ is any metric of a unit sphere [^3] In standard spherical coordinates $x^{A}=(\theta,\varphi)$, ${q_{AB}=\mathrm{diag}(1,\sin^{2}\theta)}$ with the three associated Killing vectors $\xi_{1}=\partial_{\phi}$, $\xi_{2}=\cos\phi\partial_{\theta}-\cot\theta\sin\phi\partial_{\phi}$ and $\xi_{3}=\sin\phi\partial_{\theta}+\cot\theta\cos\phi\partial_{\phi}$.
@@
 Of the field equations, only (17) would change by the substitution of the term +1 by either $+0$ or $-1$, respectively.
-The former may be interesting for plane wave front propagation, like pp waves, while the latter for hyperbolical black holes or fluids, see e.g.
+The former may be interesting for plane wave front propagation, such as pp waves, while the latter for hyperbolical black holes or fluids (see e.g.
 Herrera et al.
-(2020, 2021)..
-An understanding of the the binary parameter $\epsilon$ can be found in flat space where $V=1$ and $r=\lambda$, then for $\epsilon=-1$ the coordinate surfaces $w=const$ represents outgoing null hypersurfaces $u$, while for $\epsilon=1$ they are ingoing null hypersurfaces $v=const$.
+(2020, 2021))..
+An understanding of the binary parameter $\epsilon$ can be found in flat space, where $V=1$ and $r=\lambda$, then for $\epsilon=-1$ the coordinate surfaces $w=$const represent outgoing null hypersurfaces $u$=const, while for $\epsilon=1$ they are ingoing null hypersurfaces $v=$const.
 The inverse metric has the components
 $g^{w\lambda}=\epsilon\;\;,\;\;g^{\lambda\lambda}=V\;\;,\;\;g^{AB}=\frac{1}{r^{2}}q^{AB}\;,$ (9)
-where we used $\epsilon^{2}=1$. Furthermore, the form of the metric (8) remains invariant under the coordinate changes
+where we used $\epsilon^{2}=1$. The form of the metric (8) remains invariant under the coordinate changes
 […]
-For the metric we define two null vectors
+For the metric, we define two null vectors
 $\ell^{a}\partial_{a}=\partial_{\lambda}\;\;,\;\;n^{a}\partial_{a}=-\epsilon\partial_{w}-\frac{V}{2}\partial_{\lambda}\;,$ (11)
-whose associated one forms are $\ell_{a}dx^{a}=\epsilon dw$ and $n_{a}dx^{a}=\epsilon(V/2)dw-d\lambda$ and which are orthogonal to one another $n^{a}\ell_{a}=-1$.
-The null vector field $\ell^{a}\partial_{a}$ is tangent to affine parameterized null geodesics meaning $\ell^{a}\nabla_{a}\ell^{a}=0$, while the vector field $n^{a}\partial_{a}$ is in general tangent to non-affine parametrized null geodesics where
+whose associated one forms are $\ell_{a}dx^{a}=\epsilon dw$ and $n_{a}dx^{a}=\epsilon(V/2)dw-d\lambda$ satisfying $n^{a}\ell_{a}=-1$.
+The null vector field $\ell^{a}\partial_{a}$ is tangent to affinely parametrized null geodesics meaning $\ell^{a}\nabla_{a}\ell^{a}=0$, while the vector field $n^{a}\partial_{a}$ is in general tangent to nonaffinely parametrized null geodesics, where
 $n^{a}\nabla_{a}n^{b}=-\frac{1}{2}V_{,\lambda}n^{a}\;\;.$ (12)
@@
 $Q=-\epsilon r^{2}\alpha_{,\lambda}.$ (13)
-The non-zero components of $E_{ab}$ are
-$(E_{ww})$ ::: $0=r_{,ww}-\frac{V(r^{2}V_{,\lambda})_{,\lambda}}{4r}-\frac{\epsilon}{2}\left(r_{,w}V_{,\lambda}-r_{,\lambda}V_{,w}\right)\;,+\frac{1}{2}\kappa r\left[|\Phi_{,w}|^{2}+\frac{VQ^{2}}{8\pi r^{4}}+q^{2}\alpha^{2}|\Phi|^{2}+iq\alpha\Bigl{(}\Phi\bar{\Phi}_{,w}-\bar{\Phi}\Phi_{,w}\Bigr{)}\right]\;,$
-$(E_{w\lambda})\;$ ::: $0=r_{,w\lambda}+\frac{\epsilon}{4r}(r^{2}V_{,\lambda})_{,\lambda}-\frac{1}{2}\kappa r\left[\frac{\epsilon Q^{2}}{8\pi r^{4}}-\frac{1}{2}[\overline{\Phi}_{,w}\Phi_{,\lambda}+\overline{\Phi}_{,\lambda}\Phi_{,w}]-\frac{iq\alpha}{2}(\Phi\overline{\Phi}_{,\lambda}-\overline{\Phi}\Phi_{,\lambda})\right]\;,$
+The nonzero components of $E_{ab}$ are
+$(E_{ww})$ ::: $0=r_{,ww}-\frac{V(r^{2}V_{,\lambda})_{,\lambda}}{4r}-\frac{\epsilon}{2}\left(r_{,w}V_{,\lambda}-r_{,\lambda}V_{,w}\right)+\frac{1}{2}\kappa r\left[|\Phi_{,w}|^{2}+\frac{VQ^{2}}{8\pi r^{4}}+q^{2}\alpha^{2}|\Phi|^{2}+iq\alpha\Bigl{(}\Phi\bar{\Phi}_{,w}-\bar{\Phi}\Phi_{,w}\Bigr{)}\right]\;,$
+$(E_{w\lambda})\;$ ::: $0=r_{,w\lambda}+\frac{\epsilon}{4r}(r^{2}V_{,\lambda})_{,\lambda}-\frac{1}{2}\kappa r\left[\frac{\epsilon Q^{2}}{8\pi r^{4}}-\frac{1}{2}\left(\overline{\Phi}_{,w}\Phi_{,\lambda}+\overline{\Phi}_{,\lambda}\Phi_{,w}\right)-\frac{iq\alpha}{2}(\Phi\overline{\Phi}_{,\lambda}-\overline{\Phi}\Phi_{,\lambda})\right]$ (15)
 $(E_{\lambda\lambda})\;$ ::: $0=r_{,\lambda\lambda}+\frac{1}{2}\kappa r|\Phi_{,\lambda}|^{2}\;,$ (16)
-$(q^{AB}E_{AB})\;$ ::: $0=-\left(Vrr_{,\lambda}+2\epsilon rr_{,w}\right)_{,\lambda}+1-\frac{\kappa}{8\pi}\frac{Q^{2}}{r^{2}}$ (17)
-and those of $E_{a}$
+$(q^{AB}E_{AB})\;$ ::: $0=-\left(Vrr_{,\lambda}+2\epsilon rr_{,w}-\lambda\right)_{,\lambda}-\frac{\kappa}{8\pi}\frac{Q^{2}}{r^{2}},$ (17)
+and those of $E_{a}$ are
 $(E_{w})\;$ ::: $0=Q_{,w}+\epsilon VQ_{,\lambda}+4\pi q^{2}r^{2}\alpha|\Phi|^{2}+2i\pi qr^{2}\Bigl{(}\Phi\bar{\Phi}_{,w}-\bar{\Phi}\Phi_{,w}\Bigr{)}\;,$ (18)
@@
 $0=(r^{2}\Phi_{,w})_{,\lambda}+(r^{2}\Phi_{,\lambda})_{,w}+\epsilon(r^{2}V\Phi_{,\lambda})_{,\lambda}+2iqr\alpha(r\Phi)_{,\lambda}-i\epsilon q\Phi Q\;.$ (20)
-In spherical symmetry the Misner-Sharp mass Misner and Sharp (1964)
-$M=\frac{r}{2}(1-g^{ab}r_{,a}r_{b})=\frac{r}{2}\Bigl{(}1-2\epsilon r_{,w}r_{,\lambda}-Vr_{,\lambda}^{2}\Bigr{)}$ (21)
-is an invariant concept of quasilocal mass. Assuming asymptotically flat spacetimes, its asymptotic limit
-$m_{B}:=\lim_{\lambda\rightarrow\infty}M$ (22)
+In spherical symmetry the Misner-Sharp mass Misner and Sharp (1964),
+$M=\frac{r}{2}(1-g^{ab}r_{,a}r_{b})=\frac{r}{2}\Bigl{(}1-2\epsilon r_{,w}r_{,\lambda}-Vr_{,\lambda}^{2}\Bigr{)},$ (21)
+is an invariant concept of quasilocal mass. Assuming asymptotically flat spacetimes, its asymptotic limit,
+$m_{B}:=\lim_{\lambda\rightarrow\infty}M,$ (22)
 is the Bondi mass $m_{B}$.
@@
 III The characterisitic initial-boundary value problem
-The formulation of the an initial boundary value problem for (II) - (20) is a consequence of the twice contracted Bianchi identities, $\nabla^{a}(E_{ab}-\frac{1}{2}g_{ab}(g^{cd}E_{cd}))$ and its electromagnetic counterpart $\nabla^{a}E_{a}$ Tamburino and Winicour (1966) and it follows the idea of Bondi et al.
-(1962); Sachs (1962a, b).
-The spirit is to pick from $(E_{ab},E_{a},E)$ a set of main equations to hold everywhere, then the twice contracted Bianchi identities and this set of main equation imply that the complement equations of the main equations either hold trivially as a consequence of the main equations or they must be satisfied at a given value of $\lambda=\lambda_{\mathcal{B}}=const$.
-The former is called trivial equation and the latter are the so-called supplementary equations.
+The formulation of the CIBVP for (II) - (20) is a consequence of the twice contracted Bianchi identities, $\nabla^{a}(E_{ab}-\frac{1}{2}g_{ab}(g^{cd}E_{cd}))$ and its electromagnetic counterpart $\nabla^{a}E_{a}$ and it follows the idea of Bondi et al.
+(1962); Sachs (1962a, b); Tamburino and Winicour (1966).
+The spirit444 See e.g.
+Sachs (1962b); Tamburino and Winicour (1966) or Appendix A of Mädler (2019) for a discussion on the twice contracted Bianchi identity for a general metric adopted to a family of null hypersurfaces and the classification of main equations, supplementary equations and trivial equation.
+is to pick from $(E_{ab},E_{a},E)$ a set of main equations to hold everywhere, and then the twice contracted Bianchi identities and this set of main equations imply that the complementary equations of the main equations either hold trivially as a consequence of the main equations or they must be satisfied at a given value of $\lambda=\lambda_{\mathcal{B}}$=const.
+The former is called the trivial equation and the latter are the so-called supplementary equations.
 For the set (II) - (20), the following assignations are made:
@@
     $E_{\lambda\lambda}\;,\;\;q^{AB}E_{AB}\;,\;\;E_{\lambda}\;,\;\;E\;,$ (25)
-    which given by (16), (17), (19) and (20), respectively.
+    which are given by (16), (17), (19), and (20), respectively.
   • supplementary equations :
-    $E_{ww}\;,\;\;E_{w}$ (26)
+    $E_{ww}\;,\;\;E_{w},$ (26)
     given by (II) and (18), respectively
   • trivial equation :
-    $E_{w\lambda}$ (27)
-    represented by (LABEL:eq:E01).
-The main equations (25) do not form a typical hierarchy for the field equations like in other characteristic formulations, see e.g.
-Bondi et al.
-(1962); Newman and Penrose (1962); Mädler and Winicour (2016).
-This is due to the appearence of the $r_{,w}$ derivative in (17).
-Definition of the following new fields
+    $E_{w\lambda},$ (27)
+    represented by (15).
+In the Bondi-Sachs formulation of the CIBVP Bondi et al.
+(1962); Sachs (1962a); Mädler and Winicour (2016) the main equations form a natural hierarchical system of hypersurface equations that can be solved provided there is an initial profile of the dynamically propagating field(s).
+That is, each of the equations is solved one after another, in successive order.
+For the Einstein-scalar field equations in spherical symmetry while equipped with Bondi-Sachs coordinates, the only field propagating is the scalar field.
+There are two remaining fields in the system parametrizing the metric, the fields $\beta_{BS}$ and $V_{BS}$ (see, e.g., Gomez et al.
+(1994); Mädler and Gallo (2025)).
+Indeed, in that case [where $E_{\lambda}$ in (25) does not need to be considered], $E_{\lambda\lambda}$ is integrated first with the initial data to find $\beta_{BS}$.
+Then, with the initial data and $\beta_{BS}$, the field $V_{BS}$ is found by the integration of $q^{AB}E_{AB}$ in (25).
+Next, the wave equation $E$ in (25) is used to update the initial data from the initial null hypersurface to a null hypersurface at later times.
+However, this beautiful and beneficial hierarchy for numerical integration is destroyed in the affine-null formulation, due to the appearance of the $r_{,w}$ derivative in (17).
+This additional time derivative is a consequence of the affine-null coordinate choice.
+The affine parameter $\lambda$ is related to the Bondi-Sachs area distance coordinate $r$ via $\lambda_{,r}(w,r,x^{A})=\exp[2\beta_{BS}(w,r,x^{A})]$.
+The transformation appears simple, but it actually comes along with a change in the $\partial_{w}$ directions555 The $\partial_{A}$-directions change in addition, if the assumption of spherical symmetry is dropped..
+This results in the additional $w$-derivative in (17) Winicour (2013).
+Nevertheless, we can establish the hierarchy in the affine-null case after the introduction of the following new fields
 $\rho$ $=$ $r_{,w}\;\;,\;\;$ (28a)
 $Z$ $=$ $Vrr_{,\lambda}+2\epsilon r\rho-\lambda\;\;,\;\;$ (28b)
-$\mathcal{L}$ $=$ $\frac{2r\left(r_{,\lambda}\Phi_{,w}-\rho\Phi_{,\lambda}\right)+\epsilon(Z+\lambda)\Phi_{,\lambda}}{r_{,\lambda}}$ (28c)
-cast (25) into
+$\mathcal{L}$ $=$ $\frac{2r\left(r_{,\lambda}\Phi_{,w}-\rho\Phi_{,\lambda}\right)+\epsilon(Z+\lambda)\Phi_{,\lambda}}{r_{,\lambda}}.$ (28c)
+which cast (25) into
 00 0 $=$ $r_{,\lambda\lambda}+\frac{1}{2}\kappa r|\Phi_{,\lambda}|^{2}\;,$ (29a)
@@
 $\Phi_{,w}$ $=$ $\frac{\mathcal{L}}{2r}-\frac{\epsilon}{2}V\Phi_{,\lambda}\;,$ (29g)
-where we introduced the auxillary variable
+where we introduced the auxiliary variable
 $\mathcal{J}=i(\overline{\Phi}\Phi_{,\lambda}-\overline{\Phi}_{,\lambda}\Phi).$ (29h)
@@
 A solution of (29) requires initial data
-$\Phi(0,\lambda)=\mathcal{F}_{\Phi}(\lambda)$ (30)
+$\Phi(0,\lambda)=\mathcal{F}_{\Phi}(\lambda),$ (30)
 on a null hypersurface $w=0$ for values $\lambda\geq\lambda_{\mathcal{B}}$, where $\mathcal{F}_{\Phi}(\lambda)$ is an arbitrary function.
-It further requires the following boundary conditions at $\lambda=\lambda_{\mathcal{B}}$ for all values $w>0$
+It further requires the following boundary conditions at $\lambda=\lambda_{\mathcal{B}}$ for all values $w\neq 0$
 $\begin{split}&r(w,\lambda_{\mathcal{B}})\;,\;\;r_{,\lambda}(w,\lambda_{\mathcal{B}})\;,\;\;Q(w,\lambda_{\mathcal{B}})\;,\;\;\alpha(w,\lambda_{\mathcal{B}})\;,\;\;\\ &\mathcal{L}(w,\lambda_{\mathcal{B}})\;,\;\;V(w,\lambda_{\mathcal{B}})\;,\;\;V_{,\lambda}(w,\lambda_{\mathcal{B}})\;.\end{split}$ (31)
-In particular, the boundary conditions for the variables $r$, $Q$, $\alpha$ and $V$ follow from physical considerations, while those for auxillary fields $Z$ and $\mathcal{L}$ they follow from the definitions (28) and are given by
+In particular, the boundary conditions for the variables $r$, $Q$, $\alpha$, and $V$ follow from physical considerations, while those for auxiliary fields $Z$ and $\mathcal{L}$ follow from the definitions (28) and are given by
 $Z|_{\lambda=\lambda_{\mathcal{B}}}$ $=$ $(Vrr_{,\lambda}+2\epsilon rr_{,w}-\lambda)|_{\lambda=\lambda_{\mathcal{B}}}\;\;,\;\;$ (32a)
 $\mathcal{L}|_{\lambda=\lambda_{\mathcal{B}}}$ $=$ $\left(2r\Phi_{,w}+\epsilon rV\Phi_{,\lambda}\right)|_{\lambda=\lambda_{\mathcal{B}}}\;\;.$ (32b)
-It seems “unpleasant” that the boundary conditions (32) for $Z$ and $\mathcal{L}$, contain $w$ derivatives of the data $\Phi$ as well as the conformal factor $r$ at the boundary $\lambda=\lambda_{\mathcal{B}}$.
-We will see in the next sections that this can be (partially) resolved by specifying the nature of the boundary hypersurface at $\lambda_{\mathcal{B}}$.
+The $r_{,w}$ derivative in the boundary condition for $Z$ in (32a) can be either removed or evaluated from $\Phi_{,w}$ via a differential equation in the formulation of the CIBVP (see next sections).
+Thus, the boundary conditions depend only on the $w$-derivative of the propagating physical field i.e., the derivative $\Phi_{,w}$.
+Consequently, the only propagating degree of freedom of the formulated CIBVP is the field that is subject to a hyperbolic equation, whose characteristics are given by the $w=$const null hypersurfaces.
+For the Einstein-Maxwell-scalar-field equations in spherical symmetry only the scalar field is determined by a hyperbolic partial differential equation, i.e., the wave equation $\mathcal{D}_{a}\mathcal{D}^{a}\Phi=0.$
 Since we can always translate the origin of an affine parameter by a transformation like $\lambda\rightarrow a(w)\lambda+b(w)$, we set hereafter $\lambda_{\mathcal{B}}=0$.
-In the following three sections we discuss the asymptotic initial value problem, the initial value problem at a world line and the double null initital boundary value problem for two intersecting null hypersurfaces.
+In the following three sections we discuss the asymptotic initial value problem, the initial value problem at a world line, and the double-null initial boundary value problem for two intersecting null hypersurfaces.
 III.1 Asymptotic initial value problem for large values of the affine parameter
-Let the scalar field have the following asymptotic expansion
+To consider the asymptotic characteristic initial value problem, we assume that the fields have sufficient differentiability for large values of the affine parameter and in the limit $\lambda\rightarrow\infty$.
+An illustration of the null hypersurfaces for the parameter $\epsilon$ is given in Fig.
+1.
+The scalar field is assumed to obey the radiation condition that the limit $\lim_{\lambda\rightarrow\infty}\lambda\Phi$ is finite on $w=$const null hypersurfaces.
+Figure 1: The asymptotic null hypersurface for the cases $\epsilon=\pm 1$ are shown by spherical cuts at the location of “$\lambda=\infty$”.
+Either the idealized limit “$\lambda=\infty$” can be attached to the physical spacetime via the conformal Penrose compactification Penrose (1963) or it can be achieved by introducing a radial compactified coordinate, such as $\ell=1/\lambda$, and a proper regularization of the field equationMädler and Gallo (2025).
+The upper graphic illustrates a series of spherical cross sections of a Bondi frame, where the affine parameter $\lambda\rightarrow\infty$ on outgoing null hypersurfaces $w_{B}$=const.
+The Bondi time $w_{B}$ then corresponds to the retarded time $u_{B}$.
+The lower graphic shows the corresponding situation for ingoing null hypersurfaces toward a Bondi frame, where $\epsilon=+1$ and $w_{B}$ are ingoing null hypersurfaces.
+In this case, the Bondi time is the advanced coordinate $v_{B}$.
+In both graphics, coordinates $x^{A}$ parametrize the cuts at $w_{B}$=const and label the different rays (indicated by the arrow lines also carrying $\lambda\rightarrow\infty$ ) reaching the “$\lambda=\infty$” (null) hypersurface .
+Then the scalar field has the following asymptotic expansion
 $\Phi=\frac{\varphi_{[1]}+i\psi_{[1]}}{\lambda}+\frac{\varphi_{[2]}+i\psi_{[2]}}{\lambda^{2}}+O(\lambda^{-3})\;,$ (33)
-where the functions with subscipt indices in brackets labelling the order of $\lambda^{-i}$ are real functions depending on $w$.
-The function $\mathcal{L}$ is also decomposed into a real and imaginary part
-$\mathcal{L}=L+iP$ (34)
+where the functions with subscript indices in brackets labeling the order of $\lambda^{-i}$ are real functions depending on $w$.
+The function $\mathcal{L}$ is also decomposed into real and imaginary parts
+$\mathcal{L}=L+iP,$ (34)
 with $L$ and $P$ also being real functions. Insertion of (33) into (29h) yields
-$\mathcal{J}=\frac{2(\varphi_{[1]}\psi_{[2]}-\varphi_{[2]}\psi_{[1]})}{\lambda^{4}}+O(\lambda^{-5})$ (35)
+$\mathcal{J}=\frac{2(\varphi_{[1]}\psi_{[2]}-\varphi_{[2]}\psi_{[1]})}{\lambda^{4}}+O(\lambda^{-5}),$ (35)
 and subsequent integration of (29) gives the asymptotic solution
@@
 $V(w,\lambda)=$ $U_{\infty}\lambda+V_{\infty}+\frac{2(HZ_{\infty}-R_{\infty})+\epsilon\kappa H^{2}(L_{\infty}\varphi_{[1]}+P_{\infty}\psi_{[1]})}{2H^{3}\lambda}+O(\lambda^{-2})\;,$ (36g)
-where $H$, $R_{\infty}$, $Q_{\infty}$, $\alpha_{\infty}$, $Z_{\infty}$, $L_{\infty}$, $P_{\infty}$, $U_{\infty}$ and $V_{\infty}$ are functions of integrations depending on $w$.
-Using (17), we relate the functions integration $U_{\infty}$ and $V_{\infty}$
+where $H$, $R_{\infty}$, $Q_{\infty}$, $\alpha_{\infty}$, $Z_{\infty}$, $L_{\infty}$, $P_{\infty}$, $U_{\infty}$ and $V_{\infty}$ are functions of integration depending on $w$.
+Using (17), we relate the functions integration $U_{\infty}$ and $V_{\infty}$,
 $U_{\infty}$ $=$ $-\frac{2\epsilon H^{\prime}}{H}\;,$ (37)
-$V_{\infty}$ $=$ $\frac{1-2\epsilon HR_{\infty}^{\prime}}{H^{2}}\;\;.$ (38)
+$V_{\infty}$ $=$ $\frac{1-2\epsilon HR_{\infty}^{\prime}}{H^{2}}\;\;,$ (38)
+where denotes the total derivative with respect to the single independent argument of the respective function.
 The evolution equations (29g) relate the constants $L_{\infty}$ and $P_{\infty}$ to the $w$ derivatives of $\varphi_{[1]}$ and $\psi_{[1]}$, respectively,
@@
 $m_{B}=\frac{1}{2}(R_{\infty}-HZ_{\infty})\;.$ (41)
-From this result, alternative calculation of the Bondi mass using the hypersurface variables is
+From this result, an alternative calculation of the Bondi mass using the hypersurface variables is
 $m_{B}=-\frac{1}{2}\lim_{\lambda\rightarrow\infty}\left[\lambda^{2}\left(\frac{r}{\lambda}\right)_{,\lambda}+HZ\right]\;\;.$ (42)
@@
 The asymptotic transformation
-$x(w,\lambda)=H(w)\lambda+R_{\infty}(w)+O(\lambda^{-1})$ (44)
+$x(w,\lambda)=H(w)\lambda+R_{\infty}(w)+O(\lambda^{-1}),$ (44)
 with $w$ and the angles kept unaltered casts (43) into
@@
 which is the metric of an inertial observer for large values of the affine parameter.
-The coordinate frame in which $H=1$ is called Bondi frame and $w_{B}$ referred to as Bondi time corresponding to the inertial time of this observer.
-Consequently, in a regime where $H\neq 0$, it measures the redshift between asymptotic inertial observers and observes at finite affine parameter distance.
+The coordinate frame in which $H=1$ is called the Bondi frame and $w_{B}$ is referred to as the Bondi time, corresponding to the inertial time of this observer.
+Consequently, in a regime where $H\neq 0$, it measures the redshift between asymptotic inertial observers and observers at finite affine parameter distance.
 In particular, if $H\neq 0$, we can always find a coordinate transformation of the type (46) such that $H=1$.
-If, however, $H=0$ for some time $w=w_{E}$, then $w_{E}$ corresponds to the time of formation of an event horizon, since the surface forming rays of the surface $dw=const$ do not have an endpoint in the asymptotic region.
+If, however, $H=0$ for some time $w=w_{E}$, then $w_{E}$ corresponds to the time of formation of an event horizon, since the surface forming rays of the hypersurface $w=$const do not have an end point in the asymptotic region.
 The redshift can also conveniently be calculated via
-$H=\lim_{\lambda\rightarrow\infty}r_{,\lambda}$ (48)
+$H=\lim_{\lambda\rightarrow\infty}r_{,\lambda}.$ (48)
 In a Bondi frame, the metric fields behave as
-$\begin{split}r&=\lambda+R_{\infty}(w_{B})+O(\lambda^{-1})\\ V&=1-2\epsilon R_{\infty}^{\prime}(w_{B})+O(\lambda^{-2})\\ \end{split}$ (49)
-and the scalar field $\Phi$ has the expansion,
-$\Phi=\frac{\mathcal{C}(w_{B})}{r}+\frac{\mathcal{C}_{NP}}{r^{2}}+O(r^{-3})$ (50)
+$\begin{split}r&=\lambda+R_{\infty}(w_{B})+O(\lambda^{-1}),\\ V&=1-2\epsilon R_{\infty}^{\prime}(w_{B})+O(\lambda^{-2}),\\ \end{split}$ (49)
+and the scalar field $\Phi$ admits the expansion,
+$\Phi=\frac{\mathcal{C}(w_{B})}{r}+\frac{\mathcal{C}_{NP}}{r^{2}}+O(r^{-3}),$ (50)
 where $\mathcal{C}(w_{B})=\mathcal{C}_{\varphi}(w_{B})+i\mathcal{C}_{\psi}(w_{B})$ is the scalar monopole and $\mathcal{C}_{NP}$ is the conserved Newman-Penrose constant.
-The two can alternatively be determined via the limits
-$\mathcal{C}$ $=$ $\lim_{r\rightarrow\infty}r\Phi=\lim_{\lambda\rightarrow\infty}H\lambda\Phi=H(\varphi_{[1]}+i\psi_{[1]})$ (51)
+These two quantities can alternatively be determined via the limits
+$\mathcal{C}$ $=$ $\lim_{r\rightarrow\infty}r\Phi=\lim_{\lambda\rightarrow\infty}H\lambda\Phi$ (51)
+$=$ $H(\varphi_{[1]}+i\psi_{[1]}),$
 $\mathcal{C}_{NP}$ $=$ $-\!\lim_{r\rightarrow\infty}r^{2}(r\Phi)_{,r}=-\!\lim_{\lambda\rightarrow\infty}\frac{r^{2}(r\Phi)_{,\lambda}}{r_{,\lambda}}$ (52)
-$=$ $R_{\infty}\mathcal{C}+H^{2}(\varphi_{[2]}+i\psi_{[2]})$
+$=$ $R_{\infty}\mathcal{C}+H^{2}(\varphi_{[2]}+i\psi_{[2]}).$
 The derivative
-$\mathscr{N}(w_{B}):=\frac{d\mathcal{C}}{dw_{B}}\;\;$ (53)
+$\mathcal{N}(w_{B}):=\frac{d\mathcal{C}}{dw_{B}},\;\;$ (53)
 is the news function of the scalar field, whose naming will become justified below.
-The relation between the Bondi time $w_{B}$, and the asymptotic time $w$ is given by asymptotic transformation (46) therefore
-$\mathcal{N}(w_{B}):=\frac{d\mathcal{C}}{dw}\frac{dw}{dw_{B}}=H\frac{d\mathcal{C}}{dw}$ (54)
+The relation between the Bondi time $w_{B}$, and the asymptotic time $w$ is given by asymptotic transformation (46); therefore,
+$\mathcal{N}(w_{B}):=\frac{d\mathcal{C}}{dw}\frac{dw}{dw_{B}}=H\frac{d\mathcal{C}}{dw}.$ (54)
 The asymptotic limit of the Weyl scalar $\Psi_{2}$ gives
-$\begin{split}\lim_{\lambda\rightarrow\infty}\lambda^{3}\Psi_{2}=&\frac{2HZ_{\infty}-R_{\infty}}{2H^{3}}\\ &+\frac{\kappa\epsilon}{6H^{2}}\left[H\varphi_{[1]}\frac{d}{dw}H\varphi_{[1]}+H\psi_{[1]}\frac{d}{dw}H\psi_{[1]}\right]\end{split}$ (55)
-in a Bondi frame ($H=1$, $r=\lambda$), we can deduce that
+$\begin{split}\lim_{\lambda\rightarrow\infty}\lambda^{3}\Psi_{2}=&\frac{2HZ_{\infty}-R_{\infty}}{2H^{3}}\\ &+\frac{\kappa\epsilon}{6H^{2}}\left[H\varphi_{[1]}\frac{d}{dw}H\varphi_{[1]}+H\psi_{[1]}\frac{d}{dw}H\psi_{[1]}\right],\end{split}$ (55)
+in a Bondi frame ($H=1$, $r=\lambda$), and we can deduce that
 $\begin{split}\lim_{r\rightarrow\infty}\Psi_{2}=&-m_{B}(w_{B})\\ &+\frac{1}{12}\kappa\epsilon\left[\mathcal{C}(w_{B})\bar{\mathcal{N}}(w_{B})+\bar{\mathcal{C}}(w_{B})\mathcal{N}(w_{B})\right]\;,\end{split}$ (56)
@@
 The expression (57) corresponds to an expression presented in Bicak et al. (2010). Evaluation of the supplementary equations (18) and (II) in a Bondi frame ($H=1$) gives us the asymptotic balance laws
-$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}\left\{\left(\frac{d\varphi_{[1]}}{dw_{B}}\right)^{2}+\left(\frac{d\psi_{[1]}}{dw_{B}}\right)^{2}+2q\alpha_{\infty}\left[\phi_{[1]}\left(\frac{d\psi_{[1]}}{dw_{B}}\right)-\psi_{[1]}\left(\frac{d\phi_{[1]}}{dw_{B}}\right)\right]+q^{2}\alpha_{\infty}^{2}\left(\phi_{[1]}^{2}+\psi_{[1]}^{2}\right)\right\}$ (58a)
-$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-4\pi q^{2}\alpha_{\infty}^{2}\left(\phi_{[1]}^{2}+\psi_{[1]}^{2}\right)+4\pi q\left[\phi_{[1]}\left(\frac{d\psi_{[1]}}{dw_{B}}\right)-\psi_{[1]}\left(\frac{d\phi_{[1]}}{dw_{B}}\right)\right]$ (58b)
-which we can also write a in terms of the complex news function and complex scalar monopole
-$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}\left|\mathcal{N}+iq\alpha_{\infty}\mathcal{C}\right|^{2}$ (59a)
-$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-4\pi q^{2}\alpha_{\infty}^{2}|\mathcal{C}|^{2}-2i\pi q\left(\mathcal{N}\bar{\mathcal{C}}-\bar{\mathcal{N}}\mathcal{C}\right)$ (59b)
-As mentioned Sec.
+$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}\left\{\left(\frac{d\varphi_{[1]}}{dw_{B}}\right)^{2}+\left(\frac{d\psi_{[1]}}{dw_{B}}\right)^{2}+2q\alpha_{\infty}\left[\phi_{[1]}\left(\frac{d\psi_{[1]}}{dw_{B}}\right)-\psi_{[1]}\left(\frac{d\phi_{[1]}}{dw_{B}}\right)\right]+q^{2}\alpha_{\infty}^{2}\left(\phi_{[1]}^{2}+\psi_{[1]}^{2}\right)\right\},$ (58a)
+$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-4\pi q^{2}\alpha_{\infty}^{2}\left(\phi_{[1]}^{2}+\psi_{[1]}^{2}\right)+4\pi q\left[\phi_{[1]}\left(\frac{d\psi_{[1]}}{dw_{B}}\right)-\psi_{[1]}\left(\frac{d\phi_{[1]}}{dw_{B}}\right)\right],$ (58b)
+which we can write also in terms of the complex news function and complex scalar monopole
+$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}\left|\mathcal{N}+iq\alpha_{\infty}\mathcal{C}\right|^{2},$ (59a)
+$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-4\pi q^{2}\alpha_{\infty}^{2}|\mathcal{C}|^{2}-2i\pi q\left(\mathcal{N}\bar{\mathcal{C}}-\bar{\mathcal{N}}\mathcal{C}\right).$ (59b)
+As mentioned in Sec.
 II, there is a gauge transformation to set [^6] Since $\alpha_{\infty}=\alpha_{\infty}(w_{B})$ this gauge transformation needs to be done on each null hypersurface $w_{B}=$const for every value of $w_{B}$.
 $\alpha_{\infty}=0$.
-Using $\alpha_{\infty}=0$ in (59) casts the the Bondi mass conservation law and asymptotic the charge conservation law into the simple expressions
-$\frac{dm_{B}}{dw_{B}}=\frac{\epsilon\kappa}{2}|\mathcal{N}|^{2}\;\;,\;\;\frac{dQ_{\infty}}{dw_{B}}=-2i\pi q\left(\mathcal{N}\bar{\mathcal{C}}-\bar{\mathcal{N}}\mathcal{C}\right)\;\;$ (60)
+Using $\alpha_{\infty}=0$ in (59) casts the Bondi mass conservation law and the asymptotic charge conservation law into the simple expressions
+$\frac{dm_{B}}{dw_{B}}=\frac{\epsilon\kappa}{2}|\mathcal{N}|^{2}\;\;,\;\;\frac{dQ_{\infty}}{dw_{B}}=-2i\pi q\left(\mathcal{N}\bar{\mathcal{C}}-\bar{\mathcal{N}}\mathcal{C}\right)\;\;.$ (60)
 From the definition of the news function $\mathcal{N}$ in (53) as the derivative of the scalar monopole $\mathcal{C}$ and (60), we deduce that the asymptotic initial value problem for the charged version ($q\neq 0$ ) requires specification of three scalar values and one arbitrary function depending on $w_{B}$ while the uncharged version $(q=0)$ only requires specification of two scalars and one function depending on $w_{B}$.
-In both of these cases, the free function is the news function $\mathcal{N}$.
-Its name arises from the same motivation as in the gravitational case, see (Bondi et al., 1962, Sec.
-3), the function $\mathcal{N}$ determines the complete (temporal) behaviour of all the other physically measurable fields, $\mathcal{C}(w_{B})$, $m_{B}(w_{B})$ and $Q_{\infty}(w_{B})$.
-Both of the two versions of these asymptotic initial value problems can be discussed together.
-Given a scalar news function $\mathcal{N}$ for $w_{B}\geq 0$.
-The three scalars are specified at the initial time $w_{B}=0$.
+In both cases, the free function is the news function $\mathcal{N}$.
+Its name arises from the same motivation as in the gravitational case ( see (Bondi et al., 1962, Sec.
+3)), and the function $\mathcal{N}$ determines the complete (temporal) behavior of all the other physically measurable fields, $\mathcal{C}(w_{B})$, $m_{B}(w_{B})$ and $Q_{\infty}(w_{B})$.
+Both of the two versions of these asymptotic initial value problems can be treated together.
+Given a scalar news function $\mathcal{N}$ for $w_{B}\neq 0$, the three scalars are specified at the initial time $w_{B}=0$.
 They are $\mathcal{C}(0)$, $m_{B}(0)$ and $Q_{\infty}(0)$, which are the initial scalar monopole, the initial Bondi mass and the initial value of the global charge, respectively.
-The behavior of these fields for $w_{B}>0$ is then determined by the following hierarchy of ordinary differential equations of the asymptotic initial value problem
+The behavior of these fields for $w_{B}\neq 0$ is then determined by the following hierarchy of ordinary differential equations of the asymptotic initial value problem
 $\frac{d\mathcal{C}}{dw_{B}}$ $=$ $\mathcal{N}(w_{B})\;\;,$ (61a)
 $\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}|\mathcal{N}(w_{B})|^{2}\;\;,$ (61b)
-$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-2i\pi q\left[\mathcal{N}(w_{B})\bar{\mathcal{C}}(w_{B})-\bar{\mathcal{N}}(w_{B})\mathcal{C}(w_{B})\right]\;\;.$ (61c)
+$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-2i\pi q\left[\mathcal{N}(w_{B})\bar{\mathcal{C}}(w_{B})-\bar{\mathcal{N}}(w_{B})\mathcal{C}(w_{B})\right]\;\;.$
 In the uncharged version, where the gauge charge $q$ vanishes, the last equation of (61) implies $Q_{\infty}(w_{B})=0$ as the initial value of the global charge $Q_{\infty}(0)$ is obviously zero.
-Thus only two initial values and one free function is required for the asymptotic initial value problem in this case.
+Thus only two initial values and one free function are required for the asymptotic initial value problem in this case.
 III.2 Local initial-boundary formulation at a world line
 In spherical symmetry there exists a designated world line, which is the geodesic tracing the centers of spherical symmetry Synge (1960).
-We denote this geodesic with $c(w)$ and choose $w$ as an affine parameter for its parameterisation.
+We denote this geodesic with $c(w)$ and choose $w$ as an affine parameter for its parameterization.
 The tangent vector $c^{a}\partial_{a}$ is a unit timelike vector and $w$ measures the proper time along $c(w)$.
 This proper time is assumed to be the proper time of a freely falling Fermi observer along $c(w)$ Manasse and Misner (1963); Poisson (2009).
-Following Madler and Muller (2013), in a convex normal neighborhood of $c(w)$ we extend the definition of $w$ be constant along null cones emanating from $c(w)$.
+Following Madler and Muller (2013), in a convex normal neighborhood of $c(w)$, we extend the definition of $w$ to be constant along null cones emanating from $c(w)$.
 The surface forming rays of the $w=$const null cones are labeled by the angular coordinates $x^{A}$ within each null cone.
 The null condition on $w=$const implies $g^{ww}=0$, since $g^{ab}\nabla_{a}w\nabla_{b}w=0$.
-The co-vector $\ell_{a}=\epsilon\nabla_{a}w$ defines by metric duality the tangent vectors $\ell^{a}=\epsilon g^{ab}\nabla_{b}w=\epsilon g^{aw}$ of the surface forming null rays.
+The covector $\ell_{a}=\epsilon\nabla_{a}w$ defines by metric duality the tangent vectors $\ell^{a}=\epsilon g^{ab}\nabla_{b}w=\epsilon g^{aw}$ of the surface forming null rays.
 The null rays are labeled by the angular coordinates, $\ell^{a}\nabla_{a}x^{A}=0$ and therefore $g^{wA}=0$.
@@
 We normalize the parameter $\lambda$ along the rays such that it is an affine parameter with the normalization determined at $c(w)$.
-In particular, $c^{a}\nabla_{a}\lambda=\ell^{a}\nabla_{a}\lambda=\epsilon g^{w\lambda}=1$ with $|\epsilon|=1$ at $c(w)$ so that $\epsilon=-1$ designates that the respective null cone $w=const$ are outgoing null cones with null rays pointing into the future of $c(w)$ and $\epsilon=1$ are ingoing null cones pointing into the past of $c(w)$.
-By these conditions, the metric near $c(w)$ has the form [^4] RefMadler and Muller (2013) uses a reversed nomenclature for the binary parameter $\epsilon$. Madler and Muller (2013)
+In particular, $c^{a}\nabla_{a}\lambda=\ell^{a}\nabla_{a}\lambda=\epsilon g^{w\lambda}=1$ with $|\epsilon|=1$ at $c(w)$ so that $\epsilon=-1$ designates that the respective null cone $w=$const are outgoing null cones with null rays pointing into the future of $c(w)$ and $\epsilon=1$ are ingoing null cones pointing into the past of $c(w)$ (see Fig.
+2).
+Figure 2: A family of null cones $\mathscr{N}^{\epsilon}(w)$ at the central geodesic $c(w)$ is shown for $\epsilon=-1$ (left) and $\epsilon=+1$ (right).
+The left graphic shows the local geometry for the characteristic initial value problem where initial data are provided in the null cone $\mathscr{N}^{-}(0)$ and the right graphic illustrates it for the characteristic past value problem with the future data on $\mathscr{N}^{+}(0)$.
+The origin of the affine parameter $\lambda=0$ for each null cone $\mathscr{N}^{-}(w)$ is at the vertices of the cones traced by $c(w)$.
+Distinctive vertices are labeled with different values of $w$.
+By these conditions, the metric near $c(w)$ has the form Madler and Muller (2013) [^7] Reference Madler and Muller (2013) uses a reversed nomenclature for the binary parameter $\epsilon$.
 $\begin{split}ds^{2}=&-\left[1+O(\lambda^{2})\right](dw)^{2}+2\epsilon dwd\lambda\\ &+\lambda^{2}q_{AB}dx^{A}dx^{B}+O(\lambda^{3}).\end{split}$ (62)
@@
 $Z(w,0)=\mathcal{L}(w,0)=0\;\;,\;\;$ (64)
-The supplementary equation (II) holds at $c(w)$ due to the regularity conditions. Eq. (63) and (18) imply
-$Q_{,w}=O(\lambda^{2})\;\;\Rightarrow\;\;Q(w,0)=Q_{,\lambda}(w,0)=0$ (65)
+The supplementary equation (II) holds at $c(w)$ due to the regularity conditions. Equations (63) and (18) imply
+$Q_{,w}=O(\lambda^{2})\;\;\Rightarrow\;\;Q(w,0)=Q_{,\lambda}(w,0)=0.$ (65)
 At the vertex, we assume that the scalar field $\Phi$ has the regular Taylor expansion
 $\Phi=\Phi_{(0)}+\Phi_{(1)}\lambda+\frac{1}{2}\Phi_{(2)}\lambda^{2}+O(\lambda^{3})\;\;,$ (66)
-where the subscript $(k)$ in parenthesis indicate that the symbol carrying the subindex is the $\lambda^{k}$ coefficient function depending on $w$ of the Taylor series of the respective field evaluated at $\lambda=0$.
-The scalar field expansion (66) and (29h) give
+where the subscript $(k)$ in parentheses indicate that the symbol carrying the subindex is the $\lambda^{k}$ coefficient function depending on $w$ of the Taylor series of the respective field evaluated at $\lambda=0$.
+The scalar field expansions (66) and (29h) give
 $\begin{split}\mathcal{J}=&\mathcal{J}_{(0)}+\mathcal{J}_{(1)}\lambda+O(\lambda^{2})\\ =&i\left(\overline{\Phi}_{(1)}\Phi_{(0)}-\overline{\Phi}_{(0)}\Phi_{(1)}\right)+i\left(\overline{\Phi}_{(2)}\Phi_{(0)}-\overline{\Phi}_{(0)}\Phi_{(2)}\right)\lambda\\ &+O(\lambda^{2})\;.\end{split}$ (67)
-Insertion of these data into (29) and integration gives
+Insertion of these data into (29) and integration give
 $r(w,\lambda)$ $=$ $\lambda-\frac{\kappa\lambda^{3}}{12}|\Phi_{(1)}|^{2}-\frac{\kappa\lambda^{4}}{24}\Bigl{(}\Phi_{(2)}\overline{\Phi}_{(1)}+\Phi_{(1)}\overline{\Phi}_{(2)}\Bigr{)}+O(\lambda^{5})\;,$ (68a)
@@
 We can see that the time evolution of the coefficient $\Phi_{(i)}$ along $c(w)$ is driven by a function depending on the expansion coefficients containing all expansion coefficients up to $\Phi_{(i+1)}$.
-This is behavior is a result of the choice of $w$ being constant along a null hypersurface.
+This behavior is a result of the choice of $w$ being constant along a null hypersurface.
 The hierarchical structure of (69) is the scalar field equivalent to the similar structure in the gravitational case Madler and Muller (2013); Mädler (2013) for gravitational waves.
-Since $\Phi_{(1)}=\lim_{\lambda\rightarrow 0}\Phi_{,\lambda}$ and $\Phi_{(2)}=\lim_{\lambda\rightarrow 0}\Phi_{,\lambda\lambda}$, we learn from (68) that the evolution of the initial data $\Phi$ along $c(w)$ and in its neighbourhood is completely determined by the data on a cone $w=const$, provided the data and its derivatives are finite in the limit $\lambda\rightarrow 0$.
-Indeed it is not difficult to derive similar equations like in (69) for $\Phi_{(k)}$ with $k\geq 2$ while using an appropriate higher order expansion than (66) and integrating (29).
-From such an exercise, we can deduce the following local characteristic initial-boundary value problem at the vertices of null cones emanating from a geodesic world line: Suppose, the complex scalar field $\Phi(w,\lambda)$ on a null cone $w=0$ has a regular Taylor series expansion,
+Since $\Phi_{(1)}=\lim_{\lambda\rightarrow 0}\Phi_{,\lambda}$ and $\Phi_{(2)}=\lim_{\lambda\rightarrow 0}\Phi_{,\lambda\lambda}$, we learn from (68) that the evolution of the initial data $\Phi$ along $c(w)$ and in its neighborhood is completely determined by the data on a cone $w=$const, provided the data and its derivatives are finite in the limit $\lambda\rightarrow 0$.
+Indeed, it is not difficult to derive similar equations like in (69) for $\Phi_{(k)}$ with $k\geq 2$, while using an appropriate higher order expansion than (66) and integrating (29).
+From such an exercise, we can deduce the following local CIBVP at the vertices of null cones emanating from a geodesic world line: Suppose, the complex scalar field $\Phi(w,\lambda)$ on a null cone $w=0$ has a regular Taylor series expansion,
 $\Phi(0,\lambda)=\sum^{\infty}_{k=0}\frac{\lambda^{k}}{k!}\frac{\partial^{k}\Phi}{\partial\lambda^{k}}\Bigg{|}_{\lambda=0}\;,$ (71)
-evaluated at the world line, the boundary conditions of the fields are (63) and (64), $\alpha_{0}$ is an arbitrary function of $w$ for $w>0$, and the gauge charge $q$ is provided, then the metric, the electromagnetic potential $\alpha$ and the charge function $Q$ can be completely calculated on the cone $w=0$ using the hypersurface equations (29a)-(29f).
-Their local solution at $\lambda=0$ is of the form (68a), (68b), (68c) or (68f).
-The evolution equations (29g) determine the time derivative of the Taylor coefficients of the initial data, as seen in e.g.
-(69).
-We remark, that this regularity assumption at $w=0$ does not imply that the data necessarily remain regular for $w>0$.
+evaluated at the world line, the boundary conditions of the fields are (63) and (64), $\alpha_{0}$ is an arbitrary function of $w$ for $w\neq 0$, and the gauge charge $q$ is provided; then, the metric, the electromagnetic potential $\alpha$ and the charge function $Q$ can be completely calculated on the cone $w=0$ using the hypersurface equations (29a)-(29f).
+Their local solution at $\lambda=0$ is of the form (68a), (68b), (68c), or (68f).
+The evolution equations (29g) determine the time derivative of the Taylor coefficients of the initial data, as seen in, e.g., (69).
+We remark, that this regularity assumption at $w=0$ does not imply that the data necessarily remain regular for $w\neq 0$.
 If $r\rightarrow 0$ for $\lambda\rightarrow\lambda_{C}>0$ a (massive) caustic forms at finite value for the affine parameter.
-This caustic formation for $w>0$ is in general driven by the value of norm squared, $|\Phi_{,\lambda}|^{2}$, of the derivative of the scalar field.
-An example for caustic initical data for a real scalar field was recently given in Mädler et al.
+This caustic formation for some $w=w_{0}\neq 0$ is in general driven by the value of the norm squared of the derivative of the scalar field, $|\Phi_{,\lambda}|^{2}$.
+An example for caustic initial data was recently given in Mädler et al.
 (2024), where the data were employed in an evolution through an event horizon of a supercritical scalar field.
-The thus formulated initial value problem in spherical symmetry and with a scalar field is an analogue to the gravitational one of Friedrich Friedrich (1986).
+The formulated initial value problem along a world line for a spherically symmetric charged scalar field is an analog to the gravitational CIBVP Friedrich (1986).
 III.3 Initial-boundary value problem at a null hypersurface
 Consider a family of null hypersurfaces $w=$const designated by $\mathscr{N}^{\epsilon}(w)$, hereafter.
-The generators $\ell^{a}\partial_{a}$ of these null hypersurface are affinely parameterized with the affine parameter $\lambda$.
+The generators $\ell^{a}\partial_{a}$ of these null hypersurfaces are affinely parametrized with the affine parameter $\lambda$.
 We choose the origin, $\lambda=0$, of the affine parameter to be at a null hypersurface $\mathcal{B}^{\epsilon}$.
-The surface $\mathcal{B}^{\epsilon}$ is a null boundary for every null hypersurface $w=const$ of $\mathscr{N}^{\epsilon}$ and is the boundary surface for the double null CIBVPs to be presented, see Fig.(1).
-Figure 1: A family of null hypersurfaces $\mathscr{N}^{\epsilon}(w)$ with corresponding boundary null hypersurface $\mathcal{B}^{\epsilon}(w)$ is shown for $\epsilon=-1$ (left) and $\epsilon=-1$ (right).
+The surface $\mathcal{B}^{\epsilon}$ is a null boundary for every null hypersurface $w=$const of $\mathscr{N}^{\epsilon}$ and is the boundary surface for the double-null CIBVPs to be presented, see Fig.(3).
+Figure 3: A family of null hypersurfaces $\mathscr{N}^{\epsilon}(w)$ with corresponding boundary null hypersurface $\mathcal{B}^{\epsilon}(w)$ is shown for $\epsilon=-1$ (left) and $\epsilon=+1$ (right).
 The former is a characteristic initial value problem with the initial data surface $\mathscr{N}^{-}(0)$ and the latter gives rise to a characteristic past value problem with the past data on $\mathscr{N}^{+}(0)$.
@@
 $ds^{2}=O(\lambda)(dw)^{2}+2\epsilon dwd\lambda+r^{2}(w,0)q_{AB}dx^{A}dx^{B}+O(\lambda^{3}).$ (72)
-This implies $V(w,0)=0$ for all values of $w$ and where $\lambda=0$.
+This implies $V(w,0)=0$ for all values of $w$, where $\lambda=0$.
 Consequently, $V=O(\lambda)$ near $\mathcal{B}^{\epsilon}$ on $\mathscr{N}^{\epsilon}$.
-For the conformal factor $r$ there are two possibilities regarding its evaluation on $\mathcal{B}^{\epsilon}$: i) $r(w,0)=0$ and ii) $r(w,0)\neq 0$.
-Physically these cases mean that the null hypersurface $\mathcal{B}^{\epsilon}$ has either a trivial (case (i)) or non-trivial Misner-Sharp mass (case (ii)), which becomes clear considering (21) in the limit $\lambda\rightarrow\infty$
+For the conformal factor $r$ there are two possibilities regarding its evaluation on $\mathcal{B}^{\epsilon}$: (i) $r(w,0)=0$, and (ii) $r(w,0)\neq 0$.
+Physically these cases mean that the null hypersurface $\mathcal{B}^{\epsilon}$ has either a trivial [case (i)] or nontrivial Misner-Sharp mass [case (ii)], which becomes clear considering (21) in the limit $\lambda\rightarrow\infty$
 $\lim_{\lambda\rightarrow\infty}M=\frac{r(w,0)}{2}\left[1-2\epsilon r_{,w}(w,0)r_{,\lambda}(w,0)\right]\;\;.$ (73)
@@
 III.3.1 Initial boundary formulation at a null hypersurface with $r(w,0)=0$
-If $V(w,0)=r(w,0)=0$, the line element near the origin of the affine parameter $\lambda$ can be written like
+If $V(w,0)=r(w,0)=0$, the line element near the origin of the affine parameter $\lambda$ can be written as
 $\begin{split}ds^{2}=&-\left[V_{(1)}(w)\lambda+O(\lambda^{2})\right](dw)^{2}+2\epsilon dwd\lambda\\ &+[r_{(1)}(w)\lambda]^{2}\left[1+O(\lambda)\right]q_{AB}dx^{A}dx^{B}\;,\end{split}$ (74)
@@
 $\begin{split}ds^{2}=&-\left\{\left[r_{(1)}V_{(1)}+2\epsilon r_{(1),\hat{w}}\right]\hat{\lambda}+O(\hat{\lambda}^{2})\right\}(d\hat{w})^{2}\\ &+2\epsilon d\hat{w}d\hat{\lambda}+\hat{\lambda}^{2}\left[1+O(\hat{\lambda})\right]q_{AB}dx^{A}dx^{B}.\end{split}$ (75)
-Consequently, we can set wlog $r_{(1)}=1$ which is assumed herefter in this subsection.
-With the assumption of a regular expansion of the scalar field like in (66), we find that $Z(w,0)=\mathcal{L}(w,0)=0$.
-The residual gauge freedom of the electromagnetic field $\alpha$, allows us to set $\alpha(w,0)=0$.
-In addition, the supplementary equation (18) implies $Q(w,0)=const$ for $\lambda=0$ and we may choose $Q(w,0)=0$.
-The metric supplementary equation (II), gives in the limit $\lambda\rightarrow 0$
-$0=V_{(1),w}+\epsilon\left[-V_{(1)}^{2}+\kappa|\Phi_{(0),w}|^{2}\right]$ (76)
-This equation can be solved along the null boundary $\mathcal{B}^{\epsilon}$ for $w>0$, provided an initial value $V_{(1)}(0)$ is given at $w=0$ and provided the function $\Phi_{(0),w}$ is known for all values $w\geq 0$.
-The function $V_{(1)}$ measures the inaffinity of the null geodesics $n^{a}\partial_{a}$ ruling the null hypersurface $\mathcal{B}^{\epsilon}$, which follows from (12).
-For a non-vanishing and $w$-dependent scalar field on $\mathcal{B}^{\epsilon}$, (76) shows that the evolution of the inaffinity along $\mathcal{B}^{\epsilon}$ is driven by the free function $\Phi_{(0),w}$.
-Thus, the function $\Phi_{(0),w}$ can be seen as some kind of analogue of the news function at $\lambda=0$ instead of in the asymptotic limit $\lambda\rightarrow\infty$.
-The characteristic initial value problem can now be constructed in the following way.
-Let $w=0$ be a null hypersurface $\mathscr{N}^{\epsilon}(0)$ whose generators are tangent to affinly parameterized null geodesics with tangent $\ell^{a}\partial_{a}$ and affine parameter $\lambda$.
+Consequently, we can set wlog $r_{(1)}=1$, which is assumed hereafter in this subsection.
+With the assumption of a regular expansion of the scalar field as in (66), we find that $Z(w,0)=\mathcal{L}(w,0)=0$.
+The residual gauge freedom of the electromagnetic field $\alpha$ allows us to set $\alpha(w,0)=0$.
+In addition, the supplementary equation (18) implies $Q(w,0)$=const for $\lambda=0$ and we set wlog $Q(w,0)=0$.
+The supplementary equation (II) gives in the limit $\lambda\rightarrow 0$
+$0=V_{(1),w}+\epsilon\left[-V_{(1)}^{2}+\kappa|\Phi_{(0),w}|^{2}\right].$ (76)
+This equation can be solved along the null boundary $\mathcal{B}^{\epsilon}$ for values $w\neq 0$, provided an initial value $V_{(1)}(0)$ is given at $w=0$ and provided the function $\Phi_{(0),w}$ is known for all values $w\geq 0$.
+The function $V_{(1)}$ measures the inaffinity of the null geodesics $n^{a}\partial_{a}$ generating the null hypersurface $\mathcal{B}^{\epsilon}$, which follows from (12).
+For a nonvanishing and $w$-dependent scalar field on $\mathcal{B}^{\epsilon}$, (76) shows that the evolution of the inaffinity along $\mathcal{B}^{\epsilon}$ is driven by the free function $\Phi_{(0),w}$.
+Thus, the function $\Phi_{(0),w}$ can be seen as some kind of analog of the news function at $\lambda=0$ instead of in the asymptotic limit $\lambda\rightarrow\infty$.
+A CIBVP can now be constructed in the following way.
+Let $w=0$ be a null hypersurface $\mathscr{N}^{\epsilon}(0)$ whose generators are tangent to affinely parametrized null geodesics with tangent $\ell^{a}\partial_{a}$ and affine parameter $\lambda$.
 At the origin of the affine parameter, $\lambda=0$, the null hypersurface $\mathscr{N}^{\epsilon}(0)$ intersects a null hypersurface $\mathcal{B}^{\epsilon}$ which has the generators $n^{a}\partial_{a}$ that obey the normalization $n^{a}\ell_{a}=-1$.
-The intersection of the two null hypersurfaces $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}$ is a two dimensional spacelike space with spherical topology and unit Gaussian curvature, called $\Sigma$.
+The intersection of the two null hypersurfaces $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}$ is a two-dimensional spacelike space with spherical topology and unit Gaussian curvature, called $\Sigma$.
 The intrinsic metric of $\Sigma$ is the one of the unit sphere, $q_{AB}$.
-For this double null characteristic initial value problem, the following specification of data determines a solution of the Einstein-scalar field equations in spherical symmetry:
-  • specify the type of null hypersurface $w=0$, by choosing either $\epsilon=1$ (ingoing null hypersurface) or $\epsilon=-1$ (outgoing null hypersurface)
-  • specify the gauge charge $q$
-  • on $\mathscr{N}^{\epsilon}(0)$ specify a free complex function $\mathcal{F}_{\Phi}(\lambda)$ that can be expressed by a regular Taylor series at $\lambda=0$
-  • on $\mathcal{B}^{\epsilon}$ specify a free complex function $\mathcal{N}_{\mathcal{B}}(w)$
-  • on $\Sigma$ specify the following fields: $\Phi(0,0)$ and $V_{(1)}(0)$
-  • on $\mathcal{B}^{\epsilon}$, i.e. for $\lambda=0$, the following fields are trivial: $r$, $V$, $Q$, $\alpha$, $Z$ and $\mathcal{L}$
-  • on $\mathcal{B}^{\epsilon}$, the field $r_{,\lambda}$ is nontrivial, i. e. $\lim_{\lambda\rightarrow 0}r_{,\lambda}=1$ for all values $w\geq 0$
-With these specifications the field equations are solved in the following way:
+For this double-null CIBVP, the following specification of data determines a solution of the Einstein-scalar field equations in spherical symmetry:
+  • specify the type of null hypersurface $w=0$, by choosing either $\epsilon=1$ (ingoing null hypersurface) or $\epsilon=-1$ (outgoing null hypersurface);
+  • specify the gauge charge $q$;
+  • on $\mathscr{N}^{\epsilon}(0)$ specify a free complex function $\mathcal{F}_{\Phi}(\lambda)$ that can be expressed by a regular Taylor series at $\lambda=0$;
+  • on $\mathcal{B}^{\epsilon}$ specify a free complex function $\mathcal{N}_{\mathcal{B}}(w)$;
+  • on $\Sigma$ specify the following fields: $\Phi(0,0)$ and $V_{(1)}(0)$;
+  • on $\mathcal{B}^{\epsilon}$, i.e. for $\lambda=0$, the following fields are trivial: $r$, $V$, $Q$, $\alpha$, $Z$ and $\mathcal{L}$;
+  • on $\mathcal{B}^{\epsilon}$, the field $r_{,\lambda}$ is nontrivial, i. e. $\lim_{\lambda\rightarrow 0}r_{,\lambda}=1$ for all values $w\geq 0$.
+With these specifications, the field equations are solved in the following way:
   1. 1.
@@
   2. 2.
-    with $\mathcal{N}_{\mathcal{B}}(w)$ and $V_{(1)}(0)$ determine the inaffinity everywhere on $\mathcal{B}^{\epsilon}$ via the equation (which follows from (76))
+    With $\mathcal{N}_{\mathcal{B}}(w)$ and $V_{(1)}(0)$ determine the inaffinity everywhere on $\mathcal{B}^{\epsilon}$ via the equation [which follows from (76)]
     $V_{(1),w}=-\epsilon\left[V_{(1)}^{2}+\kappa|\mathcal{N}_{\mathcal{B}}|^{2}\right].$ (78)
   3. 3.
-    with $\mathcal{F}_{\Phi}(\lambda)$ and $\Phi(0,0)$ determine the scalar field everywhere on $\mathscr{N}^{\epsilon}$ via the equation
+    With $\mathcal{F}_{\Phi}(\lambda)$ and $\Phi(0,0)$ determine the scalar field everywhere on $\mathscr{N}^{\epsilon}(0)$ via the equation
     $\Phi_{,\lambda}(0,\lambda)=\mathcal{F}_{\Phi}(\lambda).$ (79)
   4. 4.
-    as all boundary data for $r$, $Q$, $\alpha$, $Z$, $\mathcal{L}$ and $V$ are known on $\mathcal{B}^{\epsilon}$, the hypersurface equations (29a) - (29f) can be hierarchically integrated for all the fields relevant to determine the $w$-derivative of $\Phi$ in equation (29g)
+    As all boundary data for $r$, $Q$, $\alpha$, $Z$, $\mathcal{L}$ and $V$ are known on $\mathcal{B}^{\epsilon}$, the hypersurface equations (29a) - (29f) can be hierarchically integrated for all the fields relevant to determine the $w$-derivative of $\Phi$ in Eq.
+(29g)
   5. 5.
-    integration of the $w$-derivative of $\Phi$ by equation (29g) (using for example a finite difference scheme) determines $\Phi(\Delta w,\lambda)$ at some time $w=\Delta w>0$ on a null hypersurface $\mathscr{N}^{\epsilon}(\Delta w)$
+    Integration of the $w$-derivative of $\Phi$ by Eq.
+(29g) (using, for example, a finite difference scheme) determines $\Phi(\Delta w,\lambda)$ at some time $w=\Delta w\neq 0$ on a null hypersurface $\mathscr{N}^{\epsilon}(\Delta w)$
   6. 6.
     the new values $\Phi(\Delta w,\lambda)$ serve as initial data on $\mathscr{N}^{\epsilon}(\Delta w)$ and are used to determine the metric and electromagnetic fields following step 4.
-This algorithm can be suitably implemented in a numerical algorithm.
+This algorithm can be suitably implemented numerically.
 III.3.2 Initial boundary formulation at a null hypersurface with $r(w,0)\neq 0$
-If the areal distance function is non-trivial on the bounday $\mathcal{B}^{\epsilon}$, i.e.
-$r(w,0)\neq 0$, a different boundary value formulation will come about.
-Near the boundary the metric has the behaviour
+If the areal distance function is nontrivial on the boundary $\mathcal{B}^{\epsilon}$, i.e., $r(w,0)\neq 0$, a different boundary value formulation will come about.
+Near the boundary the metric has the behavior
 $\begin{split}ds^{2}=&-\left[V_{(1)}(w)\lambda+O(\lambda^{2})\right]dw^{2}+2\epsilon dwd\lambda\\ &+\left[r_{(0)}(w)+r_{(1)}(w)\lambda+O(\lambda^{2})\right]^{2}q_{AB}dx^{A}dx^{B}\;\;,\end{split}$ (80)
@@
 that casts (80) into a form similar to the expansion of (74).
-If that were the case, the initial boundary value problem could be discussed like in the previous section.
-It is indeed possible to find a coordinate transformation for (80) such that the resulting metric has $g_{AB}=\hat{\lambda}^{2}q_{AB}+O(\hat{\lambda}^{3})$, but such a transformation gives $g_{\hat{w}\hat{w}}|_{\hat{\lambda}=0}=O(1)\neq 0$ in the new coordinates.
+If that were the case, the initial boundary value problem could be discussed as in the previous section.
+It is, indeed, possible to find a coordinate transformation for (80) such that the resulting metric has $g_{AB}=\hat{\lambda}^{2}q_{AB}+O(\hat{\lambda}^{3})$, but such a transformation gives $g_{\hat{w}\hat{w}}|_{\hat{\lambda}=0}=O(1)\neq 0$ in the new coordinates.
 This violates the null condition at the boundary $\hat{\lambda}=0$ and cannot be removed without requiring $r_{(0),w}=0$.
-Since the coefficient $r_{(0)}$ corresponds to a mass, which becomes clear considering (73), we discard the restriction $r_{(0),w}=0$ as it implies that null boundary $\mathcal{B}^{\epsilon}$ has a mass which remains constant wrt a $w$ evolution.
+Since the coefficient $r_{(0)}$ corresponds to a mass, which becomes clear considering (73), we discard the restriction $r_{(0),w}=0$ as it implies that null boundary $\mathcal{B}^{\epsilon}$ has a mass which remains constant with respect to a $w$ evolution.
 Thus, having $r_{(0)}\neq 0$ on $\mathcal{B}^{\epsilon}$ indicates that $\mathcal{B}^{\epsilon}$ is a null hypersurface with a given mass.
@@
 Although there is no local coordinate transformation to cast (80) into a form like (74), it can still be simplified because $V_{(1)}$ can be set to zero everywhere on $\mathcal{B}^{\epsilon}$.
-The null vector generating $\mathcal{B}^{\epsilon}$ is $n^{a}\partial_{a}$ and the generating curves are non-affine null geodesics, which becomes clear when projecting (12) to $\mathcal{B}^{\epsilon}$ whose only component is
+The null vector generating $\mathcal{B}^{\epsilon}$ is $n^{a}\partial_{a}$ and the generating curves are nonaffine-null geodesics, which becomes clear when projecting (12) to $\mathcal{B}^{\epsilon}$ whose only component is
 $n^{a}\nabla_{a}n^{w}=-\frac{1}{2}V_{(1)}n^{w}.$ (82)
@@
 $n^{a}\nabla_{a}n^{\hat{w}}=-\left(\frac{d}{d\hat{w}}\ln\left|\frac{dw}{d\hat{w}}\right|+\frac{1}{2}V_{(1)}\right)n^{\hat{w}}.$ (83)
-If the coordinate transformation is chosen such that the term in the parenthesis of (83) vanishes, the non-affine null geodesics ruling $\mathcal{B}^{\epsilon}$ become affine null geodesics and $g_{\hat{w}\hat{w}}=O(\lambda^{2})$ near $\mathcal{B}^{\epsilon}$.
-This removal of the inaffinity is in fact a classical feature of non-affine geodesics Schouten (1954); Synge (1960); Poisson (2009) and not restricted to null geodesics.
+If the coordinate transformation is chosen such that the term in the parentheses of (83) vanishes, the nonaffine-null geodesics ruling $\mathcal{B}^{\epsilon}$ become affine-null geodesics and $g_{\hat{w}\hat{w}}=O(\lambda^{2})$ near $\mathcal{B}^{\epsilon}$.
+This removal of the inaffinity is, in fact, a classical feature of nonaffine geodesics Schouten (1954); Synge (1960); Poisson (2009) and not restricted to null geodesics.
 We make this choice in this section.
-In addition, we make the gauge choice $\alpha=0$ everwhere on $\mathcal{B}^{\epsilon}$.
+In addition, we make the gauge choice $\alpha=0$ everywhere on $\mathcal{B}^{\epsilon}$.
 Evaluation of the supplementary equations (II) and (18) on $\mathcal{B}^{\epsilon}$ yields
@@
 00 0 $=$ $\left[2\epsilon(rr_{,\lambda})_{,w}-1+\frac{\kappa}{8\pi}\frac{Q^{2}}{r^{2}}\right]\bigg{|}_{\mathcal{B}^{\epsilon}}\;\;.$ (84c)
-Equation (84c) can be used to determine $r_{,\lambda}$ everwhere on $\mathcal{B}^{\epsilon}$ provided $r$ and $Q$ are known on $\mathcal{B}^{\epsilon}$ and $r\neq 0$ on $\mathcal{B}^{\epsilon}$.
-The start values for solving the hypersurface equations (29d) and (29e) for $Z$ and $\mathcal{L}$ follow from evaluation of their definitions on $\mathcal{B}^{\epsilon}$
+Equation (84c) can be used to determine $r_{,\lambda}$ everywhere on $\mathcal{B}^{\epsilon}$ provided $r$ and $Q$ are known on $\mathcal{B}^{\epsilon}$ and $r\neq 0$ on $\mathcal{B}^{\epsilon}$.
+The initial values for solving the hypersurface equations (29d) and (29e) for $Z$ and $\mathcal{L}$ follow from their definitions on $\mathcal{B}^{\epsilon}$
 $Z\Big{|}_{\mathcal{B}^{\epsilon}}=2\epsilon(rr_{,w})\Big{|}_{\mathcal{B}^{\epsilon}}\;\;,\;\;\mathcal{L}\Big{|}_{\mathcal{B}^{\epsilon}}=2(r\Phi_{,w})\Big{|}_{\mathcal{B}^{\epsilon}}\;\;,$ (85)
 which show that the initial values for $Z$ and $\mathcal{L}$ are known once $r$ and $\Phi_{,w}$ are known on $\mathcal{B}^{\epsilon}$.
-With the information collected above, a double null characteristic initial value problem can be formulated with a nontrivial conformal factor $r$ on the null boundary $\mathcal{B}^{\epsilon}$.
-In a spherically symmetric spacetime, let $\mathscr{N}^{\epsilon}(w)$ be a family of null hypersurface $w=const$ with generators $\ell^{a}\partial_{a}$ that are affinely parameterized with an affine parameter $\lambda$.
+With the information collected above, a double-null CIBVP can be formulated with a nontrivial conformal factor $r$ on the null boundary $\mathcal{B}^{\epsilon}$.
+In a spherically symmetric spacetime, let $\mathscr{N}^{\epsilon}(w)$ be a family of null hypersurface $w=$const with generators $\ell^{a}\partial_{a}$ that are affinely parametrized with an affine parameter $\lambda$.
 The affine parameter $\lambda$ is normalized with a parameter $\epsilon$ via $\ell^{a}\nabla_{a}\lambda=\epsilon$ with $\epsilon^{2}=1$.
@@
 The generators of $\mathcal{B}^{\epsilon}(w)$ are null vectors $n^{a}\partial_{a}$.
-These generators obey the cross normalisation $\ell^{a}n_{a}=-1$ and the integral curves (null geodesics) of $n^{a}$ are affinely parameterized with the parameter $w$ on $\mathcal{B}^{\epsilon}(w)$.
-The intersections of $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}(w)$ are spatial spherical cross sections $\Sigma(w)$ of co-dimension 2 with an intrinsic two metric $g_{AB}$.
+These generators obey the cross normalization $\ell^{a}n_{a}=-1$ and the integral curves (null geodesics) of $n^{a}$ are affinely parametrized with the parameter $w$ on $\mathcal{B}^{\epsilon}(w)$.
+The intersections of $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}(w)$ are spatial spherical cross sections $\Sigma(w)$ of codimension-2 with an intrinsic two metric $g_{AB}$.
 The volume element $\sqrt{\det(g_{AB})}=r^{2}(w,0)\sqrt{q}$ with $\sqrt{q}$ being the volume element of a unit sphere metric, $q_{AB}$, is nowhere zero on $\mathcal{B}^{\epsilon}(w)$ and for $\lambda=0$ on $\mathscr{N}^{\epsilon}(w)$.
-The function $r(w,0)$ is extended for values $\lambda>0$ on every null hypersurface $w=const$ to be be an areal distance so that the volume integral for a value $\lambda_{c}>0$ is given by $4\pi r^{2}(w,\lambda_{c})$.
+The function $r(w,0)$ is extended for values $\lambda>0$ on every null hypersurface $w=$const to be an areal distance so that the volume integral for a value $\lambda_{c}>0$ is given by $4\pi r^{2}(w,\lambda_{c})$.
 A solution of the Einstein-Maxwell-scalar field equations on the domain $\mathcal{B}^{\epsilon}(w)\cup\mathscr{N}^{\epsilon}(w)$ can be found for the double-null CIBVP with the following specifications:
-  • provide a value for $\epsilon$ to specify the nature of the initial data surface $\mathscr{N}^{\epsilon}(0)$ with $w=0$
-  • specify the value of the gauge charge $q$
-  • on $\mathscr{N}^{\epsilon}(0)$ specify a free (sufficiently differentiable) complex function $F_{\Phi}(\lambda)$
-  • on $\mathcal{B}^{\epsilon}(0)$ specify a free (sufficiently differentiable) complex function $\mathcal{N}_{\mathcal{B}}(w)$
+  • provide a value for $\epsilon$ to specify the nature of the initial data surface $\mathscr{N}^{\epsilon}(0)$ with $w=0$;
+  • specify the value of the gauge charge $q$;
+  • on $\mathscr{N}^{\epsilon}(0)$ specify a free (sufficiently differentiable) complex function $F_{\Phi}(\lambda)$;
+  • on $\mathcal{B}^{\epsilon}(0)$ specify a free (sufficiently differentiable) complex function $\mathcal{N}_{\mathcal{B}}(w)$;
   • the following fields are trivial everywhere on $\mathcal{B}^{\epsilon}(w)$
@@
     where in particular $r(0,0)>0$.
-This (double null) CIBVP can be solved in the following manner
+This double-null CIBVP can be solved in the following manner
   1. 1.
     With $\mathcal{N}_{\mathcal{B}}(w)$ and $\Phi(0,0)$ determine the scalar field everywhere on $\mathcal{B}^{\epsilon}(w)$ via the equation
-    $\Phi_{,w}(w,0)=\mathcal{N}_{\mathcal{B}}(w)$ (88)
+    $\Phi_{,w}(w,0)=\mathcal{N}_{\mathcal{B}}(w).$ (88)
   2. 2.
-    with the boundary data $\mathcal{N}_{\mathcal{B}}(w)$, the scalar field $\Phi(w,0)$ on $\mathcal{B}^{\epsilon}(w)$ and the initial values $r(0,0)$, $r_{,w}(0,0)$, $Q(0,0)$ and $r_{\lambda}(0,0)$ solve the hierarchical set of boundary equations
+    With the boundary data $\mathcal{N}_{\mathcal{B}}(w)$, the scalar field $\Phi(w,0)$ on $\mathcal{B}^{\epsilon}(w)$, and the initial values $r(0,0)$, $r_{,w}(0,0)$, $Q(0,0)$ and $r_{\lambda}(0,0)$ solve the hierarchical set of boundary equations
     00 0 $=$ $r_{,ww}(w,0)+\frac{1}{2}\kappa r(w,0)|\mathcal{N}_{\mathcal{B}}|^{2}\;\;,$ (89)
-    00 0 $=$ $Q_{,w}(w,0)+2i\pi qr^{2}(w,0)\left[\overline{\Phi}(w,0)\mathcal{N}_{\mathcal{B}}-\Phi(w,0)\overline{\mathcal{N}}_{\Phi}\right]\bigg{|}_{\mathcal{B}^{\epsilon}}\;\;.$ (90)
+    00 0 $=$ $Q_{,w}(w,0)+2i\pi qr^{2}(w,0)\left[\overline{\Phi}(w,0)\mathcal{N}_{\mathcal{B}}-\Phi(w,0)\overline{\mathcal{N}}_{\Phi}\right]\bigg{|}_{\mathcal{B}^{\epsilon}}\;\;,$ (90)
     00 0 $=$ $\left[2\epsilon(rr_{,\lambda})_{,w}-1+\frac{\kappa}{8\pi}\frac{Q^{2}}{r^{2}}\right]\bigg{|}_{\mathcal{B}^{\epsilon}}.$ (91)
@@
     On the initial null hypersurface $w=0$ utilize the initial value $\Phi(0,0)$ on $\Sigma(0)$ and the initial data $\mathcal{F}_{\Phi}$ to integrate
-    $\Phi_{,\lambda}(0,\lambda)=\mathcal{F}_{\Phi}(\lambda)$ (92)
+    $\Phi_{,\lambda}(0,\lambda)=\mathcal{F}_{\Phi}(\lambda),$ (92)
     for the initial data $\Phi(0,\lambda)$.
@@
     on $\mathscr{N}^{\epsilon}(0)$, i.e.
-the null hypersurface $w=0$, integrate hierarchically the hypersurface equations (29a)-(29f) to obtain the fields $r(0,\lambda)$, $Q(0,\lambda)$, $\alpha(0,\lambda)$, $Z(0,\lambda)$, $\mathcal{L}(0,\lambda)$ and $V(0,\lambda)$ while using $\mathcal{F}_{\Phi}(\lambda)$, $\Phi(0,\lambda)$ and the respective boundary value fields evaluated at $w=0$.
+the null hypersurface $w=0$, integrate hierarchically the hypersurface equations (29a)-(29f) to obtain the fields $r(0,\lambda)$, $Q(0,\lambda)$, $\alpha(0,\lambda)$, $Z(0,\lambda)$, $\mathcal{L}(0,\lambda)$, and $V(0,\lambda)$, while using $\mathcal{F}_{\Phi}(\lambda)$, $\Phi(0,\lambda)$ and the respective boundary value fields evaluated at $w=0$.
   5. 5.
-    with $r(0,\lambda)$, $\mathcal{L}(0,\lambda)$, $V(0,\lambda)$ and $\Phi_{,\lambda}(0,\lambda)$ algebraically calculate $\Phi_{,w}(0,\lambda)$ on $w=0$ using equation (29g), integrate this $w$ derivative (using for example a finite difference scheme) to find $\Phi(\Delta w,\lambda)$ at some time $w=\Delta w>0$ on a null hypersurface $\mathscr{N}^{\epsilon}(\Delta w)$
+    With $r(0,\lambda)$, $\mathcal{L}(0,\lambda)$, $V(0,\lambda)$, and $\Phi_{,\lambda}(0,\lambda)$ algebraically calculate $\Phi_{,w}(0,\lambda)$ on $w=0$ using Eq.
+(29g), integrate this $w$ derivative (using, for example, a finite difference scheme) to find $\Phi(\Delta w,\lambda)$ at some time $w=\Delta w\neq 0$ on a null hypersurface $\mathscr{N}^{\epsilon}(\Delta w)$
   6. 6.
     The new values $\Phi(\Delta w,\lambda)$ serve as initial data on $\mathscr{N}^{\epsilon}(\Delta w)$ and are used to determine the metric and electromagnetic fields following step 4 above but using the coordinate time $w=\Delta w$ instead of $w=0$.
-Comparing this formulation of the double null CIBVP with the one of the previous section Sec.
+Comparing this formulation of the double-null CIBVP with the one of the previous section Sec.
 III.3.1, reveals that the two are very similar.
-In fact, the only essential difference is step 2 - the determination of the nontrivial boundary fields in $\mathcal{B}^{\epsilon}(w)$.
-The reason for this difference is that for one of the formulations $r=0$ on $\mathcal{B}^{\epsilon}(w)$ while for the other one it is essentially non-trivial.
+In fact, the only essential difference is step 2 – the determination of the nontrivial boundary fields in $\mathcal{B}^{\epsilon}(w)$.
+The reason for this difference is that for one of the formulations $r=0$ on $\mathcal{B}^{\epsilon}(w)$, while for the other one it is essentially nontrivial.
 The two different formulations give rise for different areas of application.
@@
 Indeed, in Sec.
-V.1 the second scheme is used to derive the Reisner-Nordström solution of a charged black hole.
+V.1 the second scheme is used to derive the Reissner-Nordström solution of a charged black hole.
 IV Symmetry considerations
@@
 which is invariant under the action of the vector field $\xi$ that is the Lie derivative $\mathscr{L}_{\xi}z=\xi^{a}\partial_{a}z=0$.
-In principle we may choose $B=0$, because of the metric’s invariance of the coordinate transformation of the type (10).
-However, we will keep it general to allow a broader family of null affine coordinates.
+In principle, we may choose $B=0$, because of the metric’s invariance of the coordinate transformation of the type (10).
+However, we will keep it general to allow a broader family of affine-null coordinates.
 Inserting (95) into (94c) yields
@@
 $r=\hat{r}(z)\;,$ (98)
-where $\hat{r}(\cdot)$ is any non-vanishing differentiable function. Finally, inserting (95) into (94d) gives us
+where $\hat{r}(\cdot)$ is any nonvanishing differentiable function. Finally, inserting (95) into (94d) gives us
 $\begin{split}0=&V_{,w}A-V_{,\lambda}(A_{,w}\lambda-B)-2\epsilon\left(A_{,ww}\lambda-B_{,w}\right)+2VA_{,w}\end{split}$ (99)
@@
 $V=\frac{\hat{V}[z(w,\lambda)]}{A^{2}(w)}-\frac{2\epsilon(A_{,w}\lambda-B)}{A}\;,$ (100)
-in which $\hat{V}(\cdot)$ is an arbitrary differentible function.
+in which $\hat{V}(\cdot)$ is an arbitrary differentiable function.
 Some choices for $A$ and $B$ relevant in this work are listed in Table  1.
@@
 The coordinate dependence of the metric potentials $r$ and $V$ for the different expressions for Killing vectors in Table 1 is general.
-For further study of particular applications it is practical to study the norm of the Killing vector either for large values of the affine parameter or its origin.
+For further particular applications, it is practical to study the norm of the Killing vector either for large values of the affine parameter or at its origin.
 The former will give us restrictions to the metric by requiring the Killing vector be asymptotically timelike, while the latter gives us necessary conditions, if the origin of the affine parameter is on a Killing horizon with a given surface gravity.
 The norm of the Killing vector (95) is
-$\xi^{a}\xi_{a}=-VA^{2}-\epsilon(A^{2})_{,w}\lambda+2\epsilon AB$ (101)
-and in its asymptotic limit ($\lambda\rightarrow\infty$) we find using (36g), (37) and (LABEL:eq:Vinf)
+$\xi^{a}\xi_{a}=-VA^{2}-\epsilon(A^{2})_{,w}\lambda+2\epsilon AB.$ (101)
+In its asymptotic limit ($\lambda\rightarrow\infty$), we find using (36g), (37) and (38):
 $\xi^{a}\xi_{a}=$ $-2\epsilon A^{2}\left[\left(\ln\frac{A}{H}\right)_{,w}\right]\lambda$
@@
 Therefore, in order to have a finite norm of this asymptotic limit, the linear term in $\lambda$ should vanish.
-One posibility is to set $A=0$, which from (101) implies a Killing vector field with vanishing norm of the form $\xi=B\partial_{\lambda}$.
-However, taking into account (98) we see that it implies that the vector field is zero everywhere or $\xi$ is a null Killing vector with $r=r(w)$ everywhere, which is inadmissible.
+One possibility is to set $A=0$, which from (101) implies the Killing vector field $\xi=B\partial_{\lambda}$ with vanishing norm.
+However, taking into account (98), we see that it implies that the vector field is zero everywhere or $\xi$ is a null Killing vector with $r=r(w)$ everywhere, which is inadmissible.
 The other choice for the finiteness of the norm of the Killing vector for $\lambda\rightarrow\infty$ is
@@
 where $c$ is a nontrivial constant.
-If the Killing vector is required to be an asymptotically timelike unit Killing vector, one finds the condition
-$-1=\left[2\epsilon B-\frac{c-2c\epsilon H(R_{\infty})_{,w}}{H}\right]cH$ (105)
+If the Killing vector is required to be an asymptotically timelike unit vector, one finds the condition
+$-1=\left[2\epsilon B-\frac{c-2c\epsilon H(R_{\infty})_{,w}}{H}\right]cH,$ (105)
 or
@@
 The relations (104) and (106) show that the components of a timelike Killing vector can be completely fixed by the asymptotic solution of the field equation.
-In particular, it is determined by the asymptotic behavior of the areal distance $r$ and an asymptotic time like Killing vector should have the form
-$\begin{split}\xi=&cH(w)\partial_{w}+\\ &\left[-cH_{,w}\lambda+\frac{c\epsilon}{2H}-c(R_{\infty})_{,w}-\frac{\epsilon}{2cH}\right]\partial_{\lambda}+O(\lambda^{-1}).\end{split}$ (107)
-In particular, regarding the choices of $A$ and $B$ in table 1 and requiring the Killing vectors to be asymptotically timelike unit Killing vectors implies using (98), (103) and (106).
+In particular, it is determined by the asymptotic behavior of the areal distance $r$.
+An asymptotic timelike Killing vector should therefore have the form
+$\begin{split}\xi=&cH(w)\partial_{w}\\ &+\left[-cH_{,w}\lambda+\frac{c\epsilon}{2H}-c(R_{\infty})_{,w}-\frac{\epsilon}{2cH}\right]\partial_{\lambda}+O(\lambda^{-1}).\end{split}$ (107)
+In particular, regarding the choices of $A$ and $B$ in Table 1 and requiring the Killing vectors to be asymptotically timelike unit vectors implies using (98), (103) and (106):
   1. i)
-    … case $A=1$ and $B=0$, we find $c=1/H$.
+    For case $A=1$ and $B=0$, we find $c=1/H$.
 Moreover, since $r=r(\lambda)$, $H$=const and $R_{\infty}$=const.
-Consequently, $\xi^{a}\xi_{a}=-1$ in the asymptotic limit can only be archived if $H=1$, i.e.
+Consequently, $\xi^{a}\xi_{a}=-1$ in the asymptotic limit can only be achieved if $H=1$; i.e.
 the frame at large radii is a Bondi frame and the areal distance has the asymptotic expansion
@@
   2. ii)
-    …case $A=kw$ and $B=0$ gives $kw=cH(w)$.
-As Table 1 shows $r=r(w\lambda)$, the asymptotic expansion of $r$ in (36a) becomes $r(w\lambda)=\frac{kw\lambda}{c}+R_{\infty}(w)+O(\lambda^{-1})$ which requires $R_{\infty}=const$.
-But then the normalisation $\xi^{a}\xi_{a}=-1$ implies $c^{2}=1$, so that we can choose $c=1$.
+    The case $A=kw$ and $B=0$ gives $kw=cH(w)$.
+As Table 1 shows $r=r(w\lambda)$, the asymptotic expansion of $r$ in (36a) becomes $r(w\lambda)=\frac{kw\lambda}{c}+R_{\infty}(w)+O(\lambda^{-1})$, which requires $R_{\infty}$=const.
+But then the normalization $\xi^{a}\xi_{a}=-1$ implies $c^{2}=1$, so that we can choose $c=1$.
 Consequently $r$ has the asymptotic expansion
@@
   3. iii)
-    …case $A=\frac{1}{2}w^{2}(w+k)^{-2}$ and $B=1$ gives ${w^{2}=2c(w+k)^{2}H}$. Then fixing $H$ in the asymptotic behavior of $r$ gives for (36a)
+    The case $A=\frac{1}{2}w^{2}(w+k)^{-2}$ and $B=1$ gives ${w^{2}=2c(w+k)^{2}H}$. Then fixing $H$ in the asymptotic behavior of $r$ gives for (36a)
     $r(z)=\frac{w^{2}c}{2(w+k)^{2}}\lambda+R_{\infty}(w)+O(\lambda^{-1})\;\;.$ (110)
@@
     which implies by comparison with (110) that
-    $R_{\infty}(w)=-cw+r_{\infty}\;\;,r_{\infty}=const.$ (111)
-    To be consistient with the invariant, any solution for $r$ of the field equations with this Killing vector, which is in addition an asymptotical timelike unit vector, must have the large $\lambda$ expansion of the form
+    $R_{\infty}(w)=-cw+r_{\infty}\;\;,r_{\infty}=\mathrm{const}.$ (111)
+    To be consistent with the invariant, any solution for $r$ of the field equations with this Killing vector, which is in addition an asymptotical timelike unit vector, must have the large $\lambda$ expansion of the form
     $r(z)=r_{\infty}+c\left[w-\frac{w^{2}}{2(w+k)^{2}}\lambda\right]+O(\lambda^{-1}).$ (112)
-    The determination of $c$ for the present case is postponed to V.1, because of the subtlety of the behavior of the Killing vector for $w=0$.
-There, it is a non-vanishing null tangent vector to the hypersurface, consequently, $\xi|_{w=0}=\partial_{\lambda}$.
-Requiring regularity of this surface (which implies regularity of $V$ given by (100) can be used to fix the value of the constant $c$ (see next section for an explicit example in an extremal Reissner-Nordströn black hole).
-Of interest is also the behaviour of the norm of the Killing vector for $\lambda=0$, i.e on $\mathcal{B}^{\epsilon}$.
-At the origin ($\lambda=0$) of the affine parameter, the norm (101) takes the value
+    The determination of $c$ for the present case is postponed to Sec.
+V.1, because of the subtlety of the behavior of the Killing vector for $w=0$.
+There, it is a nonvanishing null tangent vector to the hypersurface, consequently, $\xi|_{w=0}=\partial_{\lambda}$.
+A regularity requirement of this surface [which implies regularity of $V$ given by (100)] can be used to fix the value of the constant $c$ (see next section for an explicit example in an extremal Reissner-Nordström black hole).
+Of interest is also the behavior of the norm of the Killing vector for $\lambda=0$, i.e on $\mathcal{B}^{\epsilon}$. At the origin ($\lambda=0$) of the affine parameter, the norm (101) takes the value
 $\xi^{a}\xi_{a}|_{\mathcal{B}^{\epsilon}}=-A^{2}V|_{\mathcal{B}^{\epsilon}}+2\epsilon AB.$ (113)
-If $V|_{\lambda=0}=0$ the surface $\lambda=0$ is a null hypersurface $\mathcal{B}^{\epsilon}$ and if $B=0$ and $A\neq 0$ the null hypersurface $\mathcal{B}^{\epsilon}$ is a Killing horizon [^5] If $A=0$, we also have $\xi^{a}\xi_{a}=0$, but it is not a Killing horizon because only a vector field of the form $\xi\propto\partial_{w}$ can be a null generator of a Killing horizon at a null surface determined by $\lambda=0$.
+If $V|_{\lambda=0}=0$ the surface $\lambda=0$ is a null hypersurface $\mathcal{B}^{\epsilon}$, and if $B=0$ and $A\neq 0$ the null hypersurface $\mathcal{B}^{\epsilon}$ is a Killing horizon [^8] If $A=0$, we also have $\xi^{a}\xi_{a}=0$, but it is not a Killing horizon because only a vector field of the form $\xi\propto\partial_{w}$ can be a null generator of a Killing horizon at a null surface determined by $\lambda=0$.
 However, the possibility $A=B=0$ can be satisfied at an eventual bifurcating Killing horizon, where the Killing vector field vanishes..
-Indeed, $B\neq 0$ in case iii) is crucial for the discussion of the (vanishing) surface gravity of an extremal black hole as is the case of extremal Reissner-Nordström black holes discussed in the next section.
+Indeed, $B\neq 0$ in case (iii) is crucial for finding the (vanishing) surface gravity of an extremal black hole as is the case of extremal Reissner-Nordström black holes discussed in the next section.
 It turns out that for case iii) there also exists a Killing horizon not necessarily located at $\lambda=0$ but given implicitly by an expression of the form $F(w,\lambda)=0$.
@@
 V.1 Charged black hole solutions in null coordinates
-For finding charged black hole solutions we work with the double null foliation of Sec. III.3 where the null boundary has nontrivial Misner-Sharp mass.
+For finding charged black hole solutions, we work with the double-null foliation of Sec. III.3, where the null boundary has nontrivial Misner-Sharp mass.
 We assume that the scalar field on the initial slice $\mathscr{N}^{\epsilon}(0)$ and the boundary $\mathcal{B}^{\epsilon}(w)$ is trivial meaning $\mathcal{F}_{\Phi}=\mathcal{N}_{\mathcal{B}}=0$.
 We use $\kappa=8\pi$ and the most general values for the remaining data at the common intersection $\Sigma(0)$
-$\begin{split}&r(0,0)=r_{0}>0\;\;,\;\;r_{,w}(0,0)=r_{N}\;\;,\;\;\\ &Q(0,0)=Q_{0}\;\;,\;\;r_{,\lambda}(0,0)=\Theta\end{split}$ (115)
+$\begin{split}&r(0,0)=r_{0}>0\;\;,\;\;r_{,w}(0,0)=r_{N}\;\;,\;\;\\ &Q(0,0)=Q_{0}\;\;,\;\;r_{,\lambda}(0,0)=\Theta.\end{split}$ (115)
 The boundary equations on $\mathcal{B}^{\epsilon}$ are
@@
 $(rr_{,\lambda})_{,w}\big{|}_{\mathcal{B}^{\epsilon}}$ $=$ $\frac{\epsilon}{2}\left(1-\frac{Q^{2}}{r^{2}}\right)\bigg{|}_{\mathcal{B}^{\epsilon}}.$ (116c)
-Integration of (116a) and (116b) the first two gives us
+Integration of (116a) and (116b) gives us
 $r(w,0)=$ $r_{\mathcal{B}}:=r_{0}+r_{N}w\;\;,$ (117a)
-$Q(w,0)=$ $Q_{0}\;\;,\;\;$ (117b)
-inserting these values into (116c) and integration yields
+$Q(w,0)=$ $Q_{0}\;\;.\;\;$ (117b)
+Insert these values into (116c) and integration yields
 $\begin{split}(r_{,\lambda})\big{|}_{\mathcal{B}^{\epsilon}}=&r_{1}(w):=\frac{r_{0}\Theta}{r_{\mathcal{B}}}+\frac{\epsilon}{2}\left(\frac{r_{0}r_{\mathcal{B}}-Q_{0}^{2}}{r_{0}r_{\mathcal{B}}^{2}}\right)\left(\frac{r_{\mathcal{B}}-r_{0}}{r_{N}}\right)\\ =&\frac{\Theta r_{0}}{r_{N}w+r_{0}}+\frac{(r_{0}r_{N}w-Q_{0}^{2}+r_{0}^{2})\epsilon w}{2r_{0}(r_{N}w+r_{0})^{2}}\;\;.\end{split}$ (118)
@@
 $Z(w,0)=2\epsilon r_{\mathcal{B}}r_{N}\;\;,\;\;\mathcal{L}(w,0)=0\;\;.$ (119)
-Now the hypersurface equation for the given initial boundary value data are
+Now the hypersurface equations for the given initial boundary value data are
 $r_{,\lambda\lambda}$ $=$ $0,$ (120a)
@@
 $V_{,\lambda\lambda}$ $=$ $-\frac{1}{\lambda}\left(\frac{\lambda^{2}}{r^{2}}{}\right)_{,\lambda}+\frac{2Zr_{,\lambda}}{r^{3}}+\frac{4Q^{2}}{r^{4}}.$ (120f)
-The integration of (120) with the boundary data (117), (118) and (119) gives us
+The integration of (120) with the boundary data (117), (118), and (119) gives us
 $r$ $=$ $r_{\mathcal{B}}+r_{1}\lambda,$ (121a)
@@
 $=$ $\left(\frac{\lambda}{r}\right)^{2}\left[\frac{Q_{0}^{2}}{r_{\mathcal{B}}^{2}}+\frac{\left(Q^{2}_{0}-r_{\mathcal{B}}^{2}+2\epsilon r_{N}r_{1}r_{\mathcal{B}}^{2}\right)r}{r_{\mathcal{B}}^{3}}\right].$
-Consequently the most general form of metric a function of the initial values reads
+Consequently, the most general form of the metric for these initial values reads
 $\begin{split}ds^{2}=&-\left(\frac{\lambda}{r}\right)^{2}\left[\frac{Q_{0}^{2}}{r_{\mathcal{B}}^{2}}+\left(Q^{2}_{0}-r_{\mathcal{B}}^{2}+2\epsilon r_{N}r_{1}r_{\mathcal{B}}^{2}\right)\left(\frac{r}{r_{\mathcal{B}}^{3}}\right)\right]dw^{2}\\ &+2\epsilon dwd\lambda+r^{2}q_{AB}dx^{A}dx^{B}\;,\\ r=&r_{0}+r_{N}w+\left[\frac{\Theta r_{0}}{r_{N}w+r_{0}}+\frac{(r_{0}r_{N}w-Q_{0}^{2}+r_{0}^{2})\epsilon w}{2r_{0}(r_{N}w+r_{0})^{2}}\right]\lambda,\end{split}$ (122)
@@
 On $\mathcal{B}^{\epsilon}$, it reduces to
-$ds^{2}=(r_{0}+r_{N}w)^{2}q_{AB}dx^{A}dx^{B}$ (123)
-and for $w=0=dw$ that is on $\mathscr{N}^{\epsilon}(0)$
-$ds^{2}=(r_{0}+\Theta\lambda)^{2}q_{AB}dx^{A}dx^{B}$ (124)
+$ds^{2}=(r_{0}+r_{N}w)^{2}q_{AB}dx^{A}dx^{B},$ (123)
+and on $\mathscr{N}^{\epsilon}(0)$, where $w=0=dw$,
+$ds^{2}=(r_{0}+\Theta\lambda)^{2}q_{AB}dx^{A}dx^{B}.$ (124)
 Since $w$ and $\lambda$ are the affine parameters on the null hypersurfaces $\mathcal{B}^{\epsilon}(w)$ and $\mathscr{N}^{\epsilon}(0)$, respectively, the initial data parameters $r_{N}$ and $\Theta$ correspond to the expansion rates of the generators of these surfaces.
 Calculation of the Misner-Sharp mass (21) gives us
-$M=$ $\frac{1}{2}\left[r_{\mathcal{B}}-2\epsilon r_{\mathcal{B}}r_{N}r_{1}+Q^{2}_{0}\left(\frac{1}{r_{\mathcal{B}}}-\frac{1}{r}\right)\right]$ (125)
+$M=$ $\frac{1}{2}\left[r_{\mathcal{B}}-2\epsilon r_{\mathcal{B}}r_{N}r_{1}+Q^{2}_{0}\left(\frac{1}{r_{\mathcal{B}}}-\frac{1}{r}\right)\right],$ (125)
 with the limiting values
@@
 so that the composite limit
-$\lim_{w\rightarrow 0\atop\lambda\rightarrow 0}M=$ $\frac{r_{0}}{2}(1-2\epsilon r_{N}\Theta)$ (128)
+$\lim_{w\rightarrow 0\atop\lambda\rightarrow 0}M=$ $\frac{r_{0}}{2}(1-2\epsilon r_{N}\Theta),$ (128)
 shows that $r_{0}=2M$ if either $r_{N}$ or $\Theta$ vanishes.
 As the limit
-$\lim_{\lambda\rightarrow\infty}M=$ $\frac{r_{\mathcal{B}}}{2}\left[1-2\epsilon r_{N}r_{1}+\frac{Q^{2}_{0}}{r_{\mathcal{B}}^{2}}\right]$ (129)
+$\lim_{\lambda\rightarrow\infty}M=$ $\frac{r_{\mathcal{B}}}{2}\left[1-2\epsilon r_{N}r_{1}+\frac{Q^{2}_{0}}{r_{\mathcal{B}}^{2}}\right],$ (129)
 gives rise to the Bondi mass, we find the most general expression for the Bondi mass of the solution to be
@@
 If either $r_{N}=0$ or $\Theta=0$, the initial value $r_{0}$ can be fixed by the Bondi mass, $m_{B}$, to find
-$r_{0,\pm}=m_{B}\pm\sqrt{m_{B}^{2}-Q^{2}_{0}}$ (131)
-where the numerical value $r_{0,+}$ is the same value as the value of the radius of the outer event horion of a Reissner-Nordström black hole and $r_{0,-}$ the value of the inner event horizon.
-V.1.1 Sub-extremal black holes
-Suppose $r_{N}=0$, then the areal distance has the form
+$r_{0,\pm}=m_{B}\pm\sqrt{m_{B}^{2}-Q^{2}_{0}},$ (131)
+where the numerical value $r_{0,+}$ is the same value as the value of the radius of the outer event horizon of a Reissner-Nordström black hole and $r_{0,-}$ the value of the inner event horizon.
+V.1.1 Subextremal black holes
+Suppose $r_{N}=0$; then the areal distance has the form
 $r=r_{0}+\left[\Theta-\epsilon\frac{Q^{2}_{0}-r^{2}_{0}}{2r_{0}^{3}}w\right]\lambda\;\;.$ (132)
-Since $w$ parameterises affine geodesics on the boundary $\mathcal{B}^{\epsilon}$ where $\lambda=0$, we have the freedom of choice in the origin of the affine parameter. The transformation
-$w\rightarrow\tilde{w}=w-\epsilon\Theta\frac{2r_{0}^{3}}{Q_{0}^{2}-r_{0}^{2}}$ (133)
+Since $w$ parametrizes affine geodesics on the boundary $\mathcal{B}^{\epsilon}$ where $\lambda=0$, we have the freedom of choice in the origin of the affine parameter. The transformation
+$w\rightarrow\tilde{w}=w-\epsilon\Theta\frac{2r_{0}^{3}}{Q_{0}^{2}-r_{0}^{2}},$ (133)
 gives the symmetric expression
-$r=\tilde{r}(\lambda\tilde{w}):=r_{0}-\epsilon\frac{Q^{2}_{0}-r^{2}_{0}}{2r_{0}^{3}}(\tilde{w}\lambda)$ (134a)
+$r=\tilde{r}(\lambda\tilde{w}):=r_{0}-\epsilon\frac{Q^{2}_{0}-r^{2}_{0}}{2r_{0}^{3}}(\tilde{w}\lambda),$ (134a)
 and
 $V$ $=-2\lambda^{2}\frac{\epsilon(\tilde{w}\lambda)(Q_{0}^{2}-r_{0}^{2})^{2}-2r_{0}^{4}(2Q^{2}_{0}-r_{0}^{2})}{\left[\epsilon(\tilde{w}\lambda)(Q_{0}^{2}-r_{0}^{2})-2r_{0}^{4}\right]^{2}}$
-$\equiv\lambda^{2}\tilde{V}(\tilde{w}\lambda)$ (134b)
+$\equiv\lambda^{2}\tilde{V}(\tilde{w}\lambda),$ (134b)
 thereby showing the two functions depend on the product $w\lambda$. For simplicity, we will write $\tilde{w}$ as $w$, hereafter.
 Therefore, the metric
-$\begin{split}ds^{2}=-\lambda^{2}\tilde{V}(\lambda w)dw^{2}+2\epsilon dwd\lambda+\tilde{r}^{2}(\lambda w)q_{AB}dx^{A}dx^{B}\end{split}$ (135)
+$\begin{split}ds^{2}=-\lambda^{2}\tilde{V}(\lambda w)dw^{2}+2\epsilon dwd\lambda+\tilde{r}^{2}(\lambda w)q_{AB}dx^{A}dx^{B},\end{split}$ (135)
 is invariant under the transformation $(w,\lambda)\rightarrow(aw,a^{-1}\lambda)$ for any number $a$, which implies a Killing vector of the form $\xi=\epsilon k(w\partial_{w}-\lambda\partial_{\lambda})$, where $k$ is a positive normalization constant determined from the calculation of the norm $\xi^{a}\xi_{a}$ and the factor $\epsilon$ is included in order to choose a future timelike Killing vector field.
 In accordance with the results of Sec.
-95, this also gives us $A(w)=kw$ and $B=0$ for the two $w-$dependent functions of the general Killing vector in equation (95).
+95, this also gives us $A(w)=kw$ and $B=0$ for the two $w-$dependent functions of the general Killing vector in Eq.
+(95).
 Calculation of the asymptotic limits of the norms gives us
@@
 $\lim_{\lambda\rightarrow\infty}\xi^{a}\xi_{a}=$ $-\left[\frac{2kr_{0}^{3}}{Q_{0}^{2}-r_{0}^{2}}\right]^{2}\;\;.$ (137)
-Therefore, the Killing vector is an asymptotic timelike unit Killing vector if the Killing vector has the form
+Therefore, the Killing vector is an asymptotic timelike unit vector if the Killing vector has the form
 $\xi=\frac{\epsilon}{2r_{0}}\left(1-\frac{Q^{2}_{0}}{r_{0}^{2}}\right)(w\partial_{w}-\lambda\partial_{\lambda})\;\;,$ (138)
-As mentioned above, if the norm of the Killing vector, $\xi^{a}\xi_{a}$ vanishes for $\lambda=0$, the null hypersurface $\mathcal{B}^{\epsilon}$ is a Killing horizon.
-If $\epsilon=1$, this boundary is the future Killing horizon $H^{+}$, and the past event horizon $H^{-}$ for the choice $\epsilon=-1$.
+As mentioned above, if the norm of the Killing vector $\xi^{a}\xi_{a}$ vanishes for $\lambda=0$, the null hypersurface $\mathcal{B}^{\epsilon}$ is a Killing horizon.
+If $\epsilon=1$, this boundary is the future Killing horizon $H^{+}$, and the past event horizon $H^{-}$ is for the choice $\epsilon=-1$.
 Its surface gravity $\kappa_{H}$ can be found from the invariant relation Wald (1984)
@@
 $\xi=\epsilon\kappa_{H}(w\partial_{w}-\lambda\partial_{\lambda}).$ (141)
-Note also that the metric (135) is regular at the null surface $w=0$ given by $\mathcal{N}^{\epsilon}(0)$.
+Note also that the metric (135) is regular at the null surface $w=0$ given by $\mathscr{N}^{\epsilon}(0)$.
 This surface is the future (past) Killing horizon $H^{+}$ ($H^{-}$) if $\epsilon=-1$ ($\epsilon=1$) with Killing vector field $\xi=\kappa_{H}\lambda\partial_{\lambda}$ ($\xi=\kappa_{H}w\partial_{w}$).
-On the other hand, $\xi$ vanishes at the 2-sphere given by $w=\lambda=0$, which constitutes the bifurcating Killing horizon.
+On the other hand, $\xi$ vanishes at the two-sphere given by $w=\lambda=0$, which is a bifurcating Killing horizon.
 V.1.2 Extremal black holes
-The surface gravity $\kappa_{H}$ in (140) vanishes if $Q_{0}=m_{B}=r_{0}$, but this goes along with the vanishing of the Killing vector in (138).
-In fact, in that limit the resulting metric associated to the $\{w,\lambda\}$ coordinates introduced in the last subsection is not the extremal Reissner-Nordström metric, but the so called Bertotti-Robinson metric Gallo et al.
+The surface gravity $\kappa_{H}$ in (140) vanishes, if $Q_{0}=m_{B}=r_{0}$, but this goes along with the vanishing of the Killing vector in (138).
+In fact, in that limit the resulting metric associated with the $\{w,\lambda\}$ coordinates introduced in the last subsection is not the extremal Reissner-Nordström metric, but the so-called Bertotti-Robinson metric Gallo et al.
 (2021).
-Thus the Killing vector (138) cannot be used to find an extremal Reissner-Nordström black hole solution where the global charge to mass ratio is unity.
+Thus the Killing vector (138) cannot be used to find an extremal Reissner-Nordström black hole solution where the global charge-to-mass ratio is unity.
 Indeed, as we will show here, the resulting extremal metric is not invariant under the transformation $(w,\lambda)\rightarrow(aw,a^{-1}\lambda)$, and therefore the resulting Killing vector field cannot be written as in (138).
-To single out an extremal black hole solution in the general solution (121) for $r$ and $V$ which also has a non-vanishing Killing vector at the horizon of the black hole, consider (130) while setting $m_{e}:=Q_{0}=m_{B}.$ This gives
-$m_{e}=\frac{m_{e}^{2}+r_{0}^{2}}{2r_{0}}-\epsilon r_{0}r_{N}\Theta$ (142)
+To single out an extremal black hole solution in the general solution (121) for $r$ and $V$, which also has a nonvanishing Killing vector at the horizon of the black hole, consider (130), while setting $m_{e}:=Q_{0}=m_{B}.$ This gives
+$m_{e}=\frac{m_{e}^{2}+r_{0}^{2}}{2r_{0}}-\epsilon r_{0}r_{N}\Theta,$ (142)
 allowing us to express any two of the initial data $(r_{0},\,r_{N},\Theta)$, by the third, i.e.
@@
 $\Theta=$ $\frac{(m_{e}-r_{0})^{2}}{2\epsilon r_{0}^{2}r_{N}}\;.$ (145)
-Since in the Schwarzschild case $r_{0}$ equates twice the Bondi mass, it is desirable to relate $r_{0}$ one-to-one with $m_{e}$.
-In (143) this is achieved only by either setting $\Theta=0$ or $r_{N}=0$ giving in both cases $r_{0}=m_{e}$.
+Since in the Schwarzschild case $r_{0}$ equals twice the Bondi mass, it is desirable to relate $r_{0}$ one-to-one with $m_{e}$.
+In (143) this is achieved only by setting either $\Theta=0$ or $r_{N}=0$ giving in both cases $r_{0}=m_{e}$.
 In (144), the choice $r_{0}=m_{e}$ gives $r_{N}=0$ and $\Theta\neq 0$ left arbitrary, while for (145) the same choice yields $\Theta=0$ and $r_{N}\neq 0$ left arbitrary.
 Thus an extremal black hole solution can be found if $r_{0}=m_{e}=Q_{0}$ and either $r_{N}=0$ and $\Theta\neq 0$ or $\Theta=0$ and $r_{N}\neq 0$.
-Now consider the solution for the areal distance and metric potential $V$ for these for cases: for the case $(r_{N}=0\;,\;\Theta\neq 0)$ we find
+Now consider the solution for the areal distance and metric potential $V$ for these cases: for the case $(r_{N}=0\;,\;\Theta\neq 0)$ we find
 $r=$ $m_{e}+\Theta\lambda$ (146a)
-$V=$ $\left(\frac{\lambda}{m_{e}+\Theta\lambda}\right)^{2}$ (146b)
+$V=$ $\left(\frac{\lambda}{m_{e}+\Theta\lambda}\right)^{2},$ (146b)
 and for the case $(\Theta=0\;,\;r_{N}\neq 0)$
@@
 $V=$ ${-\frac{4m_{e}\lambda^{2}[\epsilon\lambda r_{N}^{2}w^{3}-(m_{e}-2r_{N}w)(m_{e}+r_{N}w)^{3}]}{(r_{N}w+m_{e})[w^{4}r_{N}^{2}\lambda^{2}+4\epsilon r_{N}w^{2}(m_{e}+r_{N}nw)^{3}\lambda+4(m_{e}+r_{N}w)^{6}]}}\;.$ (148)
-The line element of first solution is
+The line element of the first solution is
 $\begin{split}ds^{2}=&-\left(\frac{\lambda}{m_{e}+\Theta\lambda}\right)^{2}dw^{2}+2\epsilon dwd\lambda\\ &+(m_{e}+\Theta\lambda)^{2}q_{AB}dx^{A}dx^{B}\;,\end{split}$ (149)
-but in fact it looks more familiar, if $r(\lambda)$ in (146a) is inverted to $\lambda(r)$, then we find
+but, in fact, it looks more familiar if $r(\lambda)$ in (146a) is inverted to $\lambda(r)$, and then we find
 $V(r)=\frac{1}{\Theta^{2}}\left(1-\frac{m_{e}}{r}\right)^{2}\;\;,\;\;$ (150)
-so that the coordinate transformation $\hat{w}=w/\Theta$, $\lambda=(r-m_{e})/\Theta$ leaves us with the well known line element
+so that the coordinate transformation $\hat{w}=w/\Theta$, $\lambda=(r-m_{e})/\Theta$ leaves us with the well-known line element
 $ds^{2}=-\left(1-\frac{m_{e}}{r}\right)^{2}d\hat{w}^{2}+2\epsilon d\hat{w}dr+r^{2}q_{AB}dx^{A}dx^{B}.$ (151)
-of an extremal Reissner-Nordström black hole with either ingoing($\epsilon=1$) or outgoing polar null coordinates.
-In such case, $\hat{w}$ is usually represented as $v$, and the past event horizon is reached as $v\to-\infty.$ For $\epsilon=1$, this metric covers the interior region of the past event horizon and the exterior region (in that case $\hat{w}$ is identified with the retarded null coordinate $u$).
+of an extremal Reissner-Nordström black hole with either ingoing ($\epsilon=1$) or outgoing ($\epsilon=-1$) polar null coordinates.
+In the first case, $\hat{w}$ is usually represented as $v$, and the past event horizon is reached as $v\to-\infty.$ For $\epsilon=1$, this metric covers the interior region of the past event horizon and the exterior region (in that case $\hat{w}$ is identified with the retarded null coordinate $u$).
 This coordinate does not cover the future horizon, which is asymptotically reached as $u\to\infty$.
 Indeed, calculation of the outgoing expansion rate of the surface forming null vector $\ell^{a}\partial_{a}=\partial_{\lambda}$ using (149) gives $2\Theta/(m_{e}+\Theta\lambda)$ which does not vanish for any value of $w$.
-Thus there is no null surface $w=const$ labelling the future event horizon.
+Thus there is no null surface $w=$const labeling the future event horizon.
 In a similar way, for ingoing null coordinates ($\epsilon=1$), the resulting coordinates cover the exterior and interior of the future event horizon.
@@
 However, we will see that the description of an extremal charged black hole with both horizons is possible for initial data with $\Theta=0$ and $r_{N}\neq 0$.
-Consider now the case $\Theta=0$ which gives (147) and (147) for the non trivial components of the metric.
+Consider now the case $\Theta=0$, which gives (147) and (148) for the non trivial components of the metric.
 From (147) and (36a), we read off the relevant functions, $H$ and $R_{\infty}$ of the asymptotic expansion of $r$ as
@@
 $R_{\infty}(w)$ $=$ $m_{e}+r_{N}w\;\;.$ (153)
-Note also, that the surface $w=0$ is a regular null surface determined by $r=m_{e}$ and where $V=\frac{\lambda^{2}}{m^{2}_{e}}$.
-On the other hand, from (152) and (107), we find the following asymptotic behaviour of the Killing vector near $w=0$
+Also note the surface $w=0$ is a regular null surface determined by $r=m_{e}$ and where $V=\frac{\lambda^{2}}{m^{2}_{e}}$.
+From (152) and (107), we find the following asymptotic behavior of the Killing vector near $w=0$
 $\xi=\left[\frac{c^{2}-1}{cr_{N}w^{2}}{(2r_{N}w+m_{e})m_{e}-\frac{r_{N}}{c}}\right]\partial_{\lambda}+O(w)\;,$ (154)
@@
 The asymptotic norm of the Killing vector gives a timelike unit vector by construction.
-Observe the difference to the nonextremal case (138), the null surface $\lambda=0$ is not a Killing horizon for (155).
+Observe the difference to the nonextremal case (138), where the null surface $\lambda=0$ is not a Killing horizon for (155).
 In fact, the boundary surface has an expansion rate proportional to $r_{N}$.
 An additional surface (to $w=0$) where the Killing vector becomes null is at
-$\lambda(w)=-\frac{2\epsilon}{w(w+m_{e})^{2}}.$ (156)
+$\lambda(w)=-\frac{2\epsilon(r_{N}w+m_{e})^{2}}{w}.$ (156)
 which follows from the calculation of (101). This is the Killing horizon $r=m_{e}$.
-Indeed, the parameter $r_{N}$ in (147) and (148) can be eliminated using the scaling relations $w\rightarrow w/r_{N}$ and $\lambda\rightarrow r_{N}\lambda$, so that the final solution for the extremal Reissner-Nordström black hole only depends on the physical parameter $m_{e}$, the Bondi mass, so that
-$r=$ $m_{e}+w+\epsilon\left[\frac{w^{2}}{2(w+m_{e})^{2}}\right]\lambda$ (157a)
-$V=$ $-\frac{4m_{e}\lambda^{2}}{(m_{e}+w)}\left[\frac{\epsilon\lambda w^{3}-(m_{e}-2w)(m_{e}+w)^{3}}{\lambda^{2}w^{4}+4\epsilon\lambda w^{2}(m_{e}+w)^{3}+4(m_{e}+w)^{6}}\right]$ (157b)
+Indeed, the parameter $r_{N}$ in (147) and (148) can be eliminated using the scaling relations $w\rightarrow w/r_{N}$ and $\lambda\rightarrow r_{N}\lambda$, so that the final solution for the extremal Reissner-Nordström black hole depends only on the physical parameter $m_{e}$, the Bondi mass, so that
+$r=$ $m_{e}+w+\epsilon\left[\frac{w^{2}}{2(w+m_{e})^{2}}\right]\lambda,$ (157a)
+$V=$ $-\frac{4m_{e}\lambda^{2}}{(m_{e}+w)}\left[\frac{\epsilon\lambda w^{3}-(m_{e}-2w)(m_{e}+w)^{3}}{\lambda^{2}w^{4}+4\epsilon\lambda w^{2}(m_{e}+w)^{3}+4(m_{e}+w)^{6}}\right],$ (157b)
 $\xi$ $=\frac{\epsilon w^{2}}{2(w+m_{e})^{2}}\partial_{w}-\left[\frac{\epsilon m_{e}(w\lambda)}{(w+m_{e})^{3}}+1\right]\partial_{\lambda}.$ (157c)
 This solution was also obtained by us in Gallo et al.
-(2021) using an alternative hierarchical system of equations, note however that the $w$ used here must be identified with the $-w$ used in that reference.
+(2021) using an alternative hierarchical system of equations; note, however, that the $w$ used here must be identified with the $-w$ used in that reference.
 In difference to the traditional outgoing/ingoing Eddington-Finkelstein coordinates, the $\{w,\lambda,x^{A}\}$ is a regular coordinate system at both horizons.
 V.2 FJNW solution in null coordinates
-The simplest Killing vector (107) is where $A=1$ and $B=0$.
+The simplest Killing vector (107) has $A=1$ and $B=0$.
 The Killing vector becomes $\xi=\partial_{w}$ and its additional requirement to be asymptotically timelike restricts the asymptotic frame to be a Bondi frame, i.e.
@@
 (1968); Wyman (1981); Virbhadra (1997); Ota et al.
-(2022) in affine null coordinates, which has not been derived explicitly as a solution of an hierarchical system of the Einstein-scalar field equations in affine null coordinates.
+(2022) in affine-null coordinates, which has not been derived explicitly as a solution of a hierarchical system of the Einstein-scalar field equations in affine-null coordinates.
 Considering $\xi=\partial_{w}$ for an uncharged ($q=0$), real scalar field, the system of equations (29) becomes
@@
 where we used that $\mathcal{L}\rightarrow L$ becomes real. The wave equation gives (20)
-$(r^{2}V\Phi_{,\lambda})_{,\lambda}=0$ (162)
+$(r^{2}V\Phi_{,\lambda})_{,\lambda}=0.$ (162)
 In the absence of a scalar field, $\Phi=0$, the solution is trivial, yielding the Schwarzschild solution $V=1+b/r$ with $r=\lambda$, where $b$ is a constant Mädler and Gallo (2023).
@@
 Although $\Phi$ in (164) is a test scalar field, we can use this ansatz for the scalar field initial data to solve the complete Einstein-scalar field equations in a consistent way.
-In particular, note that if $r^{2}V$ has the same form for the complete system like in the Schwarzschild case, the scalar field will have the same form as in Schwarzschild.
+In particular, note that if $r^{2}V$ has the same form for the complete system as in the Schwarzschild case, the scalar field will have the same form as in the Schwarzschild case.
 The solution of (158) gives us
@@
 $\delta=\sqrt{1-\frac{2\kappa a^{2}}{b^{2}}}.$ (166)
-Reality of $\delta$ requires $2\kappa a^{2}<b^{2}$. Note that if $a=0$, the scalar field vanishes, and therefore the solution must reduce to the Schwarzschild metric. In that case (165) reduces to
+The reality of $\delta$ requires $2\kappa a^{2}<b^{2}$. Note that if $a=0$, the scalar field vanishes, and therefore the solution must reduce to the Schwarzschild metric. In that case (165) reduces to
 $r=(c_{1}+c_{2})\lambda+c_{2}b.$
@@
 The asymptotic limit of (167) gives
-$r=\lambda+\frac{b}{2}\left(1-\delta\right)+O(\lambda^{-1})$ (168)
+$r=\lambda+\frac{b}{2}\left(1-\delta\right)+O(\lambda^{-1}),$ (168)
 which from (36a) implies
@@
 $R_{\infty}$ $=$ $\frac{b}{2}\left(1-\delta\right).$ (170)
-Next, using the notation as in (36d), integration of (159) yields $Z(\lambda)=Z_{\infty}=const,$ which is related to the Bondi mass by (41), i.e.
+Next, using the notation as in (36d), the integration of (159) yields $Z(\lambda)=Z_{\infty}$=const, which is related to the Bondi mass by (41), i.e.
 $Z_{\infty}=R_{\infty}-2m_{B}=\frac{1}{2}\left[b(1-\delta)-4m_{B}\right].$ (171)
@@
 The integration constant $L_{\infty}$ vanishes because of the Killing vector $\partial_{w}$ and considering (39). Next, integration of (161) gives us
-$\begin{split}V=&V_{1}\lambda+V_{0}+\frac{2\lambda}{b(\delta-1)}\left(1+\frac{b}{\lambda}\right)^{\delta}\\ &+\frac{2Z_{\infty}\left[b(\delta-1)-2\lambda\right]}{b^{2}(\delta^{2}-1)}\left(1+\frac{b}{\lambda}\right)^{\delta}\end{split}$ (173)
+$\begin{split}V=&V_{1}\lambda+V_{0}+\frac{2\lambda}{b(\delta-1)}\left(1+\frac{b}{\lambda}\right)^{\delta}\\ &+\frac{2Z_{\infty}\left[b(\delta-1)-2\lambda\right]}{b^{2}(\delta^{2}-1)}\left(1+\frac{b}{\lambda}\right)^{\delta}.\end{split}$ (173)
 Performing an asymptotic expansion of $V$ as in (36g) gives
-$\begin{split}V=&\left[V_{1}+\frac{(2\delta+2)b-4Z_{\infty}}{b^{2}(\delta^{2}-1)}\right]\lambda+\left[V_{0}+\frac{2b\delta-2Z_{\infty}}{(\delta-1)b}\right]\\ &+O(\lambda^{-1})\end{split}$ (174)
+$\begin{split}V=&\left[V_{1}+\frac{(2\delta+2)b-4Z_{\infty}}{b^{2}(\delta^{2}-1)}\right]\lambda+\left[V_{0}+\frac{2b\delta-2Z_{\infty}}{(\delta-1)b}\right]\\ &+O(\lambda^{-1}).\end{split}$ (174)
 Taking into account (37), (38) and the Killing symmetry implies $V_{\infty}=1$ and $U_{\infty}=0$, which gives us
-$V_{0}=$ $\frac{2Z_{\infty}-(1+\delta)b}{(\delta-1)b}$ (175)
-$V_{1}=$ $\frac{2(1+\delta)b-4Z_{\infty}}{b^{2}(1-\delta^{2})}$ (176)
+$V_{0}=$ $\frac{2Z_{\infty}-(1+\delta)b}{(\delta-1)b},$ (175)
+$V_{1}=$ $\frac{2(1+\delta)b-4Z_{\infty}}{b^{2}(1-\delta^{2})}.$ (176)
 These relations allow us to write $V$ as
 $\begin{split}V=&\frac{2\lambda\Bigl{[}b(\delta+1)-2Z_{\infty}\Bigr{]}}{b^{2}(\delta^{2}-1)}\left[\left(\frac{\lambda+b}{\lambda}\right)^{\delta}-1\right]\\ &+\frac{2Z_{\infty}}{b(\delta+1)}\left(\frac{\lambda+b}{\lambda}\right)^{\delta}.\end{split}$ (177)
-Finally, the asymptotic expansion of (177) which gives
+Finally, from the asymptotic expansion of (177), which gives
 $V=1+\frac{b\delta}{\lambda}+O(\lambda^{-2}),$ (178)
@@
 $Z_{\infty}=\frac{b(\delta+1)}{2}.$
-By replacing this values into (177) we arrive to our final expression for $V$
+By replacing these values into (177) we arrive to our final expression for $V$
 $V=\left(1+\frac{b}{\lambda}\right)^{\delta}.$ (179)
@@
 We discussed the CIBVP in spherical symmetry of the Einstein equations minimally coupled to a complex and charged massless scalar field.
-Charting spacetime with an affine null metric at a family of null hypersurfaces, the field equations can be cast into a hierarchical system of hypersurface-evolution equations of which the evolution equation is a first order transport equation for the complex initial data of this system.
+Charting spacetime with an affine-null metric at a family of null hypersurfaces, the field equations can be cast into a hierarchical system of hypersurface-evolution equations of which the evolution equation is a first order transport equation for the complex initial data of this system.
 The initial data consist of an arbitrary profile of the scalar field on an initial null hypersurface.
 The boundary values required to solve the hypersurface equations are determined from the type of boundary at which the family of null hypersurfaces is specified.
-Different CIBVPs were discussed; i) an asymptotic CIBVP for large values of the affine parameter, ii) a CIBVP at the central geodesic of spherical symmetry, and iii) two realizations of double null CIBVPs.
+Different CIBVPs were discussed; i) an asymptotic CIBVP for large values of the affine parameter, ii) a CIBVP at the central geodesic of spherical symmetry, and iii) two realizations of double-null CIBVPs.
 The solution of the asymptotic CIBVP is represented by two functions, one for a redshift factor $H$ and another being the complex scalar news function $\mathcal{N}_{\mathcal{B}}$ (which is the $w$ derivative of the scalar field’s monopole) and in addition there are three scalar values at a given coordinate value $w=w_{0}$, representing the complex monopole of the scalar field, the initial Bondi mass and the initial value of the global charge of the system.
 The redshift factor indicates whether an event horizon forms, i.e.
-$H\rightarrow 0$, or whether the solution approaches asymptotically Minkowkian values, $H\rightarrow 1$.
-Given a continous redshift function $H>0$ for $w\geq w_{0}$, an asymptotic inertial coordinate system can be established in which $H=1$ for $w\geq 0$.
+$H\rightarrow 0$, or whether the solution approaches asymptotically Minkowskian values, $H\rightarrow 1$.
+Given a continuous redshift function $H>0$ for subsequent values of $w$ within the region of integration, an asymptotic inertial coordinate system can be established in which $H=1$ in that region.
 Such coordinate frame is called a Bondi frame.
@@
 Evaluation of the field equations in a Bondi frame significantly simplifies the solution of the field equation.
-In such Bondi frame, it is easy to see that the news function completely determines the future($\epsilon=-1$)/past($\epsilon=+1$) behaviour of the scalar monopole, the Bondi mass and the global charge.
-This behaviour is the spherically symmetric scalar field analogue to the asymptotic CIBVP for electro-vacuum space times, see e.g.
-van der Burg (1969); Bicak and Pravdova (1998).
+In such Bondi frame, it is easy to see that the news function completely determines the future($\epsilon=-1$)/past($\epsilon=+1$) behavior of the scalar monopole, the Bondi mass and the global charge.
+This behavior is the spherically symmetric scalar field analog to the asymptotic CIBVP for electrovacuum space times (see e.g.
+van der Burg (1969); Bicak and Pravdova (1998)).
 If the boundary degenerates to the central world line of spherical symmetry so that the family of null hypersurfaces are a family of null cones whose vertices are traced by the world line, the regularity conditions at the vertices completely fix the boundary conditions to constant values along the world line.
 As a consequence, the local evolution of the scalar field along the world line is fully determined by the (first) derivative of the scalar field at the world line.
-As such, by specification of finite initial data at the vertex of an initial null cone $w=w_{0}$ and requiring the local regularity conditions for the hypersurface variables for $w>w_{0}$, one can completely determine the evolution of the scalar field and the resulting structure of the metric for all values $w>w_{0}$.
+As such, by specification of finite initial data at the vertex of an initial null cone $w=w_{0}$ and requiring the local regularity conditions for the hypersurface variables for other values of $w$ within the region of integration, one can completely determine the evolution of the scalar field and the resulting structure of the metric in that region.
 Numerical simulations of the real analog of the presented hypersurface-evolution system at a world line was presented first in Crespo et al.
@@
 (2024) also showed the regular center can be maintained in spacetimes forming black holes.
-The double null CIBVP was considered for the two cases where the Misner-Sharp mass of the null boundary surface is either trivial or non-trivial.
+The double-null CIBVP was considered for the two cases where the Misner-Sharp mass of the null boundary surface is either trivial or nontrivial.
 In both cases, the only required initial data on given null hypersurface $w=w_{0}$ is the first $\lambda-$derivative of the scalar field on the initial data surface.
 At the common intersection of the initial data surface $w=w_{0}$ and the null boundary, a value for the scalar field must be specified.
-Moreover, an analogue to the news function, referred to as boundary news function, needs to be specified to be able to find the boundary values required to solve the hypersurface-evolution system.
+Moreover, an analog to the news function, referred to as the boundary news function, needs to be specified to be able to find the boundary values required to solve the hypersurface-evolution system.
 Like for the asymptotic CIBVP, this boundary news function needs to be specified a priori and is completely independent from the initial data profile on $w_{0}$.
-For the case, where null boundary surface has vanishing Misner-Sharp mass, in fact all but two of these boundary values are zero.
-One nontrivial and constant boundary value is the first $\lambda-$derivative of the areal distance that can be set to unity for $w\geq w_{0}$.
-The other one is the first derivative of $g^{\lambda\lambda}$ evaluated at the null boundary, but this field only needs to be specified as initial value at the cross section between initial data surface and null boundary.
-This boundary value is in fact the inaffinity of the generating null vectors of the boundary surface where $w=w_{0}$.
-Its values for $w>w_{0}$ on the boundary are completely determined by the boundary news function.
-To summarize, the double-null initial-boundary value problem with a null boundary of zero Misner sharp mass requires (apart from trivial and unity values) the specification of the scalar field value and inaffinity at the intersection of the initial data surface and the null boundary, the specification of a free news function on the boundary and the specification of the $\lambda-$derivative of the scalar field on the initial data surface.
+For the case, where the null boundary surface has vanishing Misner-Sharp mass, in fact all but two of these boundary values are zero.
+One nontrivial and constant boundary value is the first $\lambda-$derivative of the areal distance that can be set to unity without loss of generality.
+The other one is the first derivative of $g^{\lambda\lambda}$ evaluated at the null boundary, but this field only needs to be specified as the initial value at the cross section between the initial data surface and the null boundary.
+This boundary value is, in fact, the inaffinity of the generating null vectors of the boundary surface, where $w=w_{0}$.
+Its values for $w\neq w_{0}$ on the boundary are completely determined by the boundary news function.
+To summarize, the double-null initial-boundary value problem with a null boundary of zero Misner-Sharp mass requires (apart from trivial and unity values) the specification of the scalar field value and inaffinity at the intersection of the initial data surface and the null boundary, the specification of a free news function on the boundary and the specification of the $\lambda-$derivative of the scalar field on the initial data surface.
 If the Misner-Sharp mass of the boundary is nonzero, the values giving rise to the mass and expansion rates of the generators of the boundary and initial data surface are to be specified at the cross section of the two surfaces.
@@
 The evolution of the boundary values is completely decoupled from the evolution of the initial data via the hypersurface evolution equations and it can therefore be done stand alone and prior to any evolution of the initial data.
-This is in contrast to the timelike null problem at the world line where the evolution of the scalar field along the world line requires data (the $\lambda$ derivative) depending on the neighborhood of the world line.
-The different initial boundary value formulations can be used for numerically investigating different physical scenaria.
-The worldline-null problem is suited for studies of charged black hole formation and critical phenomena based on regular and asymptotically flat initial data of a charged scalar field.
+This is in contrast to the timelike null problem at the world line, where the evolution of the scalar field along the world line requires data (the $\lambda$-derivative) depending on the neighborhood of the world line.
+The different CIBVP formulations can be used for numerically investigating different physical scenaria.
+The world-line-null problem is suited for studies of charged black hole formation and critical phenomena based on regular and asymptotically flat initial data of a charged scalar field.
 Currently some of these studies are ongoing as an extension of the previous work employing a real scalar field Mädler et al.
 (2024).
-This work used a coordinate compactification of the affine parameter together with some rescaling of the metric field to map null infinity to a finite value.
+This work used a coordinate compactification of the affine parameter together with some rescaling of the metric field to map null infinity to a finite value Mädler and Gallo (2025).
 Regarding the formation of charged black holes, it is natural to ask how close to unity the charge-to-mass ratio of such black holes can be found by evolving charged scalar fields.
@@
 But, previous investigations only reach values of $\sim 0.6$ for the charge-to-mass ratio Torres and Alcubierre (2014).
-The double null CIBVP with a massive null boundary is well suited to investigate scattering Baake and Rinne (2016) or perturbations of a spherically symmetric charged black hole Gelles and Pretorius (2025) since the horizon of the black hole can be taken as a boundary surface.
+The double-null CIBVP with a massive null boundary is well suited to investigate scattering Baake and Rinne (2016) or perturbations of a spherically symmetric charged black hole Gelles and Pretorius (2025) since the horizon of the black hole can be taken as a boundary surface.
 Given the recent success of using null geodesics stability analysis via Lyapynov exponents Ianniccari et al.
 (2024) to explain black hole formation, critical phenomena Choptuik (1993) or even black hole shadows Gallo and Mädler (2025), it may be interesting to explore how the presented CIBVP can be adopted to investigate such phenomena numerically.
-A formulation of a double null CIBVP at a horizon may be an option for such study.
-To better understand a spherically symmetric charged black hole in affine null coordinates, we derived the subextremal and extremal Reissner-Nordström solution using the developed affine-null metric double null CIBVP.
+A formulation of a double-null CIBVP at a horizon may be an option for such study.
+To better understand a spherically symmetric charged black hole in affine-null coordinates, we derived the subextremal and extremal Reissner-Nordström solution using the developed affine-null metric double-null CIBVP.
 For that aim, we also looked at the general solution of Killing equations.
 It turns out that there is an entire family of Killing vectors depending on two free functions depending on the $w$ coordinate.
-In deed, a restriction of the Killing vectors to be asymptotically timelike, while using the general solution of the asymptotically flat initial boundary value problem, showed that the frame at large radii does not need to necessarily be a Bondi frame, where the redshift factor is unity.
-This is in fact only the case if the Killing vector has the simple form $\partial_{w}$.
-For subextremal and extemal black holes, the results of Sec.
+Indeed, a restriction of the Killing vectors to be asymptotically timelike, while using the general solution of the asymptotically flat initial boundary value problem, showed that the frame at large radii does not necessarily need to be a Bondi frame, where the redshift factor is unity.
+This is, in fact, only the case if the Killing vector has the simple form $\partial_{w}$.
+For subextremal and extemal black holes, our results of Sec.
 V.1 show that corresponding Killing vectors have much more elaborated functional dependence.
@@
 We also thank O.
-Baake for reading and commenting on the manuscript.
+Baake for reading and commenting of the manuscript.
+DATA AVAILABILITY
+No data were created or analyzed in this study.